### v3

# Geometric measures of CHSH quantum nonlocality: characterization, quantification, and comparison by distances and operations

Abstract
We introduce a geometric framework for analyzing Bell nonlocality in Hilbert space, aiming to provide intrinsic measures of nonlocality based solely on the properties of quantum states for a given Bell inequality. Recognizing that nonlocality is inherently measurement-dependent, our approach focuses on fixing a specific Bell scenario—that is, a chosen number of measurements and outcomes per party, along with a specific Bell inequality—and, for each quantum state, maximizing the violation over all possible measurement choices within the given scenario. We then characterize the geometry of the corresponding set of local states, defined as those states for which no violation occurs for any choice of measurements. Within this framework, we define geometric and entropic measures of nonlocality based on the distance of a given quantum state from the set of local states, using quantities such as trace distance, Hilbert-Schmidt distance, and relative entropy. We first establish the general formalism, emphasizing the challenges posed by the NP-hardness of characterizing local sets for arbitrary Bell scenarios. We then specialize to the case of the Clauser-Horne-Shimony-Holt (CHSH) inequality in two-qubit systems, where the local set is fully characterized and an explicit geometric analysis is possible. In this setting, we derive geometric measures of nonlocality for Bell-diagonal and Werner states, showing that the closest local state to a Werner state remains a Werner state, and similarly for Bell-diagonal states. For each quantum states, our results offer a metric-based characterization of Bell nonlocality that is independent of specific experimental implementations and naturally extend to different Bell scenarios when the corresponding set of local states is known.
I Introduction
Quantum nonlocality is generally understood as the impossibility of describing experimental correlations using a local hidden variable model Brunner2014 . A state is considered nonlocal if there exists at least one experimental configuration that leads to a violation of a Bell-type inequality, such as the Clauser-Horne-Shimony-Holt (CHSH) inequality. However, nonlocality is not solely a property of the state itself but also depends on the chosen measurement settings, making its geometric characterization particularly challenging. For instance, it is well known that some states do not violate the CHSH inequality, where both Alice and Bob perform two dichotomic measurements (the 2-2-2-2 scenario), but do violate more general Bell inequalities, such as the 3-3-2-2 scenario.
Furthermore, there exist quantum states that do not violate any Bell inequality individually, but when multiple copies of the state are considered along with joint measurements, the resulting system can exhibit a violation of some Bell inequality Popescu1995 ; Gisin1996 ; Masanes2008 . This phenomenon is known as hidden nonlocality, or superactivation of Bell nonlocality. For instance, it has been shown that a state is distillable if and only if, after applying joint local filtering operations on an arbitrarily large number of copies, it leads to a violation of some Bell inequality Masanes2006 . Another example of nonlocality superactivation has been introduced in a recent work Bavaresco2025 , where the authors consider an appropriate (Bell-local) ancilla, a so-called catalyst, and local operations on the combined system.
A natural approach to quantify the amount of nonlocality of a given state, is to study the geometry of the set of local states within Hilbert space while fixing a particular Bell inequality. Specifically, the distance of a given quantum state from this set provides a meaningful geometric measure of its nonlocality for the chosen protocol. Various distance-based measures, such as trace distance, Hilbert-Schmidt distance, and relative entropy, can be employed to quantify the deviation of a state from the local set.
In this work, we explore the geometric and entropic characterizations of quantum correlations beyond entanglement Horodecki2009 , specifically focusing on Bell nonlocality within Hilbert space. Since entanglement and Bell nonlocality coincide for pure states, geometric approaches to entanglement naturally extend to nonlocality in such cases. However, for mixed states, while Bell nonlocality always implies entanglement, the converse does not necessarily hold, as exemplified by Werner states Werner1989 . This establishes a strict hierarchy between these two forms of quantum correlations in finite-dimensional systems.
In the geometric approach such as the one introduced in the present work, the number of measurements for each party, their possible outcomes, and the related Bell inequalities are uniquely determined. Thereafter, for any given state, one possible choice is to maximize the violation of the inequality over all possible measurements performed by the two parties. If no violation of all the different Bell inequalities occurs, the state is considered local. The corresponding geometric measure of Bell nonlocality will then depend solely on the geometry of the set of local states and on the distance of the state from the set. Alternatively, for each possible type of measurement one can proceed to define the corresponding set of local states and the corresponding, measurement-dependent geometric measure of Bell nonlocality.
In principle, the geometric approach described in the present work is very general, as it applies to any form of quantum nonlocality identified by some type of Bell inequality. What one needs to know is, for each instance, the type of measurement considered, and the corresponding set of local states. As already mentioned, this may turn out to be a very daunting task, given that in general the problem of determining the set of local states associated to a given Bell inequality is NP-hard in the number of measurements and, depending on the Hilbert space dimension, in the number of measurement outcomes.
In the following, we will illustrate the general method by restricting to the CHSH inequality with maximum violation of the inequality over all possible measurements. Notably, for two-qubit systems, a complete parameterization of the set of local states with respect to the CHSH inequality is known, enabling precise geometric analysis. Clearly, this is not the most general case, as other forms of Bell inequalities exist. On the other hand, at least in the case of two qubits, it has been shown that the general problem is NP-hard, and therefore, the analysis must necessarily be limited to a relatively small number of measurements. In this context, it is well known that the number of inequalities other that the CHSH one is very limited Collins2004 . For example, in the case of three possible outcomes on each particle, there is only one other type of Bell inequality that arises and that is distinct from the CHSH one.
We will proceed by first introducing the general theoretical framework before focusing on the violation of the CHSH inequality in two-qubit systems. For such systems, the set of local states is fully characterized, allowing us to illustrate our method effectively. We explicitly compute geometric measures of nonlocality for Bell-diagonal states and Werner states, relying on two key results that we establish: the closest local state to a Werner state remains a Werner state, and the closest local Bell-diagonal state remains a Bell-diagonal state.
The work is organized as follows: In section II, we briefly review the basic concepts and structures in the theory of Bell nonlocality. Then, in section III, we introduce the set of free operations as the set of operations that map the set of local states into itself. In section IV we introduce the geometric and entropic measures of nonlocality and discuss their properties. We address the specific case of the CHSH inequality for two qubits in sections V and VI. Finally, in section VII we discuss our results and comment on some possible upcoming research directions.
II Bell nonlocality for dichotomic observables: CHSH inequality
Quantum nonlocality naturally arises in the context of Bell inequalities Bell1964 ; Clauser1969 ; Laskowski2004 . Consider two observers, Alice and Bob, who are spatially separated. Each of them has a set of observables, with Alice having $n_{a}$ and Bob having $n_{b}$. They do not communicate and independently choose which observable to measure, recording their respective outcomes. The correlations between these observables, determined by their choices and the results obtained, are used to construct Bell inequalities.
These inequalities must hold in a classical framework, where we assume at most local hidden variable models and other commonly accepted assumptions, including choice independence, realism, and the absence of experimental loopholes. In bipartite local hidden variable theories, the conditional probability of Alice choosing observable $x$ and obtaining result $a$, and Bob choosing observable $y$ and obtaining result $b$, satisfies the following relation:
$p(a,b|x,y)=\sum_{\lambda}p(\lambda)p(a|x,\lambda)p(b|y,\lambda),$ (1)
where we can express $p(a,b|x,y)$ in quantum mechanical terms as:
$p(a,b|x,y)={\rm Tr}(\rho^{AB}M_{a|x}\otimes M_{b|y}).$ (2)
Here, $M_{a|x}$ represents a positive operator-valued measure (POVM) associated with outcome $a$ for observable $x$, and similarly, $M_{b|y}$ corresponds to outcome $b$ for observable $y$. It follows directly that for separable states, Eq. (2) implies Eq. (1). Consequently, nonlocal states must be entangled Horodecki1995 ; Werner1989 . However, some entangled states may still exhibit locality Werner1989 ; Masanes2006 .
In the case of two dichotomic observables per observer, the Bell inequality takes the CHSH form:
$E(x,y)-E(x,y^{\prime})+E(x^{\prime},y)+E(x^{\prime},y^{\prime})\leq 2,$ (3)
where $E(x,y)$ denotes the statistical expectation value of the measurement outcomes for observables $x$ and $y$ in a given quantum state. For nonlocal states, Eq. (2) does not hold, leading to a violation of the CHSH inequality, Eq. (3).
III Bell nonlocality: the free operations
The set of local states, i.e., those states that do not violate any Bell inequality, will be denoted in the text as $\mathcal{L}$. In the framework of resource theory Chitambar2019 ; Theurer2017 ; Theurer2021 ; Gallego2015 ; Wolfe2020 ; Dakic2012 ; Pirandola2014 ; Streltsov2017 ; Buscemi2020 ; Contreras2019 ; Gour2008 ; Brandao2013 ; Buscemi2023 ; Barrett2005 ; Schmid2020 ; Kuroiwa2020 ; Gonda2023 ; Bartlett2007 ; DeVicente2014 , local states are also referred to as free states, meaning that they do not possess any resource—in this case, nonlocality.
In the case of two-qubit systems and the CHSH inequality (3), the set of local states can be fully characterized Horodecki1995 ; Masanes2008 and will be discussed in detail in Section V.3 below.
The set of operations that map local states into local states, also known as the set of free operations, consists of local operations and shared randomness (LOSR) Chitambar2019 ; Schmid2020 . These operations can be generally defined as follows:
[…]
where $\Phi^{A}_{\lambda}$ and $\Phi^{B}_{\lambda}$ are completely positive and trace-preserving (CPTP) maps. These transformations cannot take a local state outside the set of local states, as they preserve the hidden variable model. Since LOSR operations are also LOCC (while the converse is not true), they inherit many properties established for the extensively studied class of LOCC transformations. There are a few general properties that free operations should satisfy Theurer2021 ; Chitambar2019 :
  • Free operations preserve the set of free states.
  • Their concatenation is still a free operation.
  • The trivial extension of a free operation is also a free operation (as a consequence, tensor products of free operations are free operations).
  • The operation of adding an auxiliary system is free.
  • The operation of discarding a subsystem (partial trace) is free.
In particular, LOSR satisfy all the properties listed above.
IV Geometric and Entropic Measures of Bell nonlocality
IV.0.1 Geometric measures
For a given quantum state $\rho$ we take as measure $\mathcal{M}(\rho)$ of Bell nonlocality any contractive distance (with the exception of the Hilbert-Schmidt distance) $D$ between $\rho$ and the closest local state $\rho_{loc}$:
$\mathcal{M}(\rho)=\min_{\rho_{loc}\in\sigma}D(\rho,\rho_{loc})\,,$ (5)
where $\sigma\equiv\{\rho_{loc}\}$ denotes the set of local states.
Physically meaningful and mathematically sound geometric measures of quantum nonlocality should satisfy the same key properties enjoyed by the bona fide entropic and geometric measures of quantum entanglement and other quantum correlations (e.g. quantum discord) introduced in the course of time Plenio2007 ; Christandl2004 ; Eisert1999 ; Wootters1998 ; Plenio2005 ; Wei2003 ; Illuminati2008 ; Uyanik2010 ; Geomdiscord2016 ; Geomdiscord2017 ; Vedral1997 ; Vedral1998 ; Baumgratz2014 ; Bengtsson2017 . In particular:
  • Property 1 – vanishing on local states. By definition, this property is automatically satisfied by the relative entropy and by any chosen distance.
  • Property 2 – monotonicity under LOSR. For any distance $D$ that is monotonically nonincreasing under an arbitrary CPTP map $\Phi$ that preserves the locality of a local state (LOSR do satisfy this property) we have
    $\mathcal{M}(\rho)=\min_{\rho_{loc}\in\sigma}D(\rho,\rho_{loc})=D(\rho,\rho_{loc}^{min})\geq D(\Phi(\rho),\Phi(\rho_{loc}^{min}))\geq\min_{\rho_{loc}\in\sigma}D(\Phi(\rho),\rho_{loc})=\mathcal{M}(\Phi(\rho)).$ (6)
    It follows that any distance that is nonincreasing under CPTP maps defines a measure of nonlocality that is nonincreasing under LOSR. Property 2 holds for the relative entropy as well, since the latter is contractive under CPTP maps by construction.
  • Property 3 – convexity (nonincreasing under mixing of quantum states). Since the set of local states is convex any distance that is jointly convex induces a measure of nonlocality that is convex. Denoting by $a$ the convexity parameter with $0\leq a\leq 1$, we have:
    $a\mathcal{M}(\rho_{1})+(1-a)\mathcal{M}(\rho_{2})$ $=$ $a\min_{\rho_{loc}\in\sigma}D(\rho_{1},\rho_{loc})+(1-a)\min_{\rho_{loc}\in\sigma}D(\rho_{2},\rho_{loc})$ (7)
    $=$ $aD(\rho_{1},\rho_{loc}^{min})+(1-a)D(\rho_{2},\rho_{loc}^{min^{\prime}})$ (8)
    $\geq$ $D(a\rho_{1}+(1-a)\rho_{2},a\rho_{loc}^{min}+(1-a)\rho_{loc}^{min^{\prime}})$ (9)
    $\geq$ $\min_{\rho_{loc}\in\sigma}D(a\rho_{1}+(1-a)\rho_{2},\rho_{loc})$ (10)
    $=$ $\mathcal{M}(a\rho_{1}+(1-a)\rho_{2}).$ (11)
    Joint convexity holds for the relative entropy by definition, and thus the latter is automatically nonincreasing under mixing.
  • Property 4 – monotonicity under subselection. We require monotonicity under selective measurements on average:
    $\mathcal{M}(\rho)\geq\sum_{i}p_{i}\mathcal{M}\left(\frac{\Lambda_{i}\rho\Lambda_{i}^{\dagger}}{p_{i}}\right),$ (12)
    where the weights $p_{i}={\rm Tr}(\Lambda_{i}\rho\Lambda_{i}^{\dagger})$ and the set $\{\Lambda_{i}\}$ of Kraus operators is such that $\sum_{i}\Lambda_{i}^{\dagger}\Lambda_{i}=\mathbb{I}$ and $\Lambda_{i}\rho\Lambda_{i}^{\dagger}\in\sigma$ $\forall\rho^{loc}\in\sigma$. In a sense, property 4 might be considered even more important than property 2, since subselection is a process often readily available under controlled experimental conditions. On the other hand, the former is usually significantly harder to verify than the latter, which is certainly satisfied by all well behaved geometric and entropic measures. It should be noted that convexity (property 3) and monotonicity under subselection (property 4), automatically imply monotonicity under LOSR (property 2). The reverse is obviously not true in general. Focusing only on the relative entropy, proof of property 4 is quite straightforward, in analogy with what has been shown to hold for the resource theory of entanglement Vedral1998 and quantum coherence Baumgratz2014 , and we provide it in Section IV.1.
  • Property 5 – computability. This property depends on the distance. We will study this subject case by case in the following by computing explicitly various measures for selected classes of two-qubit states.
In the following sections, we introduce geometric and entropic measurements and analyze the properties they satisfy.
IV.0.2 Relative entropy
Relative entropy is not a distance as it does not satisfy the triangle inequality and is not symmetric with respect to state interchange, but nonetheless it is widely used in the quantum information theory and the theory of entanglement because it enjoys nearly all of the monotonicity properties that are requested in a valid quantum resource theory. The relative entropy is defined as
$S(\rho_{1}|\rho_{2})={\rm Tr}\rho_{1}\log_{2}\rho_{1}-{\rm Tr}\rho_{1}\log_{2}\rho_{2}.$ (13)
For states that are diagonal in the same basis with eigenvalues $e^{1}_{i}$ and $e^{2}_{i}$ respectively we have
$S(\rho_{1}|\rho_{2})=\sum_{i}e^{1}_{i}\log_{2}e^{1}_{i}-\sum_{i}e^{1}_{i}\log_{2}e^{2}_{i}.$ (14)
Hence, we define the measure of Bell nonlocality based on the relative entropy as
$\mathcal{M}_{Re}(\rho)=\min_{\rho^{loc}\in\sigma}S(\rho|\rho^{loc}).$ (15)
We will show here that the measure defined by Eq. (15) fulfills monotonicity under subselection, i.e.:
$\mathcal{M}_{Re}(\rho)\geq\displaystyle\sum_{i}p_{i}\mathcal{M}_{Re}(\rho_{i}),$ (16)
where $\{\Lambda{i}\}$ is any set of Kraus operators, with weights $p_{i}={\rm Tr}(\Lambda_{i}\rho\Lambda_{i}^{\dagger})$, under the hypothesis that the involved Kraus operators are free, i.e. preserve the local states: $\sum_{i}\Lambda_{i}^{\dagger}\Lambda_{i}=\mathbb{I}$ and $\Lambda_{i}\rho\Lambda_{i}^{\dagger}\in\sigma$ $\forall\rho^{loc}\in\sigma$, where $\sigma$ is the set of local states.
To obtain the last inequality, first of all note that since the relative entropy is invariant under the same unitary or isometry applied to both arguments, we can use the isometry $I=\sum_{i}|i\rangle_{aux}\otimes|i\rangle_{aux^{\prime}}\otimes V_{i}$ for any $V_{i}$ to derive the identity $S(\rho^{(1)}|\rho^{(2)})=S(I\rho^{(1)}I^{\dagger}|I\rho^{(2}I^{\dagger})$, for any couple of states $\rho^{(1)}$ and $\rho^{(2)}$, where $\{|i\rangle_{aux}\}_{i}$ is an orthogonal basis of an auxiliary system. Next, monotonicity of the relative entropy under the partial trace over the auxiliary system $aux$ implies
$S(\rho^{(1)}|\rho^{(2)})\geq\sum_{i}S(\tilde{\rho}_{i}^{(1)}|\tilde{\rho}_{i}^{(2)}),$ (17)
where we use unnormalized states $\tilde{\rho}_{i}^{(1)}=V_{i}\rho^{(1)}V_{i}^{\dagger}$ and $\tilde{\rho}_{i}^{(2)}=V_{i}\rho^{(2)}V_{i}^{\dagger}$. Following the argument derived in Vedral1998 we have
$\sum_{i}S(\rho_{i}^{(1)}|\rho_{i}^{(2)})\geq\sum_{i}p_{i}S\left(\frac{\tilde{\rho}_{i}^{(1)}}{p_{i}}\bigg{|}\frac{\tilde{\rho}_{i}^{(2)}}{q_{i}}\right),$ (18)
where $p_{i}={\rm Tr}\ \tilde{\rho}_{i}^{(1)}$ and $q_{i}={\rm Tr}\ \tilde{\rho}_{i}^{(2)}$. Identifying $\rho\equiv\rho^{(1)}$, assuming that $\rho^{loc}\equiv\rho^{(2)}$ is a local sate and $V_{i}$ are Kraus operators which preserve locality, such that $V_{i}\rho^{loc}V_{i}^{\dagger}/q_{i}$ is still a local state, we conclude that
$\mathcal{M}_{Re}(\rho)\geq\sum_{i}p_{i}\mathcal{M}_{Re}\left(\frac{V_{i}\rho V_{i}^{\dagger}}{p_{i}}\right).$ (19)
IV.1 Distances and induced measures of nonlocality
Here we discuss the properties of the geometric based measures of Bell nonlocality as defined, respectively, by Eq. (5) and Eq. (15). Concerning the geometric measures, we will consider those associated to the Hilbert-Schmidt norm and to some of the most relevant contractive distances. In general, each distance provides a slightly different clue on the distinguishability of quantum states.
IV.1.1 Hilbert-Schmidt distance
The Hilbert-Schmidt distance $D_{HS}(\rho_{1},\rho_{2})$ between two density operators $\rho_{1}$ and $\rho_{2}$ is defined trough the Hilbert-Schmidt norm as follows:
$D_{HS}(\rho_{1},\rho_{2})\,=\,\|\rho_{1}-\rho_{2}\|_{HS}\,=\,\sqrt{{\rm Tr}\left[(\rho_{1}-\rho_{2})(\rho_{1}-\rho_{2})\right]}\,\,.$ (20)
For convenience, we define the corresponding geometric measure of Bell nonlocality $\mathcal{M}_{HS}(\rho)$ through the squared distance:
$\mathcal{M}_{HS}(\rho)\,=\,\min_{\rho^{loc}\in\sigma}D_{HS}^{2}(\rho,\rho^{loc})\,.$ (21)
The Hilbert-Schmidt (HS) distance has no direct operational meaning and in general may not monotonic under all CPTP transformations Bengtsson2017 ; Piani2012 . For instance, it increases under a trivial extension of local subsystems. On the other hand, it provides easily computable bounds on valid contractive distances that are often more difficult to compute.
IV.1.2 Hellinger distance
The Hellinger distance is defined as the Euclidean two-norm of the difference of the square roots of the density matrices (taking squares for computational convenience):
$D_{He}(\rho_{1},\rho_{2})=\|\sqrt{\rho_{1}}-\sqrt{\rho_{2}}\|_{2}.$ (22)
Given a pair of states that are diagonal in the same basis with, respectively, eigenvalues $\{e^{1}_{i}\}$ and $\{e^{2}_{i}\}$, one has:
$D_{He}(\rho_{1},\rho_{2})=\sqrt{2-2\sum_{i}\sqrt{e^{(1)}_{i}e^{(2)}_{i}}}=\sqrt{\displaystyle\sum_{i=1}^{4}\bigg{(}\sqrt{e_{i}^{(1)}}-\sqrt{e_{i}^{(2)}}\bigg{)}^{2}}.$ (23)
The above is the Euclidean norm in the $\sqrt{e_{i}}-$space and coincides with the classical Hellinger distance quantifying the distance between the two classical probability distributions Ref_Hellinger1909 . The corresponding Hellinger measure of Bell nonlocality is thus defined as
$\mathcal{M}_{He}(\rho)=\min_{\rho^{loc}\in\sigma}D_{He}^{2}(\rho,\rho^{loc})\,.$ (24)
IV.1.3 Bures distance
The Bures distance is defined as Bengtsson2017
$D_{B}=\sqrt{2\big{[}1-\sqrt{F(\rho_{1},\rho_{2}})\big{]}}\,,$ (25)
where $F$ is the Uhlmann fidelity given by Uhlmann1976 ; Jozsa1994 ; Nielsen2005
$F(\rho_{1},\rho_{2})=\bigg{(}\text{Tr}\bigg{[}\sqrt{\sqrt{\rho_{1}}\rho_{2}\sqrt{\rho_{1}}}\bigg{]}\bigg{)}^{2}\,.$ (26)
In the case of commuting density operators, the Bures and Hellinger distances coincide. The corresponding Bures measure of Bell nonlocality reads
$\mathcal{M}_{B}(\rho)=\min_{\rho^{loc}\in\sigma}D_{B}(\rho,\rho^{loc})\,.$ (27)
IV.1.4 Trace distance
The trace distance is defined as
$D_{Tr}(\rho_{1},\rho_{2})=\frac{1}{2}\|\rho_{1}-\rho_{2}\|_{1}=\frac{1}{2}{\rm Tr}\left[\sqrt{(\rho_{1}-\rho_{2})(\rho_{1}-\rho_{2})}\right]\,,$ (28)
where $\|\mathcal{O}\|_{1}\equiv\sqrt{\mathcal{O}^{\dagger}\mathcal{O}}$. For states that are diagonal in the same basis with eigenvalues $\{e^{1}_{i}\}$ and $\{e^{2}_{i}\}$ respectively, we have
$D_{Tr}(\rho_{1},\rho_{2})=\frac{1}{2}{\rm Tr}|\rho_{1}-\rho_{2}|=\frac{1}{2}\sum_{i}|e_{i}^{1}-e^{2}_{i}|\,.$ (29)
The corresponding trace-distance measure of Bell nonlocality reads
$\mathcal{M}_{Tr}(\rho)=\min_{\rho^{loc}\in\sigma}D_{Tr}(\rho,\rho^{loc}).$ (30)
The trace distance defines the probability of error in the Helstrom measurement which is the optimal measurement to distinguish two quantum states. It is jointly convex and nonincreasing under CPTP maps Bengtsson2017 ; Nielsen2005 . As a consequence, the corresponding measure of Bell nonlocality Eq. (30) satisfies monotonicity under LOSR and monotonicity under mixing (convexity).
V Application to two-qubit states: Formalism and CHSH locality
In this section we specialize to two-qubit states in relation to the CHSH inequality Eq. (3). Before proceeding, we recollect some basic facts and derive some theorems that are necessary in order to apply the general framework outlined in the previous sections to the concrete instance of two-qubit states. An arbitrary two-qubit state $\rho$ can be represented as
$\rho=\frac{1}{4}\sum_{i,j=0}^{3}\alpha_{ij}\sigma_{i}\otimes\sigma_{j},$ (31)
where $\sigma_{0}=\mathbf{1}$ and the $\sigma_{i}$ ($i=1,2,3$) are the Pauli matrices satisfying ${\rm Tr}\sigma_{i}=0$, ${\rm Tr}\sigma_{i}\sigma_{j}=\delta_{ij}$, and the superposition coefficients read $\alpha_{00}=1$, $\alpha_{0i}=\text{Tr}\big{[}\rho\big{(}\mathbb{I}\otimes\sigma_{i}\big{)}\big{]}$, $\alpha_{i0}=\text{Tr}\big{[}\rho\big{(}\sigma_{i}\otimes\mathbb{I}\big{)}\big{]}$ and $\alpha_{ij}=\text{Tr}\big{[}\rho\big{(}\sigma_{i}\otimes\sigma_{j}\big{)}\big{]}$ for $i=1,2,3$; they can be recast in matrix form as follows:
$\alpha=\begin{bmatrix}1&\vec{\alpha}_{0i}\\ \vec{\alpha}_{i0}&A\end{bmatrix},$ (32)
where all entries are real, $\vec{\alpha}_{0i}$ is a row three-vector of elements $\{\alpha_{0i}\}$, $\vec{\alpha}_{i0}$ is a column three-vector of elements $\{\alpha_{i0}\}$, and $A$ is a $3\times 3$ real matrix. Explicitly:
$\rho=\frac{1}{4}\bigg{(}\mathbf{1}+\sum_{i=1}^{3}\alpha_{0i}\mathbf{1}\otimes\sigma_{i}+\sum_{i=1}^{3}\alpha_{i0}\sigma_{i}\otimes\mathbf{1}+\sum_{i,j=1}^{3}A_{ij}\sigma_{i}\otimes\sigma_{j}\bigg{)}\,.$ (33)
The condition $\text{Tr}[\rho^{2}]\leq 1$ leads to
$\sum_{i}(\alpha_{0i}^{2}+\alpha_{i0}^{2})+\displaystyle\sum_{i,j}\alpha_{ij}^{2}\leq 3\,.$ (34)
It is useful to consider the singular value decomposition of the matrix $A$:
$A=O^{1}DO^{2},$ (35)
where $O^{1}$ and $O^{2}$ are orthogonal matrices and $D$ is diagonal and non-negative. The orthogonal transformations correspond to local changes of basis:
$\rho=\frac{1}{4}\bigg{(}\mathbf{1}+\sum_{i=1}^{3}\alpha_{0i}\mathbf{1}\otimes\sigma_{i}+\sum_{i=1}^{3}\alpha_{i0}\sigma_{i}\otimes\mathbf{1}+\sum_{k=1}^{3}D_{k}\bigg{(}\sum_{i}O^{1}_{ik}\sigma_{i}\bigg{)}\otimes\bigg{(}\sum_{j}O^{2}_{kj}\sigma_{j}\bigg{)}\bigg{)}.$ (36)
V.1 Bell-diagonal states
Given the Bell basis formed by the four maximally entangled Bell states
[…]
their convex combinations define the Bell-diagonal states:
$\rho_{BD}=\displaystyle\sum_{i=1}^{4}e_{i}\ket{\Psi_{i}}\bra{\Psi_{i}}\,,$ (38)
where $\{0\leq e_{i}\leq 1\}$, with $\sum_{i=1}^{4}e_{i}=1$. By resorting to the notation of Eq. (33), a generic Bell-diagonal state Eq. (38) can be rewritten as
$\rho_{BD}=\frac{1}{4}\bigg{[}\mathcal{I}_{2}\otimes\mathcal{I}_{2}+\displaystyle\sum_{i=1}^{3}a_{i}\sigma_{i}\otimes\sigma_{i}\bigg{]}\,,$ (39)
where $\mathcal{I}_{2}$ denotes the identity matrix in two dimensions. Matrix $\alpha$ of Eq. (32) can then be recast in the form
$\alpha_{BD}=\begin{bmatrix}1&0&0&0\\ 0&a_{1}&0&0\\ 0&0&a_{2}&0\\ 0&0&0&a_{3}\end{bmatrix}\,.$ (40)
Figure 1: Semitransparent tetrahedron: set of Bell-diagonal states. The set is parameterized by the matrix elements $(a_{1},a_{2},a_{3})$ in Eq. (40). Solid figure inside the tetrahedron: set of local states. Solid grey line: set of Werner states.
Collecting the above results, for any Bell-diagonal state it is rather straightforward to derive the following formula relating the coefficients $\vec{a}=(1,a_{1},a_{2},a_{3})^{T}$ of the matrix $\alpha$ and the eigenvalues $\vec{e}=(e_{1},...,e_{4})^{T}$ of the density matrix:
$\begin{bmatrix}1\\ a_{1}\\ a_{2}\\ a_{3}\end{bmatrix}=\begin{bmatrix}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ -1&1&1&-1\end{bmatrix}\begin{bmatrix}e_{1}\\ e_{2}\\ e_{3}\\ e_{4}\end{bmatrix}\,.$ (41)
The matrix in the above expression is unitary upon multiplication by $1/2$. It follows that
$\begin{bmatrix}e_{1}\\ e_{2}\\ e_{3}\\ e_{4}\end{bmatrix}=\frac{1}{4}\begin{bmatrix}1&1&1&-1\\ 1&1&-1&1\\ 1&-1&1&1\\ 1&-1&-1&-1\end{bmatrix}\begin{bmatrix}1\\ a_{1}\\ a_{2}\\ a_{3}\end{bmatrix}\,,$ (42)
so that positivity and normalization of the sum of the eigenvalues allow to represent the set of all Bell-diagonal states parameterized by the three-dimensional vector $(a_{1},a_{2},a_{3})$ as shown in Fig. (1).
V.2 Werner states
A particular subclass of the Bell-diagonal states are the Werner states $\rho_{w}$, described by one free parameter $|a_{1}|=|a_{2}|=|a_{3}|\equiv w$. Formally, a Werner state is a convex combination of the maximally mixed state with any one of the four Bell states. This class of states, first introduced by R. F. Werner Werner1989 , is important as it provides the paramount example of a set of states that for $\frac{1}{3}\leq w\leq\frac{1}{\sqrt{2}}$ are entangled but nevertheless local.
V.3 CHSH-local states and their properties
Let $(d_{1},d_{2},d_{3})$ be the singular values in the decomposition Eq. (35) and, without loss of generality let $d_{1}\geq d_{2}\geq d_{3}$. As proved in Ref. Horodecki1995 , a state is CHSH-local, i.e. it does not violate the CHSH inequality Eq. (3), if and only if
$d_{1}^{2}+d_{2}^{2}\leq 1.$ (43)
We prove the following:
Lemma 1. If $\rho$ is a CHSH-local state with given $(\alpha_{0i},\alpha_{i0},A)$ then the state $\rho^{-}$ given by $(-\alpha_{0i},-\alpha_{i0},A)$ is also CHSH-local.
Proof. It suffices to consider a local transformation that for $i=1,2,3$ changes $\sigma_{i}$ into $-\sigma_{i}$.
Lemma 2. For any CHSH-local state $\rho$ with given $(\alpha_{0i},\alpha_{i0},A)$ there exists a CHSH-local state $\rho^{\prime}$ with the same $A$ but with $\alpha_{0i}=\alpha_{i0}=0$
Proof. Consider the convex combination
$\rho^{\prime}=\frac{1}{2}\rho+\frac{1}{2}\rho^{-}.$ (44)
Indeed, for such a combination we have
$\alpha=\begin{bmatrix}\alpha_{00}&0\\ 0&A\end{bmatrix}.$ (45)
Now, convex combinations of local states are local: by virtue of the property of shared randomness, if there exists a hidden variable model for each of the states in the combination, then there must exist a hidden variable model for their classical statistical mixture.
V.4 Properties of symmetric states
Proposition 1. For any jointly convex and unitarily invariant functional $D$ of pairs of two-qubit density matrices
$\min_{\rho^{loc}\in\sigma}D(\rho_{w},\rho^{loc})=\min_{\rho^{loc}_{w}\in\sigma}D(\rho_{w},\rho^{loc}_{w})\,,$ (46)
where $\rho_{w}$ is a Werner state, $\rho^{loc}$ is a CHSH-local state, and $\rho^{loc}_{w}$ is a CHSH-local Werner state.
Proof. The proof is based on the following observations:
  • Obs. 1. Invariance of Werner states under the same local unitary transformations on both subsystems:
    $\forall\ {\rm unitary}\ U:\qquad\left(U\otimes U\right)\rho_{w}\left(U^{\dagger}\otimes U^{\dagger}\right)=\rho_{w}.$ (47)
  • Obs. 2. Local unitary transformations are resource non-generating (RNG) for nonlocality; in particular, $\left(U\otimes U\right)\rho^{loc}\left(U^{\dagger}\otimes U^{\dagger}\right)$ remains local.
  • Obs. 3. Any convex combination of local states is a local state.
  • Obs. 4. Twirling always projects on Werner states. Therefore, we have that for any given class of density matrices $\rho^{a}$
[…]
    where the integral is taken over the Haar measure on the set of unitary transformations.
From the unitary invariance of $D$ we have
$D(\rho_{w},\rho^{loc})$ $=$ $D\left(U\otimes U\rho_{w}U^{\dagger}\otimes U^{\dagger}\,,\,U\otimes U\rho^{loc}U^{\dagger}\otimes U^{\dagger}\right)$ (49)
[…]
$=$ $D(\rho_{w},\rho^{loc}_{w})\,.$ (52)
Here the first and second lines come from unitary invariance of $D$; the third follows from the joint convexity; and the fourth from Obs. 1 - 4 above.
Similar propositions hold for general Bell-diagonal states; however, in the general case the proof holds for fairly smaller symmetry classes.
Proposition 2. For any jointly convex and unitarily invariant functional $D$ of pairs of two-qubit density matrices, the closest CHSH-local state to a Bell-diagonal state is Bell-diagonal:
$\min_{\rho_{loc}\in\sigma}f(\rho_{BD},\rho_{loc})=\min_{\rho_{BD,loc}\in\sigma}f(\rho_{BD},\rho_{BD,loc})\,,$ (53)
where $\rho_{BD}$ denotes a Bell-diagonal state, $\rho_{loc}$ a CHSH-local state, and $\rho_{BD,loc}$ a CHSH-local Bell-diagonal state.
Proof. The symmetry class that we consider contains the simultaneous local $\pi$-rotations of qubits around the three Pauli axes. Rotation $R_{1}$ around $\sigma_{1}$ changes $\sigma_{2}$ and $\sigma_{3}$ to $-\sigma_{2}$ and $-\sigma_{3}$ respectively. Rotations $R_{2}$ and $R_{3}$ operate analogously. These transformations are local and unitary, therefore they cannot change the locality of a quantum state. Moreover, these transformations preserve Bell-diagonal states. Indeed, let us consider the general state representation Eq. (31). Under rotation $R_{1}$ we have the following transformation:
$\alpha=\begin{pmatrix}1&a_{01}&a_{02}&a_{03}\\ a_{10}&a_{11}&a_{12}&a_{13}\\ a_{20}&a_{21}&a_{22}&a_{23}\\ a_{30}&a_{31}&a_{32}&a_{33}\end{pmatrix}\,\,\longrightarrow\,\,\alpha_{1}=\begin{pmatrix}1&a_{01}&-a_{02}&-a_{03}\\ a_{10}&a_{11}&-a_{12}&-a_{13}\\ -a_{20}&-a_{21}&a_{22}&a_{23}\\ -a_{30}&-a_{31}&a_{32}&a_{33}\end{pmatrix}.$ (54)
The average $\rho^{\prime}=1/2(\rho+R_{1}(\rho))$ enjoys the representation
$\alpha_{1}^{\prime}=\begin{pmatrix}1&a_{01}&0&0\\ a_{10}&a_{11}&0&0\\ 0&0&a_{22}&a_{23}\\ 0&0&a_{32}&a_{33}\end{pmatrix}.$ (55)
Applying rotation $R_{2}$ to $\rho^{\prime}$ and averaging the result with $\rho^{\prime}$ leads to the state with diagonal representation Eq. (39), i.e. a Bell-diagonal state. The total transformation that projects a given state to a Bell-diagonal state takes the compact form
$\rho\,\,\longrightarrow\,\,\frac{1}{2}\left(\frac{1}{2}(\rho+R_{1}(\rho))+R_{2}\left(\frac{1}{2}(\rho+R_{1}(\rho))\right)\right).$ (56)
This transformation obviously preserves the Bell-diagonal states, does not change the locality, and is a composition of joint local unitaries and averaging. The proof of Proposition 2 is thus immediately completed by straightforwardly adapting the reasoning followed in the proof of Proposition 1.
VI CHSH nonlocality of two-qubit states
Exploiting the results derived in the previous section, we can proceed to quantify the CHSH nonlocality of two-qubit states using the Hilbert space based distance and relative entropy measures with respect to the set of CHSH-local two-qubit states. Explicit examples will include Werner states, Bell-diagonal states, and reduced convex combinations of Bell states. These important classes of states are routinely considered in the study of quantum resource theories, starting from the pioneering investigations on entanglement and discord Wootters1997 ; Wootters1998 ; Henderson_Vedral2001 .
VI.1 Werner states
VI.1.1 Hilbert-Schmidt distance
Consider the squared Hilbert-Schmidt (HS) distance defined on a generic pair $\{\rho_{1},\rho_{2}\}$ of two-qubit states as induced by the HS norm Eq. (20). Using Eq. (31) and the properties of the Pauli matrices:
$\|\rho_{1}-\rho_{2}\|_{HS}\,=\,\|\alpha_{1}-\alpha_{2}\|_{HS}\,.$ (57)
It follows from Proposition 1, Eq. (46), that the closest CHSH-local state to a Werner state is also a Werner state with $w\leq 1/\sqrt{2}$. Hence, for CHSH-nonlocal Werner states, i.e. with $w\geq 1/\sqrt{2}$ :
$\mathcal{M}_{HS}(\rho_{w})=\min_{\rho^{loc}_{w}\in\sigma}\left\|\rho_{w}-\rho^{loc}\right\|_{HS}^{2}=\left\|\frac{1}{\sqrt{2}}\mathbf{1}-w\mathbf{1}\right\|_{HS}^{2}=3\left(\frac{1}{\sqrt{2}}-w\right)^{2}\,,$ (58)
where $\mathbf{1}$ is the identity matrix and $w\in\left(1/\sqrt{2},1\right]$. Correctly, the maximum is achieved by the maximally entangled Bell states corresponding to $w=1$ :
$\mathcal{M}_{HS}^{max}\,=\,\frac{3}{2}\left(3-2\sqrt{2}\right)\,.$ (59)
The behavior of the normalized HS geometric measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{HS}=\mathcal{M}_{HS}(\rho_{w})/\mathcal{M}_{HS}^{max}$ as a function of the $w$ parameter is reported in Fig. 2. The plot is a parabola, as the the HS distance is simply the Euclidean distance in the three-dimensional space of Fig. 1.
VI.1.2 Hellinger distance
Recalling again the consequences of Proposition 1, Eq. (46), the measure of CHSH nonlocality for Werner states based on the (squared) Hellinger distance, Eq. (23), reads
$\mathcal{M}_{He}(\rho_{w})=\min_{\rho_{w}^{loc}\in\sigma}D^{2}_{He}(\rho_{w},\rho_{w}^{loc})=2-\frac{1}{2}\left(3\sqrt{(1-w)\left(1-\frac{1}{\sqrt{2}}\right)}+\sqrt{(1+3w)\left(1+\frac{3}{\sqrt{2}}\right)}\right)\,,$ (60)
where $w\in(1/\sqrt{2},1]$. Here we have used the relation between the eigenvalues of the state density matrix and the Werner parameter $w$ which is derived directly from Eq. (42). As $|a_{1}|=|a_{2}|+|a_{3}|=w$, the signs of the matrix elements $a_{i}$ identify the states corresponding to different corners of the state tetrahedron in Fig. 1. The Hellinger measure is maximized by the Bell states at $w=1$ :
$\mathcal{M}_{He}^{max}=2-\sqrt{1+\frac{3}{\sqrt{2}}}\,\,.$ (61)
The behavior of the normalized Hellinger geometric measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{He}=\mathcal{M}_{He}(\rho_{w})/\mathcal{M}_{He}^{max}$ as a function of the $w$ parameter is shown in Fig. 2.
VI.1.3 Bures distance
In the case of commuting operators the Bures distance and the Hellinger distance coincide. This is indeed the case for Werner density matrices. Therefore, the Bures and Hellinger measures of CHSH nonlocality coincide on Werner states $\rho_{w}$:
$\mathcal{M}_{Bu}(\rho_{w})=\mathcal{M}_{He}(\rho_{w})\,.$ (62)
VI.1.4 Trace distance
The geometric measure of CHSH nonlocality of Werner states based on the trace distance Eq. (28) reads:
$\mathcal{M}_{Tr}(\rho_{w})=\min_{\rho_{w}^{loc}\in\sigma}D_{Tr}(\rho_{w},\rho_{w}^{loc})=\frac{3}{2}\left|w-\frac{1}{\sqrt{2}}\right|=\frac{3}{2}\left(w-\frac{1}{\sqrt{2}}\right)\,,$ (63)
where the last equality follows from the fact that without loss of generality we can limit the analysis to the interval $w\in(1/\sqrt{2},1]$. The maximum value is achieved at $w=1$ by the Bell states:
$\mathcal{M}_{Tr}^{max}=\frac{3}{2}\left(1-\frac{1}{\sqrt{2}}\right)\,.$ (64)
The normalized trace measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{Tr}=\mathcal{M}_{Tr}(\rho_{w})/\mathcal{M}_{Tr}^{max}$ as a function of the $w$ parameter is shown in figure 2.
VI.1.5 Relative entropy
It is straightforward to verify that for a nonlocal Werner state in the interval $w\geq 1/\sqrt{2}\,$ its relative entropy Eq. (14) with respect to the set $\sigma$ of local states is minimized by the local state with $w=1/\sqrt{2}$ , so that:
$\mathcal{M}_{Re}(\rho_{w})=\min_{\rho_{w}^{loc}\in\sigma}S(\rho_{w}|\rho_{w}^{loc})$ $=$ $\frac{3(1-w)}{4}\log_{2}\left(\frac{3(1-w)}{4}\right)+\frac{1+3w}{4}\log_{2}\left(\frac{1+3w}{4}\right)+$ (65)
-- $\frac{3(1-w)}{4}\log_{2}\left(\frac{\sqrt{2}-1}{4\sqrt{2}}\right)-\frac{1+3w}{4}\log_{2}\left(\frac{\sqrt{2}+3}{4\sqrt{2}}\right)\,.$
The maximum value achieved by the Bell states at $w=1$ reads
$\mathcal{M}_{Re}^{max}=2-\log_{2}{\bigg{(}1+\frac{3}{\sqrt{2}}\bigg{)}}\,.$ (66)
The normalized relative entropy measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{Re}=\mathcal{M}_{Re}(\rho_{w})/\mathcal{M}_{Re}^{max}$ as a function of the $w$ parameter is reported in Fig. 2.
Figure 2: Normalized geometric measures of CHSH nonlocality for Werner states as a function of the Werner parameter $w$ in the interval $\frac{1}{\sqrt{2}}\leq w\leq 1$. Black dashed straight line: Trace measure ~$\widetilde{\mathcal{M}}_{Tr}$. Grey dashed curve: Hilbert-Schmidt measure ~$\widetilde{\mathcal{M}}_{HS}$. Black solid curve: Relative entropy measure ~$\widetilde{\mathcal{M}}_{Re}$. Grey solid curve: Hellinger measure ~$\widetilde{\mathcal{M}}_{He}$. All plotted quantities are dimensionless.
VI.2 Bell-diagonal states
VI.2.1 Hilbert-Schmidt distance
All Bell-diagonal states can be represented as points in a tetrahedron, as shown in Fig. 1, with vertices representing Bell states. The points are given by three dimensional vectors $\vec{\alpha}$ that appear in Eq. (57). This equation implies that the Euclidean distance in the space represented in Fig. 1 coincides with the HS distance in the space of density matrices of Bell-diagonal states. Moreover, for all Bell-diagonal states the closest local state is a Bell-diagonal state, as proved in Proposition 2, Eq. (53). The set of local states is represented by points in the solid figure inscribed inside the tetrahedron, see Fig. 1. Hence, the Euclidean distance between the points representing nonlocal Bell-diagonal states and the boundary of the solid figure yields the HS-based geometric measure of CHSH nonlocality. Exploiting the Horodecki locality criterion Eq. (43), the separation surface between local and nonlocal Bell-diagonal state is identified by the solutions of the equation
$\max_{i,ji\neq j}\{a_{i}^{2}+a_{j}^{2}\}=1\,.$ (67)
Combining the above with Proposition 2, Eq. (53), the minimum of the squared HS distance is determined as
$\min_{\rho_{BD}^{loc}\in\sigma}\left\|\rho_{BD}-\rho_{BD}^{loc}\right\|_{HS}^{2}=\min_{a_{i}^{loc}}\frac{1}{2}\displaystyle\sum_{i=1}^{3}(a_{i}-a_{i}^{loc})^{2}\,,$ (68)
with the provision that the states $\rho_{BD}$ belong to the surface Eq. (67).
Considering first the case $|a_{1}^{loc}|,|a_{2}^{loc}|>|a_{3}^{loc}|$, we can proceed to determine the minimum by introducing a Lagrange multiplier $\lambda$. The associated Lagrange variational function reads
$\Gamma(a_{1}^{loc},a_{2}^{loc},a_{3}^{loc},\lambda)\equiv\displaystyle\sum_{i=1}^{3}(a_{i}-a_{i}^{loc})^{2}+\lambda[(a_{1}^{loc})^{2}+(a_{2}^{loc})^{2}-1]\,.$ (69)
Vanishing of the gradient of $\Gamma$ leads to
$\begin{cases}(a_{1}-a_{1}^{loc})+\lambda a_{1}^{loc}=0,\\ (a_{2}-a_{2}^{loc})+\lambda a_{2}^{loc}=0,\\ (a_{3}-a_{3}^{loc})=0,\\ (a_{1}^{loc})^{2}+(a_{2}^{loc})^{2}-1=0.\end{cases}$
Denoting the minima by $\tilde{a}_{i}^{loc}$, the solution reads
$\tilde{a}_{1}^{loc}=\frac{-a_{2}\pm\sqrt{a_{2}^{2}+4}}{2},$
$\tilde{a}_{2}^{loc}=\frac{-a_{1}\pm\sqrt{a_{1}^{2}+4}}{2},$
$\tilde{a}_{3}^{loc}=a_{3}.$ (70)
By Eq. (68) the (un-normalized) HS geometric measure of CHSH nonlocality for Bell-diagonal states is thus
$\mathcal{M}_{HS}(\rho_{BD})=\frac{1}{4}\left[(a_{1}-a_{2})^{2}+\frac{a_{1}^{2}+a_{2}^{2}}{2}+(a_{2}-a_{1})\bigg{(}\sqrt{a_{1}^{2}+4}-\sqrt{a_{2}^{2}+4}\bigg{)}+2\right].$ (71)
The remaining cases $|a_{2}^{loc}|,|a_{3}^{loc}|>|a_{1}^{loc}|$ and $|a_{1}^{loc}|,|a_{3}^{loc}|>|a_{2}^{loc}|$ can be immediately found by symmetry and the global minimum is determined over all possible cases. For the maximally nonlocal Bell states the minimum is attained when $|a_{1}^{loc}|=|a_{2}^{loc}|=|a_{3}^{loc}|\equiv w^{2}$, with $w$ the Werner parameter. The constraints in Eq. (69) thus reduce to $12w^{2}=$ 1, i.e. $w=1/\sqrt{2}$, and $\mathcal{M}_{HS}$ achieves its maximum $\mathcal{M}_{HS}^{max}$, Eq. (59). Using Eqs. (41) and (42) the transformation from the matrix elements $\{a_{i}\}$ to the eigenvalues $\{e_{i}\}$ of the density matrix follows immediately. The behavior of the normalized HS measure of CHSH nonlocality ~$\widetilde{\mathcal{M}}_{HS}=\mathcal{M}_{HS}/\mathcal{M}_{HS}^{max}$ for Bell-diagonal states as a function of the eigenvalues $e_{1}$, $e_{2}$ in the case of $e_{4}=0$ is reported in Fig. 3.
Figure 3: Normalized Hilbert-Schmidt measure ~$\widetilde{\mathcal{M}}_{HS}$ of CHSH nonlocality for Bell-diagonal states as a function of the two free parameters $e_{1}$ and $e_{2}$, with $e_{4}=0$, $e_{3}=1-e_{1}-e_{2}$, and normalization coefficient $\mathcal{M}_{HS}^{max}$. The maximum ~$\widetilde{\mathcal{M}}_{HS}=1$ is reached in the three cases $(e_{1},e_{2},e_{3})=(1,0,0)$; $(e_{1},e_{2},e_{3})=(0,1,0)$; and $(e_{1},e_{2},e_{3})=(0,0,1)$. Each case corresponds to a different Bell state. The central region of minima is composed by local states, i.e. states with ~$\widetilde{\mathcal{M}}_{HS}=0$. All plotted quantities are dimensionless.
VI.2.2 Hellinger distance
Each of the eigenvalues $\{e_{i}^{k}\}$ of a density matrix $\rho_{k}$ is the probability that a system in state $\rho_{k}$ collapses in the eigenstate relative to that eigenvalue. For a Bell-diagonal state $\rho_{BD}$ the eigenstates are obviously the four Bell states. Recalling Proposition 2, Eq. (53), and Eq. (23) the minimization process on the Hellinger distance squared yields the Hellinger geometric measure of CHSH nonlocality $\mathcal{M}_{He}(\rho_{BD})$ for Bell-diagonal states:
$\mathcal{M}_{He}(\rho_{BD})=\min_{\rho_{BD}^{loc}\in\sigma}D^{2}_{He}(\rho_{BD},\rho_{BD}^{loc})=\min_{\rho^{loc}\in\sigma}\left\{\displaystyle\sum_{i=1}^{4}\bigg{(}\sqrt{e_{i}}-\sqrt{e_{i}^{loc}}\bigg{)}^{2}\right\}\,.$ (72)
Using Eq.(42) we can rewrite the condition Eq. (67) in terms of the eigenvalues $\{e_{i}^{loc}\}$. Fixing, for instance, the case $|a_{1}|,|a_{2}|>|a_{3}|$, one has:
$(e_{1}-e_{3})^{2}+(e_{2}-e_{4})^{2}\,>\,\frac{1}{2}\,.$ (73)
For probabilities, their positivity $e_{i}>0$ and normalization condition $\sum_{i}e_{i}=1$ define a clove of a unitary $4-$sphere in the $\sqrt{e_{i}}-$space. Therefore, the set of nonlocal states with $|a_{1}|,|a_{2}|>|a_{3}|$ is represented in the $\sqrt{e_{i}}-$space by the intersection of the hypervolume Eq. (73) with the unitary semi-sphere centered in the origin. Since the distance in this space is the Euclidean one, the minimum in Eq. (72) in the case $|a_{1}^{loc}|,|a_{2}^{loc}|>|a_{3}^{loc}|$ is determined by imposing the constraint
$(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}\,.$ (74)
Recalling Eq. (23), the above relation identifying the region of separation between local and nonlocal states is obtained when $\sum_{i=1}^{4}\sqrt{e_{i}e_{i}^{loc}}$ is maximum. We can thus proceed by the method of Lagrange multipliers and introduce the following Lagrangian function
$\Gamma(e_{i}^{loc},\lambda_{1},\lambda_{2})\equiv\displaystyle\sum_{i=1}^{4}\sqrt{e_{i}e_{i}^{loc}}+\lambda_{1}\bigg{[}(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}-\frac{1}{2}\bigg{]}+\lambda_{2}\left(\displaystyle\sum_{i=1}^{4}e_{i}^{loc}-1\right)\,.$ (75)
Taking the gradient of $\Gamma$ and imposing its vanishing yields
$\begin{cases}\frac{1}{2}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}+2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0\\ \vspace{0.2cm}\par\frac{1}{2}\sqrt{\frac{e_{2}}{e_{2}^{loc}}}+2\lambda_{1}(e_{2}^{loc}-p_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{1}{2}\sqrt{\frac{e_{3}}{e_{3}^{loc}}}-2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\par\frac{1}{2}\sqrt{\frac{e_{4}}{e_{4}^{loc}}}-2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\ \vspace{0.2cm}\par\par(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}\,.\end{cases}$
From the above relations one immediately determines the Lagrange multipliers:
$\lambda_{1}=-\frac{\sqrt{\frac{e_{1}}{e_{1}^{loc}}}-\sqrt{\frac{e_{3}}{e_{3}^{loc}}}}{8(e_{1}^{loc}-e_{3}^{loc})}=-\frac{\sqrt{\frac{e_{2}}{e_{2}^{loc}}}-\sqrt{\frac{e_{4}}{e_{4}^{loc}}}}{8(e_{2}^{loc}-e_{4}^{loc})}\,,$ (76)
$\lambda_{2}=-\frac{1}{4}\bigg{(}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}+\sqrt{\frac{e_{3}}{e_{3}^{loc}}}\bigg{)}=-\frac{1}{4}\bigg{(}\sqrt{\frac{e_{2}}{e_{2}^{loc}}}+\sqrt{\frac{e_{4}}{e_{4}^{loc}}}\bigg{)}\,,$ (77)
so that
$\begin{cases}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}+\sqrt{\frac{e_{3}}{e_{3}^{loc}}}=\sqrt{\frac{e_{2}}{e_{2}^{loc}}}+\sqrt{\frac{e_{4}}{e_{4}^{loc}}}\,,\\ \vspace{0.2cm}\par\par\bigg{(}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}-\sqrt{\frac{e_{3}}{e_{3}^{loc}}}\bigg{)}(e_{2}^{loc}-e_{4}^{loc})=\bigg{(}\sqrt{\frac{e_{2}}{e_{2}^{loc}}}-\sqrt{\frac{e_{4}}{e_{4}^{loc}}}\bigg{)}(e_{1}^{loc}-e_{3}^{loc})\,,\\ \vspace{0.2cm}\par\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1\,,\\ \vspace{0.2cm}\par\par(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}\,.\end{cases}$
The remaining cases can be treated symmetrically; the absolute minimum is determined by the smallest value with respect to all the possible cases. In the case of Bell states, e.g. $e_{1}=1$, $e_{2}=e_{3}=e_{4}=0$, it is straightforward to verify that the minimum is reached for $e_{1}^{loc}=\frac{1+3t}{4}$, $e_{2}^{loc}=e_{3}^{loc}=e_{4}^{loc}=\frac{1-t}{4}$ , with $t=\frac{1}{\sqrt{2}}$. Therefore, as it must be, we find that the maximum value $\mathcal{M}_{He}^{max}$ of the Hellinger geometric measure of CHSH nonlocality Eq. (72) coincides with that of the Hellinger measure for Werner states, Eq. (60), with Werner parameter $w=1$.
On the other hand, we see that the above minimization procedure cannot always yield a closed formula of the Hellinger measure of CHSH nonlocality for generic Bell-diagonal states. This is implied by the type of bound on the set of local states and the form of the function one needs to minimize. Indeed, the above algebraic systems yields that in order to find the minimum one must determine the roots of a polynomial of the fifth order. Therefore, from the Abel-Ruffini-Seralian theorem it follows that it is impossible to obtain a closed analytical formula by quadratures for the roots in terms of the coefficients of the polynomial. Of course, the solutions can be determined to any desired degree of precision by standard approximation procedures such as, e.g., the Newton-Raphson method or the Laguerre iterative scheme. Alternatively, we can proceed analytically in special but relevant instances, such as the Bell states or, as we will show later on, the Bell-diagonal states with one free parameter.
VI.2.3 Bures distance
In analogy with the case of Werner states, also in the case of Bell-diagonal states the density operators commute and the Bures and Hellinger distances coincide, so that
$\mathcal{M}_{Bu}(\rho_{BD})=\mathcal{M}_{He}(\rho_{BD})\,.$ (78)
VI.2.4 Relative entropy
In the case of the relative entropy we assume that in order to determine the minimum in Eq. (14) we can continue to impose the constraint Eq. (74). In other words, we expect that also for the relative entropy of nonlocality of Bell-diagonal states the closest local state belongs to the same surface as for the Hellinger and Bures measures. Assuming the above hypothesis, the minimum in Eq. (14) is realized when
$\sum_{i=1}^{3}e_{i}\log_{2}{e_{i}^{loc}}$ (79)
achieves its maximum.
Fixing as usual, to begin with, the case $|a_{1}|,|a_{2}|>|a_{3}|$, the Lagrange function associated to the Lagrange multipliers takes the form
$\Gamma(e_{i}^{loc},\lambda_{1},\lambda_{2})\equiv\displaystyle\sum_{i=1}^{4}e_{i}\log_{2}{e_{i}^{loc}}+\lambda_{1}[(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}-\frac{1}{2}]+\lambda_{2}\left(\displaystyle\sum_{i=1}^{4}e_{ij}^{loc}-1\right)\,,$ (80)
which leads to
$\begin{cases}\frac{e_{1}}{e_{1}^{loc}\ln{2}}+2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{p_{01}}{e_{2}^{loc}\ln{2}}+2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{e_{3}}{e_{3}^{loc}\ln{2}}-2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{e_{4}}{e_{4}^{loc}\ln{2}}-2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\ \vspace{0.2cm}\par(e_{1}^{loc}-e_{2}^{loc})^{2}+(e_{3}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\end{cases}$
It is straightforward to show that the above relations reduce to
$\begin{cases}\frac{e_{1}}{e_{1}^{loc}}+\frac{e_{3}}{e_{3}^{loc}}=\frac{e_{2}}{e_{2}^{loc}}+\frac{e_{4}}{e_{4}^{loc}},\\ \vspace{0.2cm}\par\bigg{(}\frac{e_{3}}{e_{3}^{loc}}-\frac{e_{1}}{e_{1}^{loc}}\bigg{)}(e_{2}^{loc}-e_{4}^{loc})=\bigg{(}\frac{e_{4}}{e_{4}^{loc}}-\frac{e_{2}}{e_{2}^{loc}}\bigg{)}(e_{1}^{loc}-e_{3}^{loc}),\\ \vspace{0.2cm}\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\ \vspace{0.2cm}\par(e_{1}^{loc}-e_{2}^{loc})^{2}+(e_{3}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\par\end{cases}$
In analogy with the case of the Hellinger measure, the above algebraic system is of higher order and not amenable to a solution by elementary quadratures. Therefore, in the following subsection we will consider the particular instance of one-parameter Bell-diagonal states that allows for the full analytical evaluation of the entire set of geometric and entropic measures of CHSH nonlocality.
VI.3 Bell-diagonal states with one free parameter
Let us consider Bell-diagonal states that are convex combinations of only two of the four Bell states. For instance, let us choose $e_{2}=e_{4}=0$ , $e_{1}=p$ , and $e_{3}=1-p$ , with $p$ free parameter in the interval $[0,1]$. Such Bell-diagonal states represent all possible convex combinations of the first and third Bell states of Eq. (37). Moving to symplectic space and recalling Eqs. (42), this instance corresponds to $a_{1}=a$, $a_{2}=-a$, and $a_{3}=1$, with $a=2p-1$.
For this class of states the locality condition Eq. (43) is fulfilled only for $p=\frac{1}{2}$; therefore, such states are nonlocal for $0\leq p\leq 1$, with the exception of $p=\frac{1}{2}$. The behavior of the four normalized measures of CHSH nonlocality as functions of the convex combination parameter $p$ in the half-interval $p\geq\frac{1}{2}$ is reported in Fig. 4. We see that all measures vanish at $p=\frac{1}{2}$ and reach their maximum at $p=1$, i.e. on the first Bell state. For all measures, the normalization coefficients turn out to be the same as the ones of the nonlocality measures for Werner states, in agreement with Proposition 1, Eq. (46), and Proposition 2, Eq. (53). The case $p\leq\frac{1}{2}$ is readily obtained by symmetry.
It is straightforward to verify that the above-described behavior is generic, i.e. it does not depend on the pair of Bell states that are chosen as entries of the convex combination.
Figure 4: Normalized measures of CHSH nonlocality for the convex combination of two Bell states as functions of the convex combination parameter $p$ in the interval $p\geq\frac{1}{2}$. The same behavior holds in the symmetric case $p\leq\frac{1}{2}$, with the maximum achieved in $p=0$. Black dashed curve: Trace measure ~$\widetilde{\mathcal{M}}_{Tr}$. Black solid curve: Relative entropy measure ~$\widetilde{\mathcal{M}}_{Re}$. Grey dashed curve: Hellinger measure ~$\widetilde{\mathcal{M}}_{He}$. Grey solid curve: Hilbert-Schmidt measure ~$\widetilde{\mathcal{M}}_{HS}$. All plotted quantities are dimensionless.
VII Discussion and outlook
We introduced a geometric measure for Bell nonlocality in a fixed scenario, whose defining elements are the state vectors in Hilbert space and their distances from the set of local states. In the case of the CHSH inequality, we have identified the basic traits of such approach, including the set of local states and the free operations, within the framework based on Local Operations and Shared Randomness (LOSR), and we have discussed the concept of maximally nonlocal states Buscemi2012 ; Wolfe2020 ; Schmid2020b ; Rosset2020 ; Lipka2021 ; Gallego2017 .
The quantification of Bell nonlocality for quantum states in Hilbert space is naturally introduced in terms of geometric measures, in close methodological analogy with the standard theory of geometric entanglement Wei2003 ; Uyanik2010 and the more recent geometric approaches to the quantification of other types of nonlocal correlations such as, e.g., the quantum discord Geomdiscord2016 ; Geomdiscord2017 . Geometric measures define a structure and an ordering in the space of quantum states, and we have provided a detailed investigation of such structure in the case of the CHSH inequality for different classes of contractive distances (trace, Hellinger, Bures), in the case of two-qubit states, by proving and exploiting the result that for contractive distances the closest local state to a Werner state is still a Werner state, and that an analogous result holds for Bell-diagonal states. We have confronted the ordering provided by contractive geometric measures of CHSH nonlocality with the non-contractive Hilbert-Schmidt distance, and we have introduced the relative entropy of nonlocality as a further benchmark.
An important aspect of our approach is that, within a fixed Bell scenario (defined by the number of measurements, outcomes, and a Bell inequality), we perform a maximization over all possible measurement settings compatible with the given constraints. This procedure ensures that geometric measures of nonlocality depend only on the quantum state and the structure of the chosen Bell scenario, and not on the particular experimental setting. On the other hand, one could envisage an alternative strategy, in which also the measurement settings are fixed from the outset. In this case, the set of local states would be defined relative to these fixed measurements, and the resulting geometric measure would capture the nonlocality of a state specifically with respect to the selected experimental configuration, without any apriori extremization. Such a fixed-measurements approach might offer advantages in operational contexts, where the measurements are experimentally constrained or predetermined.
Generalizations and applications of the conceptual framework introduced in the present work are conceivable along different directions, such as the identification of the set of local states and the ordering of nonlocal states in finite-dimensional Hilbert spaces of higher dimension and generalized Bell inequalities, the extension to field-theoretical settings of elementary particle physics, and the use of the Hilbert space-based measures of Bell nonlocality as tools in the investigation and characterization of quantum collective phenomena and quantum phase transitions. Concerning the latter, a characterization of the ground states of quantum many-body systems in terms of Hilbert-space-based measures of Bell nonlocality could provide novel and potentially profound insights on the degree of quantumness and quantumness hierarchies between different systems of quantum matter, in analogy with and complementing the very successful use of entanglement in condensed matter theory Laflorencie2016 ; Illuminati2022 ; Illuminati2023 . Indeed, some recent preliminary studies on the structure of Bell nonlocality in quantum many-body systems hint at complex patterns of quantum nonlocality beyond entanglement Nonlocmanybody2015 ; Belldepthmanybody2019 whose precise entropic and geometric quantifications might provide important insights on the structure and hierarchy of quantum collective phenomena.
The Hilbert-space characterization and quantification that we have introduced paves the way to the study of several important questions concerning quantum nonlocality. In entanglement theory, it is known that different entanglement measures can give rise to different orderings of quantum states; for instance, the entropy-based entanglement of formation and the entanglement negativity based on the PPT criterion induce a different ordering on the set of entangled Gaussian states Adesso2005 . Similar ordering issues might arise when determining the degree of Bell nonlocality of quantum states as quantified by the different geometric or entropic measures that we have introduced in the present work. A further important question concerns the characterization and quantification of extremal nonlocality in mixed states. In entanglement theory, maximally entangled mixed states are typically defined at fixed energy, or fixed degree of local and global entropies Adesso2003 ; Adesso2004 . An investigation along similar lines could lead to the identification of the maximally nonlocal mixed states at fixed global and/or local purities.
In our opinion, a particularly challenging open problem whose eventual solution would bear far-reaching consequences revolves around the question whether the analogs of Werner states, i.e., mixed states that are local and yet entangled, exist in infinite-dimensional Hilbert spaces. In particular, it would be very interesting to determine the existence of Werner Gaussian states of two- and multimode continuous-variable systems, and to characterize their properties.
Acknowledgements.
W.R. and M.T. are supported by JST Moonshot R&D Grant No. JPMJMS2061 and JST COI-NEXT Grant No. JPMJPF2221. F.I. acknowledges funding from the Italian Ministry of University and Research, call PRIN PNRR 2022, project “Harnessing topological phases for quantum technologies”, code P202253RLY, CUP D53D23016250001, and PNRR-NQSTI project ”ECoN: End-to-end long-distance entanglement in quantum networks”, CUP J13C22000680006;

### v4

@@
 Abstract
-We introduce a geometric framework for analyzing Bell nonlocality in Hilbert space, aiming to provide intrinsic measures of nonlocality based solely on the properties of quantum states for a given Bell inequality.
+We introduce a geometric framework for the study of Bell nonlocality in Hilbert space, aiming to provide intrinsic measures based solely on the properties of quantum states for a given Bell inequality.
 Recognizing that nonlocality is inherently measurement-dependent, our approach focuses on fixing a specific Bell scenario—that is, a chosen number of measurements and outcomes per party, along with a specific Bell inequality—and, for each quantum state, maximizing the violation over all possible measurement choices within the given scenario.
-We then characterize the geometry of the corresponding set of local states, defined as those states for which no violation occurs for any choice of measurements.
-Within this framework, we define geometric and entropic measures of nonlocality based on the distance of a given quantum state from the set of local states, using quantities such as trace distance, Hilbert-Schmidt distance, and relative entropy.
+We then characterize the geometry of the corresponding set of local states, defined as those states for which no violation occurs for any choice of the measurements.
+Within this framework, we define geometric measures of nonlocality based on the distance of a given quantum state from the set of local states, using quantities such as the trace distance, the Hilbert-Schmidt distance, and the relative entropy.
 We first establish the general formalism, emphasizing the challenges posed by the NP-hardness of characterizing local sets for arbitrary Bell scenarios.
 We then specialize to the case of the Clauser-Horne-Shimony-Holt (CHSH) inequality in two-qubit systems, where the local set is fully characterized and an explicit geometric analysis is possible.
-In this setting, we derive geometric measures of nonlocality for Bell-diagonal and Werner states, showing that the closest local state to a Werner state remains a Werner state, and similarly for Bell-diagonal states.
-For each quantum states, our results offer a metric-based characterization of Bell nonlocality that is independent of specific experimental implementations and naturally extend to different Bell scenarios when the corresponding set of local states is known.
+In this setting, we derive geometric measures of nonlocality for Bell-diagonal and Werner states, showing that the local state closest to a Werner state is a Werner state, and analogously for Bell-diagonal states.
+For each quantum state, our results offer a metric-based characterization of Bell nonlocality that is independent of specific experimental implementations and naturally extend to different Bell scenarios whenever the corresponding set of local states is accessible.
 I Introduction
@@
 Denoting the minima by $\tilde{a}_{i}^{loc}$, the solution reads
-$\tilde{a}_{1}^{loc}=\frac{-a_{2}\pm\sqrt{a_{2}^{2}+4}}{2},$
-$\tilde{a}_{2}^{loc}=\frac{-a_{1}\pm\sqrt{a_{1}^{2}+4}}{2},$
-$\tilde{a}_{3}^{loc}=a_{3}.$ (70)
+$a_{1}^{loc}=\pm\frac{a_{1}}{\sqrt{a_{1}^{2}+a_{2}^{2}}}{2},$
+$a_{2}^{loc}=\pm\frac{a_{2}}{\sqrt{a_{2}^{2}+a_{3}^{2}}},$
+$a_{3}^{loc}=a_{3}.$ (70)
 By Eq. (68) the (un-normalized) HS geometric measure of CHSH nonlocality for Bell-diagonal states is thus
-$\mathcal{M}_{HS}(\rho_{BD})=\frac{1}{4}\left[(a_{1}-a_{2})^{2}+\frac{a_{1}^{2}+a_{2}^{2}}{2}+(a_{2}-a_{1})\bigg{(}\sqrt{a_{1}^{2}+4}-\sqrt{a_{2}^{2}+4}\bigg{)}+2\right].$ (71)
+$\mathcal{M}_{HS}(\rho_{BD})=\frac{3}{4}\sum\limits_{i=x,y,z}(a_{i}-a_{i}^{loc})^{2}=\frac{3}{4}\left(\sqrt{a_{1}^{2}+a_{2}^{2}}\mp 1\right)^{2}$ (71)
+whose minimum clearly correspond to the minus sign (the plus sign in (71)), since $t_{x}^{2}+t_{y}^{2}\geq 1$.
 The remaining cases $|a_{2}^{loc}|,|a_{3}^{loc}|>|a_{1}^{loc}|$ and $|a_{1}^{loc}|,|a_{3}^{loc}|>|a_{2}^{loc}|$ can be immediately found by symmetry and the global minimum is determined over all possible cases.
@@
 (59).
+Note that this analytical expression is not always valid, as no constraint has been imposed on the positivity of the density operator associated with the set $\{a_{i}^{loc}\}$.
+Therefore, in order to use expression (71), one must also take such constraints into account.
+If the local state does not satisfy these constraints, the distance must be computed numerically.
 Using Eqs.

### v5

@@
-Geometric measures of CHSH quantum nonlocality: characterization, quantification, and comparison by distances and operations
+Geometric measures of quantum nonlocality: characterization, quantification, and comparison by distances and operations
 Abstract
-We introduce a geometric framework for the study of Bell nonlocality in Hilbert space, aiming to provide intrinsic measures based solely on the properties of quantum states for a given Bell inequality.
-Recognizing that nonlocality is inherently measurement-dependent, our approach focuses on fixing a specific Bell scenario—that is, a chosen number of measurements and outcomes per party, along with a specific Bell inequality—and, for each quantum state, maximizing the violation over all possible measurement choices within the given scenario.
-We then characterize the geometry of the corresponding set of local states, defined as those states for which no violation occurs for any choice of the measurements.
-Within this framework, we define geometric measures of nonlocality based on the distance of a given quantum state from the set of local states, using quantities such as the trace distance, the Hilbert-Schmidt distance, and the relative entropy.
-We first establish the general formalism, emphasizing the challenges posed by the NP-hardness of characterizing local sets for arbitrary Bell scenarios.
-We then specialize to the case of the Clauser-Horne-Shimony-Holt (CHSH) inequality in two-qubit systems, where the local set is fully characterized and an explicit geometric analysis is possible.
-In this setting, we derive geometric measures of nonlocality for Bell-diagonal and Werner states, showing that the local state closest to a Werner state is a Werner state, and analogously for Bell-diagonal states.
-For each quantum state, our results offer a metric-based characterization of Bell nonlocality that is independent of specific experimental implementations and naturally extend to different Bell scenarios whenever the corresponding set of local states is accessible.
-I Introduction
-Quantum nonlocality is generally understood as the impossibility of describing experimental correlations using a local hidden variable model Brunner2014 .
-A state is considered nonlocal if there exists at least one experimental configuration that leads to a violation of a Bell-type inequality, such as the Clauser-Horne-Shimony-Holt (CHSH) inequality.
-However, nonlocality is not solely a property of the state itself but also depends on the chosen measurement settings, making its geometric characterization particularly challenging.
-For instance, it is well known that some states do not violate the CHSH inequality, where both Alice and Bob perform two dichotomic measurements (the 2-2-2-2 scenario), but do violate more general Bell inequalities, such as the 3-3-2-2 scenario.
-Furthermore, there exist quantum states that do not violate any Bell inequality individually, but when multiple copies of the state are considered along with joint measurements, the resulting system can exhibit a violation of some Bell inequality Popescu1995 ; Gisin1996 ; Masanes2008 .
-This phenomenon is known as hidden nonlocality, or superactivation of Bell nonlocality.
-For instance, it has been shown that a state is distillable if and only if, after applying joint local filtering operations on an arbitrarily large number of copies, it leads to a violation of some Bell inequality Masanes2006 .
-Another example of nonlocality superactivation has been introduced in a recent work Bavaresco2025 , where the authors consider an appropriate (Bell-local) ancilla, a so-called catalyst, and local operations on the combined system.
-A natural approach to quantify the amount of nonlocality of a given state, is to study the geometry of the set of local states within Hilbert space while fixing a particular Bell inequality.
-Specifically, the distance of a given quantum state from this set provides a meaningful geometric measure of its nonlocality for the chosen protocol.
-Various distance-based measures, such as trace distance, Hilbert-Schmidt distance, and relative entropy, can be employed to quantify the deviation of a state from the local set.
-In this work, we explore the geometric and entropic characterizations of quantum correlations beyond entanglement Horodecki2009 , specifically focusing on Bell nonlocality within Hilbert space.
-Since entanglement and Bell nonlocality coincide for pure states, geometric approaches to entanglement naturally extend to nonlocality in such cases.
-However, for mixed states, while Bell nonlocality always implies entanglement, the converse does not necessarily hold, as exemplified by Werner states Werner1989 .
-This establishes a strict hierarchy between these two forms of quantum correlations in finite-dimensional systems.
-In the geometric approach such as the one introduced in the present work, the number of measurements for each party, their possible outcomes, and the related Bell inequalities are uniquely determined.
-Thereafter, for any given state, one possible choice is to maximize the violation of the inequality over all possible measurements performed by the two parties.
-If no violation of all the different Bell inequalities occurs, the state is considered local.
-The corresponding geometric measure of Bell nonlocality will then depend solely on the geometry of the set of local states and on the distance of the state from the set.
-Alternatively, for each possible type of measurement one can proceed to define the corresponding set of local states and the corresponding, measurement-dependent geometric measure of Bell nonlocality.
-In principle, the geometric approach described in the present work is very general, as it applies to any form of quantum nonlocality identified by some type of Bell inequality.
-What one needs to know is, for each instance, the type of measurement considered, and the corresponding set of local states.
-As already mentioned, this may turn out to be a very daunting task, given that in general the problem of determining the set of local states associated to a given Bell inequality is NP-hard in the number of measurements and, depending on the Hilbert space dimension, in the number of measurement outcomes.
-In the following, we will illustrate the general method by restricting to the CHSH inequality with maximum violation of the inequality over all possible measurements.
-Notably, for two-qubit systems, a complete parameterization of the set of local states with respect to the CHSH inequality is known, enabling precise geometric analysis.
-Clearly, this is not the most general case, as other forms of Bell inequalities exist.
-On the other hand, at least in the case of two qubits, it has been shown that the general problem is NP-hard, and therefore, the analysis must necessarily be limited to a relatively small number of measurements.
-In this context, it is well known that the number of inequalities other that the CHSH one is very limited Collins2004 .
-For example, in the case of three possible outcomes on each particle, there is only one other type of Bell inequality that arises and that is distinct from the CHSH one.
-We will proceed by first introducing the general theoretical framework before focusing on the violation of the CHSH inequality in two-qubit systems.
-For such systems, the set of local states is fully characterized, allowing us to illustrate our method effectively.
-We explicitly compute geometric measures of nonlocality for Bell-diagonal states and Werner states, relying on two key results that we establish: the closest local state to a Werner state remains a Werner state, and the closest local Bell-diagonal state remains a Bell-diagonal state.
-The work is organized as follows: In section II, we briefly review the basic concepts and structures in the theory of Bell nonlocality.
-Then, in section III, we introduce the set of free operations as the set of operations that map the set of local states into itself.
-In section IV we introduce the geometric and entropic measures of nonlocality and discuss their properties.
-We address the specific case of the CHSH inequality for two qubits in sections V and VI.
-Finally, in section VII we discuss our results and comment on some possible upcoming research directions.
-II Bell nonlocality for dichotomic observables: CHSH inequality
-Quantum nonlocality naturally arises in the context of Bell inequalities Bell1964 ; Clauser1969 ; Laskowski2004 .
-Consider two observers, Alice and Bob, who are spatially separated.
-Each of them has a set of observables, with Alice having $n_{a}$ and Bob having $n_{b}$.
-They do not communicate and independently choose which observable to measure, recording their respective outcomes.
-The correlations between these observables, determined by their choices and the results obtained, are used to construct Bell inequalities.
-These inequalities must hold in a classical framework, where we assume at most local hidden variable models and other commonly accepted assumptions, including choice independence, realism, and the absence of experimental loopholes.
-In bipartite local hidden variable theories, the conditional probability of Alice choosing observable $x$ and obtaining result $a$, and Bob choosing observable $y$ and obtaining result $b$, satisfies the following relation:
-$p(a,b|x,y)=\sum_{\lambda}p(\lambda)p(a|x,\lambda)p(b|y,\lambda),$ (1)
-where we can express $p(a,b|x,y)$ in quantum mechanical terms as:
-$p(a,b|x,y)={\rm Tr}(\rho^{AB}M_{a|x}\otimes M_{b|y}).$ (2)
-Here, $M_{a|x}$ represents a positive operator-valued measure (POVM) associated with outcome $a$ for observable $x$, and similarly, $M_{b|y}$ corresponds to outcome $b$ for observable $y$.
-It follows directly that for separable states, Eq.
+We introduce a geometric framework for studying Bell nonlocality in Hilbert space, where, for a given quantum state, nonlocality is quantified by the distance between the state and the set of local states.
+This approach applies to any Bell inequality and any measurement scenario.
+Whenever the local set is characterized, the proposed nonlocality measure can be computed explicitly.
+As a general result, we prove that for any scenario in arbitrary dimension the closest local state to a Werner state is itself a Werner state, and analogously, the closest local state to an isotropic state is again isotropic.
+In the two-qubit case, we further show that the closest local state to a Bell-diagonal state is Bell-diagonal as well.
+These structural results are independent of the specific Bell inequality considered, thus revealing intrinsic geometric features of these families of states and providing significant simplifications for computing the proposed measures.
+For the Clauser–Horne–Shimony–Holt (CHSH) inequality in two-qubit systems and the Collins–Gisin–Linden–Massar–Popescu (CGLMP) inequality for two qudits of arbitrary finite dimension, we derive explicit geometric measures of nonlocality for Bell-diagonal, Werner, and isotropic states using various distance metrics, including the trace, Hellinger, Hilbert–Schmidt distances, and relative entropy.
+Furthermore, we prove in all generality that for all scenarios in which the local set is not fully characterized, the geometric measures provide rigorous lower bounds on nonlocality.
+1 Introduction
+Quantum nonlocality is generally understood as the impossibility of describing experimental correlations using a local hidden variable (LHV) model [1, 2].
+A multipartite state is considered nonlocal if there exists at least one experimental configuration, involving non-communicating parties that can at most share randomness, which leads to a violation of a Bell-type inequality.
+It is clear that a state that appears local with respect to one Bell inequality may still violate a different Bell inequality [3, 4, 5].
+It can also happen that a local quantum state can violate a Bell inequality after being preprocessed by local filtering operations, revealing a hidden nonlocality [6, 7].
+Furthermore, there exist quantum states that do not violate any Bell inequality individually, but when multiple copies of the state are considered along with joint measurements, the resulting system can exhibit a violation of some Bell inequality.
+This phenomenon is known as superactivation of Bell nonlocality [8, 9].
+Determining whether a state is local, that is, whether its correlations admit a local hidden variable (LHV) description, can be challenging as it requires testing the state against an enormous number of possible experimental scenarios.
+For example, it has been shown that there exist two-qubit Werner states [10] that do not violate the Clauser–Horne–Shimony–Holt (CHSH) inequality [11] but nevertheless exhibit nonlocality when considering more than two measurements per party [12].
+Therefore, for a fixed scenario involving $n$ observers, each allowed to perform $m_{i}$ measurements labeled by $x_{i}$ ($i=1,\dots,n$), and each measurement having $\Delta_{i}$ possible outcomes $a_{i}$, we define the set $\mathcal{L}_{\vec{m},\vec{\Delta}}$ of local states, i.e.
+the set of states that admit a local hidden variable (LHV) description, as follows:
+$p(\vec{a}|\vec{x})=\sum_{\lambda}p(\lambda)\prod_{i=1}^{n}p(a_{i}|\lambda,x_{i}),$ (1)
+for any choice of $\vec{a}$ and $\vec{x}$, where
+$p(\vec{a}|\vec{x})={\rm Tr}\!\left[\rho\bigotimes_{i=1}^{n}M_{a_{i}|x_{i}}\right].$ (2)
+Here, $M_{a_{i}|x_{i}}$ denotes the positive operator-valued measure (POVM) element associated with outcome $a_{i}$ for measurement setting $x_{i}$.
+It follows directly that, for separable states, Eq.
 (2) implies Eq.
 (1).
-Consequently, nonlocal states must be entangled Horodecki1995 ; Werner1989 .
-However, some entangled states may still exhibit locality Werner1989 ; Masanes2006 .
-In the case of two dichotomic observables per observer, the Bell inequality takes the CHSH form:
-$E(x,y)-E(x,y^{\prime})+E(x^{\prime},y)+E(x^{\prime},y^{\prime})\leq 2,$ (3)
-where $E(x,y)$ denotes the statistical expectation value of the measurement outcomes for observables $x$ and $y$ in a given quantum state.
-For nonlocal states, Eq.
-(2) does not hold, leading to a violation of the CHSH inequality, Eq.
-(3).
-III Bell nonlocality: the free operations
-The set of local states, i.e., those states that do not violate any Bell inequality, will be denoted in the text as $\mathcal{L}$.
-In the framework of resource theory Chitambar2019 ; Theurer2017 ; Theurer2021 ; Gallego2015 ; Wolfe2020 ; Dakic2012 ; Pirandola2014 ; Streltsov2017 ; Buscemi2020 ; Contreras2019 ; Gour2008 ; Brandao2013 ; Buscemi2023 ; Barrett2005 ; Schmid2020 ; Kuroiwa2020 ; Gonda2023 ; Bartlett2007 ; DeVicente2014 , local states are also referred to as free states, meaning that they do not possess any resource—in this case, nonlocality.
-In the case of two-qubit systems and the CHSH inequality (3), the set of local states can be fully characterized Horodecki1995 ; Masanes2008 and will be discussed in detail in Section V.3 below.
-The set of operations that map local states into local states, also known as the set of free operations, consists of local operations and shared randomness (LOSR) Chitambar2019 ; Schmid2020 .
+Consequently, any nonlocal state must necessarily be entangled.
+More generally, the set of local states $\mathcal{L}$ may coincide with the intersection of several sets defined for fixed scenarios.
+As mentioned previously, nonlocality expresses the incompatibility of probability distributions obtained in quantum mechanics, or more general nonlocal theories, with those allowed by classical theories.
+This observation leads naturally to a geometric quantification of quantum nonlocality in terms of the trace distance between classical and quantum probability distributions [13].
+Such a geometrical approach provides a direct characterization of Bell nonlocality and allows to introduce well-behaved nonlocality monotones, in close analogy with the resource theory of entanglement.
+In this work, inspired by resource-theoretical approaches [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], we introduce a natural geometric quantifier of Bell nonlocality based on the distance of a given quantum state to the set of local states.
+This approach provides a perspective on nonlocality as an intrinsic property of the state and the considered measurement scenario, rather than as a feature associated solely with the resulting probability distribution.
+In the case where the convex set of local states is defined as the intersection of all possible $\mathcal{L}_{\vec{m},\vec{\Delta}}$, we obtain geometric measures of nonlocality that depend solely on the quantum state itself.
+We introduce different distance-based monotones, showing that they satisfy the same key properties enjoyed by the bona fide entropic and geometric measures of quantum entanglement and other quantum correlations, e.g.
+quantum discord, introduced in the course of time [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46].
+We prove that for Werner states the closest local state is always a Werner state, and analogously, for isotropic states the closest local state is itself isotropic.
+In the two-qubit case, we also show that the closest local state to a Bell-diagonal state is Bell-diagonal.
+These results hold independently of the specific Bell inequality under consideration, and therefore capture intrinsic geometric properties of these families of states.
+In practice, the set of local states is known only for specific scenarios.
+In such cases, the resulting geometric measure not only quantifies nonlocality within the given scenario but also provides a lower bound for more general ones.
+The work is organized as follows: In Sect.
+2, we introduce the set of free operations as the set of operations that map the set of local states into itself.
+In Sect.
+3 we introduce the geometric and entropic measures of nonlocality and discuss their properties.
+In Sect.
+4 we address the specific case of two qubit-systems, focusing in Sect.
+5 on the CHSH inequality.
+In Sect.
+6 we discuss isotropic and Werner states in arbitrary dimensions, including both bipartite and multipartite systems.
+As relevant examples, we compute the exact expressions of the geometric measures of nonlocality relative to the bipartite Collins–Gisin–Linden–Massar–Popescu (CGLMP) inequality for two-qudit isotropic states and we present a preliminary discussion of the multipartite Mermin inequality in arbitrary dimensions.
+Finally, in section 7 we review our results and comment on some possible upcoming research directions in the geometric theory of quantum nonlocality.
+2 Bell nonlocality: the free operations
+The set of local states will be denoted in the text as $\mathcal{L}$.
+In the framework of resource theory, local states are also referred to as free states, meaning that they do not possess any resource – in this case, nonlocality.
+The set of operations that map local states into local states, also known as the set of free operations, consists of local operations and shared randomness (LOSR) [14, 27].
 These operations can be generally defined as follows:
-[…]
+$\int_{\lambda}d\lambda\Phi^{A}_{\lambda}\otimes\Phi^{B}_{\lambda},$ (3)
 where $\Phi^{A}_{\lambda}$ and $\Phi^{B}_{\lambda}$ are completely positive and trace-preserving (CPTP) maps.
@@
 Since LOSR operations are also LOCC (while the converse is not true), they inherit many properties established for the extensively studied class of LOCC transformations.
-There are a few general properties that free operations should satisfy Theurer2021 ; Chitambar2019 :
+There are a few general properties that free operations should satisfy [14]:
   • Free operations preserve the set of free states.
@@
 In particular, LOSR satisfy all the properties listed above.
-IV Geometric and Entropic Measures of Bell nonlocality
-IV.0.1 Geometric measures
-For a given quantum state $\rho$ we take as measure $\mathcal{M}(\rho)$ of Bell nonlocality any contractive distance (with the exception of the Hilbert-Schmidt distance) $D$ between $\rho$ and the closest local state $\rho_{loc}$:
-$\mathcal{M}(\rho)=\min_{\rho_{loc}\in\sigma}D(\rho,\rho_{loc})\,,$ (5)
-where $\sigma\equiv\{\rho_{loc}\}$ denotes the set of local states.
-Physically meaningful and mathematically sound geometric measures of quantum nonlocality should satisfy the same key properties enjoyed by the bona fide entropic and geometric measures of quantum entanglement and other quantum correlations (e.g.
-quantum discord) introduced in the course of time Plenio2007 ; Christandl2004 ; Eisert1999 ; Wootters1998 ; Plenio2005 ; Wei2003 ; Illuminati2008 ; Uyanik2010 ; Geomdiscord2016 ; Geomdiscord2017 ; Vedral1997 ; Vedral1998 ; Baumgratz2014 ; Bengtsson2017 .
-In particular:
+3 Geometric and Entropic Measures of Bell nonlocality
+3.1 Geometric measures
+For a given quantum state $\rho$ we take as measure $\mathcal{M}(\rho)$ of Bell nonlocality any contractive distance (with the exception of the Hilbert-Schmidt distance) $D$ between $\rho$ and the closest local state $\rho^{loc}$:
+$\mathcal{M}(\rho)=\min_{\rho^{loc}\in\mathcal{L}}D(\rho,\rho^{loc})\,$, (4)
+where $\mathcal{L}\equiv\{\rho^{loc}\}$ denotes the set of local states.
+Physically meaningful and mathematically sound geometric measures of a quantum resource should satisfy the following key properties [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]:
   • Property 1 – vanishing on local states. By definition, this property is automatically satisfied by the relative entropy and by any chosen distance.
@@
 For any distance $D$ that is monotonically nonincreasing under an arbitrary CPTP map $\Phi$ that preserves the locality of a local state (LOSR do satisfy this property) we have
-    $\mathcal{M}(\rho)=\min_{\rho_{loc}\in\sigma}D(\rho,\rho_{loc})=D(\rho,\rho_{loc}^{min})\geq D(\Phi(\rho),\Phi(\rho_{loc}^{min}))\geq\min_{\rho_{loc}\in\sigma}D(\Phi(\rho),\rho_{loc})=\mathcal{M}(\Phi(\rho)).$ (6)
+    $\mathcal{M}(\rho)$ $=$ $\min_{\rho^{loc}\in\mathcal{L}}D(\rho,\rho^{loc})=D(\rho,\rho^{loc}_{min})$ (5)
+    $\geq$ $D(\Phi(\rho),\Phi(\rho^{loc}_{min}))\geq\min_{\rho^{loc}\in\mathcal{L}}D(\Phi(\rho),\rho^{loc})=\mathcal{M}(\Phi(\rho)).$
     It follows that any distance that is nonincreasing under CPTP maps defines a measure of nonlocality that is nonincreasing under LOSR.
@@
 Denoting by $a$ the convexity parameter with $0\leq a\leq 1$, we have:
-    $a\mathcal{M}(\rho_{1})+(1-a)\mathcal{M}(\rho_{2})$ $=$ $a\min_{\rho_{loc}\in\sigma}D(\rho_{1},\rho_{loc})+(1-a)\min_{\rho_{loc}\in\sigma}D(\rho_{2},\rho_{loc})$ (7)
-    $=$ $aD(\rho_{1},\rho_{loc}^{min})+(1-a)D(\rho_{2},\rho_{loc}^{min^{\prime}})$ (8)
-    $\geq$ $D(a\rho_{1}+(1-a)\rho_{2},a\rho_{loc}^{min}+(1-a)\rho_{loc}^{min^{\prime}})$ (9)
-    $\geq$ $\min_{\rho_{loc}\in\sigma}D(a\rho_{1}+(1-a)\rho_{2},\rho_{loc})$ (10)
-    $=$ $\mathcal{M}(a\rho_{1}+(1-a)\rho_{2}).$ (11)
+    $\hskip-56.9055pta\mathcal{M}(\rho_{1})+(1-a)\mathcal{M}(\rho_{2})$ $=$ $a\min_{\rho^{loc}\in\mathcal{L}}D(\rho_{1},\rho^{loc})+(1-a)\min_{\rho^{loc}\in\mathcal{L}}D(\rho_{2},\rho^{loc})$ (6)
+    $=$ $aD(\rho_{1},\rho^{loc}_{min})+(1-a)D(\rho_{2},\rho^{{}^{\prime}loc}_{min})$
+    $\geq$ $D(a\rho_{1}+(1-a)\rho_{2},a\rho^{loc}_{min}+(1-a)\rho^{{}^{\prime}loc}_{min})$
+    $\geq$ $\min_{\rho^{loc}\in\mathcal{L}}D(a\rho_{1}+(1-a)\rho_{2},\rho^{loc})$
+    $=$ $\mathcal{M}(a\rho_{1}+(1-a)\rho_{2}).$
     Joint convexity holds for the relative entropy by definition, and thus the latter is automatically nonincreasing under mixing.
@@
     $\mathcal{M}(\rho)\geq\sum_{i}p_{i}\mathcal{M}\left(\frac{\Lambda_{i}\rho\Lambda_{i}^{\dagger}}{p_{i}}\right),$ (7)
-    where the weights $p_{i}={\rm Tr}(\Lambda_{i}\rho\Lambda_{i}^{\dagger})$ and the set $\{\Lambda_{i}\}$ of Kraus operators is such that $\sum_{i}\Lambda_{i}^{\dagger}\Lambda_{i}=\mathbb{I}$ and $\Lambda_{i}\rho\Lambda_{i}^{\dagger}\in\sigma$ $\forall\rho^{loc}\in\sigma$.
+    where the weights $p_{i}={\rm Tr}(\Lambda_{i}\rho\Lambda_{i}^{\dagger})$ and the set $\{\Lambda_{i}\}$ of Kraus operators is such that $\sum_{i}\Lambda_{i}^{\dagger}\Lambda_{i}=\mathds{1}$ and $\Lambda_{i}\rho\Lambda_{i}^{\dagger}\in\mathcal{L}$ $\forall\rho^{loc}\in\mathcal{L}$.
 In a sense, property 4 might be considered even more important than property 2, since subselection is a process often readily available under controlled experimental conditions.
@@
 The reverse is obviously not true in general.
-Focusing only on the relative entropy, proof of property 4 is quite straightforward, in analogy with what has been shown to hold for the resource theory of entanglement Vedral1998 and quantum coherence Baumgratz2014 , and we provide it in Section IV.1.
+Focusing only on the relative entropy, proof of property 4 is quite straightforward, in analogy with what has been shown to hold for the resource theory of entanglement [44] and quantum coherence [45], and we provide it in Section 3.2.
   • Property 5 – computability.
@@
 We will study this subject case by case in the following by computing explicitly various measures for selected classes of two-qubit states.
-In the following sections, we introduce geometric and entropic measurements and analyze the properties they satisfy.
-IV.0.2 Relative entropy
+In the following, we introduce geometric and entropic measurements and analyze the properties they satisfy.
+3.1.1 Hilbert-Schmidt distance
+The Hilbert-Schmidt distance $D_{\mathrm{HS}}(\rho_{1},\rho_{2})$ between two density operators $\rho_{1}$ and $\rho_{2}$ is defined through the Hilbert-Schmidt norm as follows:
+$D_{\mathrm{HS}}(\rho_{1},\rho_{2})\,=\,\|\rho_{1}-\rho_{2}\|_{2}\,=\,\sqrt{{\rm Tr}\left[(\rho_{1}-\rho_{2})(\rho_{1}-\rho_{2})\right]}\,\,.$ (8)
+and the corresponding measure is
+$\mathcal{M}_{\mathrm{HS}}(\rho)\,=\,\min_{\rho^{loc}\in\mathcal{L}}D_{\mathrm{HS}}(\rho,\rho^{loc})\,.$ (9)
+The Hilbert-Schmidt (HS) distance has no direct operational meaning and, in general, may not monotonic under all CPTP transformations [46, 47].
+For instance, it increases under a trivial extension of local subsystems.
+On the other hand, it provides easily computable bounds on valid contractive distances that are often more difficult to compute.
+3.1.2 Hellinger distance
+The Hellinger distance is defined as the Euclidean two-norm of the difference of the square roots of the density matrices (taking squares for computational convenience):
+$D_{\mathrm{He}}(\rho_{1},\rho_{2})=\|\sqrt{\rho_{1}}-\sqrt{\rho_{2}}\|_{2}.$ (10)
+Given a pair of states that are diagonal in the same basis with, respectively, eigenvalues $\{e^{1}_{i}\}$ and $\{e^{2}_{i}\}$, one has:
+$D_{\mathrm{He}}(\rho_{1},\rho_{2})=\sqrt{2-2\sum_{i}\sqrt{e^{(1)}_{i}e^{(2)}_{i}}}=\sqrt{\displaystyle\sum_{i=1}^{4}\bigg(\sqrt{e_{i}^{(1)}}-\sqrt{e_{i}^{(2)}}\bigg)^{2}}.$ (11)
+The above is the Euclidean norm in the $\sqrt{e_{i}}$-space and coincides with the classical Hellinger distance quantifying the distance between the two classical probability distributions [48].
+The corresponding Hellinger measure of Bell nonlocality is thus defined as
+$\mathcal{M}_{\mathrm{He}}(\rho)=\min_{\rho^{loc}\in\mathcal{L}}D_{\mathrm{He}}(\rho,\rho^{loc})\,.$ (12)
+3.1.3 Bures distance
+The Bures distance is defined as [46]
+$D_{\mathrm{B}}(\rho_{1},\rho_{2})=\sqrt{2\big[1-\sqrt{F(\rho_{1},\rho_{2}})\big]}\,$, (13)
+where $F$ is the Uhlmann fidelity given by [49, 50]
+$F(\rho_{1},\rho_{2})=\bigg(\mathrm{Tr}\bigg[\sqrt{\sqrt{\rho_{1}}\rho_{2}\sqrt{\rho_{1}}}\bigg]\bigg)^{2}\,.$ (14)
+In the case of commuting density operators, the Bures and Hellinger distances coincide. The corresponding Bures measure of Bell nonlocality reads
+$\mathcal{M}_{\mathrm{B}}(\rho)=\min_{\rho^{loc}\in\mathcal{L}}D_{\mathrm{B}}(\rho,\rho^{loc})\,.$ (15)
+3.1.4 Trace distance
+The trace distance is defined as
+$D_{\mathrm{Tr}}(\rho_{1},\rho_{2})=\frac{1}{2}\|\rho_{1}-\rho_{2}\|_{1}=\frac{1}{2}{\rm Tr}\left[\sqrt{(\rho_{1}-\rho_{2})(\rho_{1}-\rho_{2})}\right]\,$, (16)
+where $\|\mathcal{O}\|_{1}\equiv\sqrt{\mathcal{O}^{\dagger}\mathcal{O}}$. For states that are diagonal in the same basis with eigenvalues $\{e^{1}_{i}\}$ and $\{e^{2}_{i}\}$ respectively, we have
+$D_{\mathrm{Tr}}(\rho_{1},\rho_{2})=\frac{1}{2}{\rm Tr}|\rho_{1}-\rho_{2}|=\frac{1}{2}\sum_{i}|e_{i}^{1}-e^{2}_{i}|\,.$ (17)
+The corresponding trace-distance measure of Bell nonlocality reads
+$\mathcal{M}_{\mathrm{Tr}}(\rho)=\min_{\rho^{loc}\in\mathcal{L}}D_{\mathrm{Tr}}(\rho,\rho^{loc}).$ (18)
+The trace distance defines the probability of error in the Helstrom measurement which is the optimal measurement to distinguish two quantum states.
+It is jointly convex and nonincreasing under CPTP maps [46, 50].
+As a consequence, the corresponding measure of Bell nonlocality Eq.
+(18) satisfies monotonicity under LOSR and monotonicity under mixing (convexity).
+3.2 Relative entropy
 Relative entropy is not a distance as it does not satisfy the triangle inequality and is not symmetric with respect to state interchange, but nonetheless it is widely used in the quantum information theory and the theory of entanglement because it enjoys nearly all of the monotonicity properties that are requested in a valid quantum resource theory.
 The relative entropy is defined as
-$S(\rho_{1}|\rho_{2})={\rm Tr}\rho_{1}\log_{2}\rho_{1}-{\rm Tr}\rho_{1}\log_{2}\rho_{2}.$ (13)
+$S(\rho_{1}||\rho_{2})={\rm Tr}\rho_{1}\log_{2}\rho_{1}-{\rm Tr}\rho_{1}\log_{2}\rho_{2}.$ (19)
 For states that are diagonal in the same basis with eigenvalues $e^{1}_{i}$ and $e^{2}_{i}$ respectively we have
-$S(\rho_{1}|\rho_{2})=\sum_{i}e^{1}_{i}\log_{2}e^{1}_{i}-\sum_{i}e^{1}_{i}\log_{2}e^{2}_{i}.$ (14)
+$S(\rho_{1}||\rho_{2})=\sum_{i}e^{1}_{i}\log_{2}e^{1}_{i}-\sum_{i}e^{1}_{i}\log_{2}e^{2}_{i}.$ (20)
 Hence, we define the measure of Bell nonlocality based on the relative entropy as
-$\mathcal{M}_{Re}(\rho)=\min_{\rho^{loc}\in\sigma}S(\rho|\rho^{loc}).$ (15)
+$\mathcal{M}_{\mathrm{Re}}(\rho)=\min_{\rho^{loc}\in\mathcal{L}}S(\rho||\rho^{loc}).$ (21)
 We will show here that the measure defined by Eq. (21) fulfills monotonicity under subselection, i.e.:
-$\mathcal{M}_{Re}(\rho)\geq\displaystyle\sum_{i}p_{i}\mathcal{M}_{Re}(\rho_{i}),$ (16)
+$\mathcal{M}_{\mathrm{Re}}(\rho)\geq\displaystyle\sum_{i}p_{i}\mathcal{M}_{\mathrm{Re}}(\rho_{i}),$ (22)
 where $\{\Lambda{i}\}$ is any set of Kraus operators, with weights $p_{i}={\rm Tr}(\Lambda_{i}\rho\Lambda_{i}^{\dagger})$, under the hypothesis that the involved Kraus operators are free, i.e.
-preserve the local states: $\sum_{i}\Lambda_{i}^{\dagger}\Lambda_{i}=\mathbb{I}$ and $\Lambda_{i}\rho\Lambda_{i}^{\dagger}\in\sigma$ $\forall\rho^{loc}\in\sigma$, where $\sigma$ is the set of local states.
-To obtain the last inequality, first of all note that since the relative entropy is invariant under the same unitary or isometry applied to both arguments, we can use the isometry $I=\sum_{i}|i\rangle_{aux}\otimes|i\rangle_{aux^{\prime}}\otimes V_{i}$ for any $V_{i}$ to derive the identity $S(\rho^{(1)}|\rho^{(2)})=S(I\rho^{(1)}I^{\dagger}|I\rho^{(2}I^{\dagger})$, for any couple of states $\rho^{(1)}$ and $\rho^{(2)}$, where $\{|i\rangle_{aux}\}_{i}$ is an orthogonal basis of an auxiliary system.
-Next, monotonicity of the relative entropy under the partial trace over the auxiliary system $aux$ implies
-$S(\rho^{(1)}|\rho^{(2)})\geq\sum_{i}S(\tilde{\rho}_{i}^{(1)}|\tilde{\rho}_{i}^{(2)}),$ (17)
-where we use unnormalized states $\tilde{\rho}_{i}^{(1)}=V_{i}\rho^{(1)}V_{i}^{\dagger}$ and $\tilde{\rho}_{i}^{(2)}=V_{i}\rho^{(2)}V_{i}^{\dagger}$.
-Following the argument derived in Vedral1998 we have
-$\sum_{i}S(\rho_{i}^{(1)}|\rho_{i}^{(2)})\geq\sum_{i}p_{i}S\left(\frac{\tilde{\rho}_{i}^{(1)}}{p_{i}}\bigg{|}\frac{\tilde{\rho}_{i}^{(2)}}{q_{i}}\right),$ (18)
+preserve the local states: $\sum_{i}\Lambda_{i}^{\dagger}\Lambda_{i}=\mathds{1}$ and $\Lambda_{i}\rho\Lambda_{i}^{\dagger}\in\mathcal{L}$ $\forall\rho^{loc}\in\mathcal{L}$, where $\mathcal{L}$ is the set of local states.
+To obtain the last inequality, first of all note that since the relative entropy is invariant under the same unitary or isometry applied to both arguments, we can use the isometry $I=\sum_{i}|i\rangle_{aux}\otimes|i\rangle_{aux^{\prime}}\otimes V_{i}$ for any $V_{i}$ to derive the identity $S(\rho^{(1)}|\rho^{(2)})=S(I\rho^{(1)}I^{\dagger}|I\rho^{(2)}I^{\dagger})$, for any couple of states $\rho^{(1)}$ and $\rho^{(2)}$, where $\{|i\rangle_{aux}\}_{i}$ is an orthogonal basis of an auxiliary system.
+Next, monotonicity of the relative entropy under the partial trace over the auxiliary system auxaux implies
+$S(\rho^{(1)}||\rho^{(2)})\geq\sum_{i}S(\tilde{\rho}_{i}^{(1)}||\tilde{\rho}_{i}^{(2)}),$ (23)
+where we use unnormalized states $\tilde{\rho}_{i}^{(1)}=V_{i}\rho^{(1)}V_{i}^{\dagger}$ and $\tilde{\rho}_{i}^{(2)}=V_{i}\rho^{(2)}V_{i}^{\dagger}$. Following the argument derived in [44] we have
+$\sum_{i}S(\rho_{i}^{(1)}||\rho_{i}^{(2)})\geq\sum_{i}p_{i}S\left(\frac{\tilde{\rho}_{i}^{(1)}}{p_{i}}\bigg|\frac{\tilde{\rho}_{i}^{(2)}}{q_{i}}\right),$ (24)
 where $p_{i}={\rm Tr}\ \tilde{\rho}_{i}^{(1)}$ and $q_{i}={\rm Tr}\ \tilde{\rho}_{i}^{(2)}$.
-Identifying $\rho\equiv\rho^{(1)}$, assuming that $\rho^{loc}\equiv\rho^{(2)}$ is a local sate and $V_{i}$ are Kraus operators which preserve locality, such that $V_{i}\rho^{loc}V_{i}^{\dagger}/q_{i}$ is still a local state, we conclude that
-$\mathcal{M}_{Re}(\rho)\geq\sum_{i}p_{i}\mathcal{M}_{Re}\left(\frac{V_{i}\rho V_{i}^{\dagger}}{p_{i}}\right).$ (19)
-IV.1 Distances and induced measures of nonlocality
-Here we discuss the properties of the geometric based measures of Bell nonlocality as defined, respectively, by Eq.
-(5) and Eq.
-(15).
-Concerning the geometric measures, we will consider those associated to the Hilbert-Schmidt norm and to some of the most relevant contractive distances.
-In general, each distance provides a slightly different clue on the distinguishability of quantum states.
-IV.1.1 Hilbert-Schmidt distance
-The Hilbert-Schmidt distance $D_{HS}(\rho_{1},\rho_{2})$ between two density operators $\rho_{1}$ and $\rho_{2}$ is defined trough the Hilbert-Schmidt norm as follows:
-$D_{HS}(\rho_{1},\rho_{2})\,=\,\|\rho_{1}-\rho_{2}\|_{HS}\,=\,\sqrt{{\rm Tr}\left[(\rho_{1}-\rho_{2})(\rho_{1}-\rho_{2})\right]}\,\,.$ (20)
-For convenience, we define the corresponding geometric measure of Bell nonlocality $\mathcal{M}_{HS}(\rho)$ through the squared distance:
-$\mathcal{M}_{HS}(\rho)\,=\,\min_{\rho^{loc}\in\sigma}D_{HS}^{2}(\rho,\rho^{loc})\,.$ (21)
-The Hilbert-Schmidt (HS) distance has no direct operational meaning and in general may not monotonic under all CPTP transformations Bengtsson2017 ; Piani2012 .
-For instance, it increases under a trivial extension of local subsystems.
-On the other hand, it provides easily computable bounds on valid contractive distances that are often more difficult to compute.
-IV.1.2 Hellinger distance
-The Hellinger distance is defined as the Euclidean two-norm of the difference of the square roots of the density matrices (taking squares for computational convenience):
-$D_{He}(\rho_{1},\rho_{2})=\|\sqrt{\rho_{1}}-\sqrt{\rho_{2}}\|_{2}.$ (22)
-Given a pair of states that are diagonal in the same basis with, respectively, eigenvalues $\{e^{1}_{i}\}$ and $\{e^{2}_{i}\}$, one has:
-$D_{He}(\rho_{1},\rho_{2})=\sqrt{2-2\sum_{i}\sqrt{e^{(1)}_{i}e^{(2)}_{i}}}=\sqrt{\displaystyle\sum_{i=1}^{4}\bigg{(}\sqrt{e_{i}^{(1)}}-\sqrt{e_{i}^{(2)}}\bigg{)}^{2}}.$ (23)
-The above is the Euclidean norm in the $\sqrt{e_{i}}-$space and coincides with the classical Hellinger distance quantifying the distance between the two classical probability distributions Ref_Hellinger1909 .
-The corresponding Hellinger measure of Bell nonlocality is thus defined as
-$\mathcal{M}_{He}(\rho)=\min_{\rho^{loc}\in\sigma}D_{He}^{2}(\rho,\rho^{loc})\,.$ (24)
-IV.1.3 Bures distance
-The Bures distance is defined as Bengtsson2017
-$D_{B}=\sqrt{2\big{[}1-\sqrt{F(\rho_{1},\rho_{2}})\big{]}}\,,$ (25)
-where $F$ is the Uhlmann fidelity given by Uhlmann1976 ; Jozsa1994 ; Nielsen2005
-$F(\rho_{1},\rho_{2})=\bigg{(}\text{Tr}\bigg{[}\sqrt{\sqrt{\rho_{1}}\rho_{2}\sqrt{\rho_{1}}}\bigg{]}\bigg{)}^{2}\,.$ (26)
-In the case of commuting density operators, the Bures and Hellinger distances coincide. The corresponding Bures measure of Bell nonlocality reads
-$\mathcal{M}_{B}(\rho)=\min_{\rho^{loc}\in\sigma}D_{B}(\rho,\rho^{loc})\,.$ (27)
-IV.1.4 Trace distance
-The trace distance is defined as
-$D_{Tr}(\rho_{1},\rho_{2})=\frac{1}{2}\|\rho_{1}-\rho_{2}\|_{1}=\frac{1}{2}{\rm Tr}\left[\sqrt{(\rho_{1}-\rho_{2})(\rho_{1}-\rho_{2})}\right]\,,$ (28)
-where $\|\mathcal{O}\|_{1}\equiv\sqrt{\mathcal{O}^{\dagger}\mathcal{O}}$. For states that are diagonal in the same basis with eigenvalues $\{e^{1}_{i}\}$ and $\{e^{2}_{i}\}$ respectively, we have
-$D_{Tr}(\rho_{1},\rho_{2})=\frac{1}{2}{\rm Tr}|\rho_{1}-\rho_{2}|=\frac{1}{2}\sum_{i}|e_{i}^{1}-e^{2}_{i}|\,.$ (29)
-The corresponding trace-distance measure of Bell nonlocality reads
-$\mathcal{M}_{Tr}(\rho)=\min_{\rho^{loc}\in\sigma}D_{Tr}(\rho,\rho^{loc}).$ (30)
-The trace distance defines the probability of error in the Helstrom measurement which is the optimal measurement to distinguish two quantum states.
-It is jointly convex and nonincreasing under CPTP maps Bengtsson2017 ; Nielsen2005 .
-As a consequence, the corresponding measure of Bell nonlocality Eq.
-(30) satisfies monotonicity under LOSR and monotonicity under mixing (convexity).
-V Application to two-qubit states: Formalism and CHSH locality
-In this section we specialize to two-qubit states in relation to the CHSH inequality Eq.
-(3).
-Before proceeding, we recollect some basic facts and derive some theorems that are necessary in order to apply the general framework outlined in the previous sections to the concrete instance of two-qubit states.
+Identifying $\rho\equiv\rho^{(1)}$, assuming that $\rho^{loc}\equiv\rho^{(2)}$ is a local state and $V_{i}$ are Kraus operators which preserve locality, such that $V_{i}\rho^{loc}V_{i}^{\dagger}/q_{i}$ is still a local state, we conclude that
+$\mathcal{M}_{\mathrm{Re}}(\rho)\geq\sum_{i}p_{i}\mathcal{M}_{\mathrm{Re}}\left(\frac{V_{i}\rho V_{i}^{\dagger}}{p_{i}}\right).$ (25)
+4 Application to two-qubit states
 An arbitrary two-qubit state $\rho$ can be represented as
 $\rho=\frac{1}{4}\sum_{i,j=0}^{3}\alpha_{ij}\,\sigma_{i}\otimes\sigma_{j},$ (26)
-where $\sigma_{0}=\mathbf{1}$ and the $\sigma_{i}$ ($i=1,2,3$) are the Pauli matrices satisfying ${\rm Tr}\sigma_{i}=0$, ${\rm Tr}\sigma_{i}\sigma_{j}=\delta_{ij}$, and the superposition coefficients read $\alpha_{00}=1$, $\alpha_{0i}=\text{Tr}\big{[}\rho\big{(}\mathbb{I}\otimes\sigma_{i}\big{)}\big{]}$, $\alpha_{i0}=\text{Tr}\big{[}\rho\big{(}\sigma_{i}\otimes\mathbb{I}\big{)}\big{]}$ and $\alpha_{ij}=\text{Tr}\big{[}\rho\big{(}\sigma_{i}\otimes\sigma_{j}\big{)}\big{]}$ for $i=1,2,3$; they can be recast in matrix form as follows:
-$\alpha=\begin{bmatrix}1&\vec{\alpha}_{0i}\\ \vec{\alpha}_{i0}&A\end{bmatrix},$ (32)
-where all entries are real, $\vec{\alpha}_{0i}$ is a row three-vector of elements $\{\alpha_{0i}\}$, $\vec{\alpha}_{i0}$ is a column three-vector of elements $\{\alpha_{i0}\}$, and $A$ is a $3\times 3$ real matrix.
-Explicitly:
-$\rho=\frac{1}{4}\bigg{(}\mathbf{1}+\sum_{i=1}^{3}\alpha_{0i}\mathbf{1}\otimes\sigma_{i}+\sum_{i=1}^{3}\alpha_{i0}\sigma_{i}\otimes\mathbf{1}+\sum_{i,j=1}^{3}A_{ij}\sigma_{i}\otimes\sigma_{j}\bigg{)}\,.$ (33)
-The condition $\text{Tr}[\rho^{2}]\leq 1$ leads to
-$\sum_{i}(\alpha_{0i}^{2}+\alpha_{i0}^{2})+\displaystyle\sum_{i,j}\alpha_{ij}^{2}\leq 3\,.$ (34)
-It is useful to consider the singular value decomposition of the matrix $A$:
+where $\sigma_{0}=\mathds{1}_{2}$ and $\sigma_{i}$ ($i=1,2,3$) denote the Pauli matrices, satisfying ${\rm Tr}[\sigma_{i}]=0$ and ${\rm Tr}[\sigma_{i}\sigma_{j}]=2\delta_{ij}$.
+The expansion coefficients are given by $\alpha_{00}=1$, $\alpha_{0i}={\rm Tr}\big[\rho\,(\mathds{1}_{2}\otimes\sigma_{i})\big]$, $\alpha_{i0}={\rm Tr}\big[\rho\,(\sigma_{i}\otimes\mathds{1}_{2})\big]$, and $\alpha_{ij}={\rm Tr}\big[\rho\,(\sigma_{i}\otimes\sigma_{j})\big]$ for $i=1,2,3$.
+These coefficients can be arranged in matrix form as
+$\alpha=\left[\begin{array}[]{cc}1&\vec{\alpha}_{0i}\\ \vec{\alpha}_{i0}&A\end{array}\right],$ (27)
+where all entries are real, $\vec{\alpha}_{0i}$ is a row vector with components $\{\alpha_{0i}\}$, $\vec{\alpha}_{i0}$ is a column vector with components $\{\alpha_{i0}\}$, and $A$ is a $3\times 3$ real matrix.
+As first result, we prove the following:
+Result 1. If $\rho$ is a local state with given $(\alpha_{0i},\alpha_{i0},A)$ then the state $\rho^{-}$ given by $(-\alpha_{0i},-\alpha_{i0},A)$ is also local.
+Proof. It suffices to consider a local transformation that for $i=1,2,3$ changes $\sigma_{i}$ into $\sigma_{i}$. $\square$
+Result 2.
+For any local state $\rho$ with given $(\alpha_{0i},\alpha_{i0},A)$ there exists a local state $\rho^{\prime}$ with the same $A$ but with maximally mixed reduced states, i.e., with $\alpha_{0i}=\alpha_{i0}=0$
+Proof. Consider the convex combination
+$\rho^{\prime}=\frac{1}{2}\rho+\frac{1}{2}\rho^{-}.$ (28)
+Indeed, for such a combination we have
+$\alpha=\left[\begin{array}[]{cc}\alpha_{00}&0\\ 0&A\end{array}\right].$ (29)
+Now, convex combinations of local states are local and this ends the proof. $\square$
+Two particularly relevant families of states are the Bell-diagonal states, i.e., convex mixtures of the four Bell states (see A for details):
+[…]
+and the subclass known as Werner states, which are convex combinations of the maximally mixed state and the singlet Bell state.
+They are parametrized by a single parameter $\omega$ and were introduced as the first example of states that are entangled yet do not necessarily violate the CHSH inequality [10].
+For this class of states we prove the following two propositions, holding for every scenario in which a set of Bell inequalities (just “Bell inequality” in the following) defines the convex set $\mathcal{L}$ of local states in that scenario:
+Proposition 1. For any Bell inequality, any jointly convex and unitarily invariant functional $D$ of pairs of two-qubit density matrices
+$\min_{\rho^{loc}\in\mathcal{L}}D(\rho_{w},\rho^{loc})=\min_{\rho^{loc}_{w}\in\mathcal{L}}D(\rho_{w},\rho^{loc}_{w})\,$, (31)
+where $\rho_{w}$ is a Werner state, $\rho^{loc}$ is a local state, and $\rho^{loc}_{w}$ is a local Werner state.
+Proof.
+The proof of Proposition 1 is shown in section 6 for isotropic states and generalized Werner states in higher dimensions.
+Here we just sketch the relevant steps.
+Starting from Eq.
+(4), we can apply the same $U\otimes U$ to both arguments of $D$ and make use of the unitarily invariance of $D$; then, we bound this quantity from below by twirling both arguments and using jointly convexity; finally we recognize that twirling projects $\rho_{w}$ itself and $\rho^{loc}$ into another (local) Werner state $\rho^{loc}_{w}$.
+So $D(\rho_{w},\rho^{loc}_{w})\leq D(\rho_{w},\rho^{loc})$, hence the closest state can be taken to be a Werner.
+$\square$
+Proposition 2.
+For any Bell inequality, any jointly convex and unitarily invariant functional $D$ of pairs of two-qubit density matrices, the closest local state to a Bell-diagonal state is Bell-diagonal:
+$\min_{\rho^{loc}\in\mathcal{L}}D(\rho_{\mathrm{BD}},\rho^{loc})=\min_{\rho_{\mathrm{BD}}^{loc}\in\mathcal{L}}D(\rho_{\mathrm{BD}},\rho_{\mathrm{BD}}^{loc})\,$, (32)
+where $\rho_{\mathrm{BD}}$ denotes a Bell-diagonal state, $\rho^{loc}$ a local state, and $\rho_{\mathrm{BD}}^{loc}$ a local Bell-diagonal state.
+Proof.
+The symmetry class that we consider contains the simultaneous local $\pi$-rotations of qubits around the three Pauli axes.
+Rotation $R_{1}$ around $\sigma_{1}$ changes $\sigma_{2}$ and $\sigma_{3}$ to $\sigma_{2}$ and $\sigma_{3}$, respectively.
+Rotations $R_{2}$ and $R_{3}$ operate analogously.
+These transformations are local and unitary, therefore, they cannot change the locality of a quantum state.
+Moreover, these transformations preserve Bell-diagonal states.
+Indeed, let us consider the general state representation Eq.
+(26).
+Under rotation $R_{1}$, we have the following transformation:
+$\hskip-28.45274pt\alpha=\left(\begin{array}[]{cccc}1&a_{01}&a_{02}&a_{03}\\ a_{10}&a_{11}&a_{12}&a_{13}\\ a_{20}&a_{21}&a_{22}&a_{23}\\ a_{30}&a_{31}&a_{32}&a_{33}\end{array}\right)\,\,\longrightarrow\,\,\alpha_{1}=\left(\begin{array}[]{cccc}1&a_{01}&-a_{02}&-a_{03}\\ a_{10}&a_{11}&-a_{12}&-a_{13}\\ -a_{20}&-a_{21}&a_{22}&a_{23}\\ -a_{30}&-a_{31}&a_{32}&a_{33}\end{array}\right).$ (33)
+The average $\rho^{\prime}=1/2(\rho+R_{1}(\rho))$ enjoys the representation
+$\hskip-28.45274pt\alpha_{1}^{\prime}=\left(\begin{array}[]{cccc}1&a_{01}&0&0\\ a_{10}&a_{11}&0&0\\ 0&0&a_{22}&a_{23}\\ 0&0&a_{32}&a_{33}\end{array}\right).$ (34)
+Applying rotation $R_{2}$ to $\rho^{\prime}$ and averaging the result with $\rho^{\prime}$ leads to the state with diagonal representation Eq.
+(30), i.e., a Bell-diagonal state.
+The total transformation that projects a given state to a Bell-diagonal state takes the compact form
+$\hskip-28.45274pt\rho\,\,\longrightarrow\,\,\frac{1}{2}\left(\frac{1}{2}(\rho+R_{1}(\rho))+R_{2}\left(\frac{1}{2}(\rho+R_{1}(\rho))\right)\right).$ (35)
+This transformation obviously preserves the Bell-diagonal states, does not change the locality, and is a composition of joint local unitaries and averaging.
+The proof of Proposition 2 is thus immediately completed by straightforwardly adapting the reasoning followed in the proof of Proposition 3. $\square$
+Note that the previous result is useful for quantifying nonlocality of any two-qubit system with maximally-mixed subsystems, since those states can be transformed in Bell-diagonal states by means of local unitary transformations which do not change the nonlocality of the state.
+Indeed, if one considers the singular value decomposition of the matrix $A$:
 $A=O^{1}DO^{2},$ (36)
 where $O^{1}$ and $O^{2}$ are orthogonal matrices and $D$ is diagonal and non-negative. The orthogonal transformations correspond to local changes of basis:
-$\rho=\frac{1}{4}\bigg{(}\mathbf{1}+\sum_{i=1}^{3}\alpha_{0i}\mathbf{1}\otimes\sigma_{i}+\sum_{i=1}^{3}\alpha_{i0}\sigma_{i}\otimes\mathbf{1}+\sum_{k=1}^{3}D_{k}\bigg{(}\sum_{i}O^{1}_{ik}\sigma_{i}\bigg{)}\otimes\bigg{(}\sum_{j}O^{2}_{kj}\sigma_{j}\bigg{)}\bigg{)}.$ (36)
-V.1 Bell-diagonal states
+$\rho=\frac{1}{4}$ $\bigg($ $\mathds{1}_{4}+\sum_{i=1}^{3}\alpha_{0i}\mathds{1}_{2}\otimes\sigma_{i}+\sum_{i=1}^{3}\alpha_{i0}\sigma_{i}\otimes\mathds{1}_{2}$ (37)
+$+$ $\sum_{k=1}^{3}D_{k}\bigg(\sum_{i}O^{1}_{ik}\sigma_{i}\bigg)\otimes\bigg(\sum_{j}O^{2}_{kj}\sigma_{j}\bigg)\bigg).$
+5 CHSH-nonlocality of two-qubit states
+Exploiting the results derived in the previous section, we can proceed to quantify the CHSH nonlocality of two-qubit states using the Hilbert space based distance and relative entropy measures with respect to the set of CHSH-local two-qubit states.
+In the case of two dichotomic observables per observer ($m=2$, $\Delta=2$), the Bell inequality takes the CHSH form:
+$E(x,y)-E(x,y^{\prime})+E(x^{\prime},y)+E(x^{\prime},y^{\prime})\leq 2,$ (38)
+where $(x,x^{\prime})$ and $(y,y^{\prime})$ are the observables of Alice and Bob, respectively.
+$E(x,y)$ denotes the statistical expectation value of the measurement outcomes for observables $x$ and $y$ in a given quantum state.
+For nonlocal states, Eq.
+(2) does not hold, leading to a violation of the CHSH inequality, Eq.
+(38).
+Let $(d_{1},d_{2},d_{3})$ be the singular values in the decomposition Eq.
+(36), and without loss of generality, let $d_{1}\geq d_{2}\geq d_{3}$.
+As proved in Ref.
+[51], a state is CHSH-local, i.e., it does not violate the CHSH inequality Eq.
+(38), if and only if
+$d_{1}^{2}+d_{2}^{2}\leq 1.$ (39)
+Explicit examples will include Werner states, Bell-diagonal states, and reduced convex combinations of Bell states.
+These important classes of states are routinely considered in the study of quantum resource theories, starting from the pioneering investigations on entanglement and discord [52, 36, 53].
+5.1 Werner states
+5.1.1 Hilbert-Schmidt distance
+Consider the Hilbert-Schmidt (HS) distance between a generic pair $\{\rho_{1},\rho_{2}\}$ of two-qubit states as induced by the HS norm Eq.
+(8).
+Using Eq.
+(26) and the properties of the Pauli matrices:
+$\hskip-28.45274pt\|\rho_{1}-\rho_{2}\|_{2}\,=\frac{1}{2}\,\|\alpha_{1}-\alpha_{2}\|_{2}\,.$ (40)
+It follows from Proposition 1, Eq.
+(31), that the closest CHSH-local state to a Werner state is also a Werner state with $w\leq 1/\sqrt{2}$.
+Hence, for CHSH-nonlocal Werner states, i.e., with $w\geq 1/\sqrt{2}$ :
+$\hskip-28.45274pt\mathcal{M}_{\mathrm{HS}}(\rho_{w})=\min_{\rho^{loc}_{w}\in\mathcal{L}}\left\|\rho_{w}-\rho^{loc}\right\|_{2}=\frac{\sqrt{3}}{2}\left(w-\frac{1}{\sqrt{2}}\right)\,$, (41)
+where $\mathds{1}$ is the identity matrix and $w\in\left(1/\sqrt{2},1\right]$. Correctly, the maximum is achieved by the maximally entangled Bell states corresponding to $w=1$ :
+$\mathcal{M}_{\mathrm{HS}}^{max}\,=\,{\frac{\sqrt{3}}{2}\left(1-\frac{1}{\sqrt{2}}\right)}\,.$ (42)
+The behavior of the normalized HS geometric measure of CHSH nonlocality for Werner states $\widetilde{\mathcal{M}}_{\mathrm{HS}}=\mathcal{M}_{\mathrm{HS}}(\rho_{w})/\mathcal{M}_{\mathrm{HS}}^{max}$ as a function of the $w$ parameter is reported in Fig.
+1.
+The plot is a parabola, as the HS distance is simply the Euclidean distance in the three-dimensional space of Fig.
+4.
+5.1.2 Hellinger distance
+Recalling again the consequences of Proposition 1, Eq. (31), the measure of CHSH nonlocality for Werner states based on the (squared) Hellinger distance, Eq. (11), reads
+$\hskip-28.45274pt\mathcal{M}_{\mathrm{He}}(\rho_{w})$ $=$ $\min_{\rho_{w}^{loc}\in\mathcal{L}}D^{2}_{\mathrm{He}}(\rho_{w},\rho_{w}^{loc})$ (43)
+$=$ $2-\frac{1}{2}\left(3\sqrt{(1-w)\left(1-\frac{1}{\sqrt{2}}\right)}+\sqrt{(1+3w)\left(1+\frac{3}{\sqrt{2}}\right)}\right)\,$,
+where $w\in(1/\sqrt{2},1]$.
+Here we have used the relation between the eigenvalues of the state density matrix and the Werner parameter $w$, which is derived directly from Eq.
+(70).
+As $|a_{1}|=|a_{2}|+|a_{3}|=w$, the signs of the matrix elements $a_{i}$ identify the states corresponding to different corners of the state tetrahedron in Fig.
+4.
+The Hellinger measure is maximized by the Bell states at $w=1$ :
+$\mathcal{M}_{\mathrm{He}}^{max}=2-\sqrt{1+\frac{3}{\sqrt{2}}}\,\,.$ (44)
+The behavior of the normalized Hellinger geometric measure of CHSH nonlocality for Werner states $\widetilde{\mathcal{M}}_{\mathrm{He}}=\mathcal{M}_{\mathrm{He}}(\rho_{w})/\mathcal{M}_{\mathrm{He}}^{max}$ as a function of the $w$ parameter is shown in Fig.
+1.
+5.1.3 Bures distance
+In the case of commuting operators, the Bures distance and the Hellinger distance coincide.
+This is indeed the case for Werner density matrices.
+Therefore, the Bures and Hellinger measures of CHSH nonlocality coincide on Werner states $\rho_{w}$:
+$\mathcal{M}_{\mathrm{Bu}}(\rho_{w})=\mathcal{M}_{\mathrm{He}}(\rho_{w})\,.$ (45)
+5.1.4 Trace distance
+The geometric measure of CHSH nonlocality of Werner states based on the trace distance Eq. (16) reads:
+$\mathcal{M}_{\mathrm{Tr}}(\rho_{w})=\min_{\rho_{w}^{loc}\in\mathcal{L}}D_{\mathrm{Tr}}(\rho_{w},\rho_{w}^{loc})=\frac{3}{4}\left|w-\frac{1}{\sqrt{2}}\right|=\frac{3}{4}\left(w-\frac{1}{\sqrt{2}}\right)\,$, (46)
+where the last equality follows from the fact that, without loss of generality, we can limit the analysis to the interval $w\in(1/\sqrt{2},1]$.
+The maximum value is achieved at $w=1$ by the Bell states:
+$\mathcal{M}_{\mathrm{Tr}}^{max}=\frac{3}{4}\left(1-\frac{1}{\sqrt{2}}\right)\,.$ (47)
+The normalized trace measure of CHSH nonlocality for Werner states $\widetilde{\mathcal{M}}_{\mathrm{Tr}}=\mathcal{M}_{\mathrm{Tr}}(\rho_{w})/\mathcal{M}_{\mathrm{Tr}}^{max}$ as a function of the $w$ parameter is shown in figure 1.
+5.1.5 Relative entropy
+It is straightforward to verify that for a nonlocal Werner state in the interval $w\geq 1/\sqrt{2}$, its relative entropy Eq.
+(20) with respect to the set $\sigma$ of local states is minimized by the local state with $w=1/\sqrt{2}$ , so that:
+$\hskip-56.9055pt\mathcal{M}_{\mathrm{Re}}(\rho_{w})=\min_{\rho_{w}^{loc}\in\mathcal{L}}S(\rho_{w}|\rho_{w}^{loc})$ $=$ $\frac{3(1-w)}{4}\log_{2}\left(\frac{3(1-w)}{4}\right)+\frac{1+3w}{4}\log_{2}\left(\frac{1+3w}{4}\right)+$ (48)
+- $\frac{3(1-w)}{4}\log_{2}\left(\frac{\sqrt{2}-1}{4\sqrt{2}}\right)-\frac{1+3w}{4}\log_{2}\left(\frac{\sqrt{2}+3}{4\sqrt{2}}\right)\,.$
+The maximum value achieved by the Bell states at $w=1$ reads
+$\mathcal{M}_{\mathrm{Re}}^{max}=2-\log_{2}{\bigg(1+\frac{3}{\sqrt{2}}\bigg)}\,.$ (49)
+The normalized relative entropy measure of CHSH nonlocality for Werner states $\widetilde{\mathcal{M}}_{\mathrm{Re}}=\mathcal{M}_{\mathrm{Re}}(\rho_{w})/\mathcal{M}_{\mathrm{Re}}^{max}$ as a function of the $w$ parameter is reported in Fig.
+1.
+Figure 1: Normalized geometric measures of CHSH nonlocality for Werner states as a function of the Werner parameter $w$ in the interval $\frac{1}{\sqrt{2}}\leq w\leq 1$.
+Dashed straight line: Trace and Hilbert-Schmidt measures $\widetilde{\mathcal{M}}_{\mathrm{Tr}}=\widetilde{\mathcal{M}}_{\mathrm{HS}}$.
+Dashed-dotted curve: Hellinger and Bures measure $\widetilde{\mathcal{M}}_{\mathrm{He}}=\widetilde{\mathcal{M}}_{\mathrm{Bu}}$.
+Solid curve: Relative entropy measure $\widetilde{\mathcal{M}}_{\mathrm{Re}}$.
+Dashed-dotted curve: Hellinger measure $\widetilde{\mathcal{M}}_{\mathrm{He}}$.
+All plotted quantities are dimensionless.
+5.2 Bell-diagonal states
+As mentioned before, the results found in this section can be applied to any two-qubit state with maximally disordered subsystems, since they can be transformed into Bell-diagonal states by local unitary transformations that do not change their nonlocality.
+Since Bell-diagonal states are diagonal in the Bell basis, the evaluation of the previously introduced distances is straightforward for this class of states.
+For a given state $\rho$, the closest local state can be identified through numerical analysis by exploiting the Horodecki criterion,
+$\max_{i,j\,,\,i\neq j}\{a_{i}^{2}+a_{j}^{2}\}=1\,$, (50)
+(see B for details).
+Here, let us consider Bell-diagonal states that are convex combinations of only two of the four Bell states.
+For instance, let us choose $e_{2}=e_{4}=0$, $e_{1}=p$, and $e_{3}=1-p$, with $p$ as a free parameter in the interval $[0,1]$.
+Such Bell-diagonal states represent all possible convex combinations of the first and third Bell states of Eq.
+(65).
+Moving to symplectic space and recalling Eqs.
+(70), this instance corresponds to $a_{1}=a$, $a_{2}=-a$, and $a_{3}=1$, with a=2p1a=2p-1.
+For this class of states, the locality condition Eq.
+(39) is fulfilled only for $p=\frac{1}{2}$; therefore, such states are nonlocal for $0\leq p\leq 1$, with the exception of $p=\frac{1}{2}$.
+The behavior of the four normalized measures of CHSH nonlocality as functions of the convex combination parameter $p$ in the half-interval $p\geq\frac{1}{2}$ is reported in Fig.
+2.
+We see that all measures vanish at $p=\frac{1}{2}$ and reach their maximum at $p=1$, i.e., on the first Bell state.
+For all measures, the normalization coefficients turn out to be the same as the ones of the nonlocality measures for Werner states, in agreement with Proposition 1, Eq.
+(31), and Proposition 2, Eq.
+(32).
+The case $p\leq\frac{1}{2}$ is readily obtained by symmetry.
+It is straightforward to verify that the above-described behavior is generic, i.e. it does not depend on the pair of Bell states that are chosen as entries of the convex combination.
+In the case of the Hilbert-Schmidt norm, Fig. 3 displays the corresponding nonlocality measure for a convex combination of two Bell states.
+Figure 2: Normalized measures of CHSH nonlocality for the convex combination of two Bell states as functions of the convex combination parameter $p$ in the interval $p\geq\frac{1}{2}$.
+The same behavior holds in the symmetric case $p\leq\frac{1}{2}$, with the maximum achieved in $p=0$.
+Black dashed curve: Trace measure $\widetilde{\mathcal{M}}_{\mathrm{Tr}}$.
+Black solid curve: Relative entropy measure $\widetilde{\mathcal{M}}_{\mathrm{Re}}$.
+Grey dashed curve: Hellinger measure $\widetilde{\mathcal{M}}_{\mathrm{He}}$.
+Grey solid curve: Hilbert-Schmidt measure $\widetilde{\mathcal{M}}_{\mathrm{HS}}$.
+All plotted quantities are dimensionless.
+Figure 3: Normalized Hilbert-Schmidt measure $\widetilde{\mathcal{M}}_{\mathrm{HS}}$ of CHSH nonlocality for Bell-diagonal states as a function of the two free parameters $e_{1}$ and $e_{2}$, with $e_{4}=0$, $e_{3}=1-e_{1}-e_{2}$, and normalization coefficient $\mathcal{M}_{\mathrm{HS}}^{max}$.
+The maximum $\widetilde{\mathcal{M}}_{\mathrm{HS}}=1$ is reached in the three cases $(e_{1},e_{2},e_{3})=(1,0,0)$; $(e_{1},e_{2},e_{3})=(0,1,0)$; and $(e_{1},e_{2},e_{3})=(0,0,1)$.
+Each case corresponds to a different Bell state.
+The central region of minima is composed by local states, i.e.
+states with $\widetilde{\mathcal{M}}_{\mathrm{HS}}=0$.
+All plotted quantities are dimensionless.
+6 Higher dimensions and multipartite systems
+The CHSH scenario has been generalized to two-qudit systems, for which some relevant inequalities are the CGLMP inequalities [54, 55] ($m=2,\Delta=d$), and to an arbitrary number of qubits, for $m=2$, $\Delta=2$ [57, 56, 58] and for arbitrary $m$ [4].
+These inequalities are tight, and so they completely characterize the set of local correlations for the selected scenario.
+For more complex scenarios, while it’s often the case that some Bell inequalities are known, they are generally not tight.
+6.1 Bipartite isotropic states
+[…]
+For any local isotropic state, the following result holds:
+Proposition 3. For any Bell inequality and for any jointly convex and unitarily invariant functional $D$ of pairs of two-qudit density matrices,
+$\min_{\rho^{loc}\in\mathcal{L}}D(\rho_{iso},\rho^{loc})=\min_{\rho^{loc}_{iso}\in\mathcal{L}}D(\rho_{iso},\rho_{iso}^{loc}),$ (51)
+where $\rho_{iso}$ is an isotropic state, $\rho^{loc}$ is a local state, and $\rho^{loc}_{iso}$ is a local isotropic state.
+Proof: The proof is based on the following observations:
+  1. 1.
+    Obs. 1: Invariance of isotropic states under $U\otimes U^{*}$ transformations:
+    $(U\otimes U^{*})\,\rho_{iso}\,(U^{\dagger}\otimes{U^{*}}^{\dagger})=\rho_{iso},\quad\forall\,U\ \mathrm{unitary}.$ (52)
+  2. 2.
+    Obs. 2: Local unitary transformations are resource non-generating (RNG) for nonlocality; in particular,
+    $(U\otimes U^{*})\,\rho^{loc}\,(U^{\dagger}\otimes{U^{*}}^{\dagger})\in\mathcal{L},$ (53)
+    where $\mathcal{L}$ denotes the set of local states.
+  3. 3.
+    Obs. 3: Any convex combination of local states is itself a local state.
+  4. 4.
+    Any state $\rho$ can be transformed into an isotropic state by applying the twirling operation:
+    $\int dU\,(U\otimes U^{*})\,\rho\,(U^{\dagger}\otimes{U^{*}}^{\dagger})=\rho_{iso}.$ (54)
+From the unitary invariance of $D$, we have
+$D(\rho_{iso},\rho^{loc}_{iso})$ $=$ $D(V\,\rho_{iso}\,V^{\dagger},V\,\rho^{loc}\,V^{\dagger})$ (55)
+$=$ $\int dV\,D(V\,\rho_{iso}\,V^{\dagger},V\,\rho^{loc}\,V^{\dagger})$ (56)
+$\geq$ $D\Bigg(\int dV\,V\,\rho_{iso}\,V^{\dagger},\int dV\,V\,\rho^{loc}\,V^{\dagger}\Bigg)$ (57)
+$=$ $D(\rho_{iso},\rho^{loc}_{iso})\,$, (58)
+where, in going from the third line to the fourth, we defined $V=U\otimes U^{*}$.
+Proposition 3 is general, since we do not specify any Bell inequality.
+In practice, for isotropic states, the set of local states that do not violate the CGLMP inequality is known [55, 54].
+This inequality is violated when $\omega>\frac{2}{I_{d}(QM)}$, with
+$I_{d}(QM)=4d\sum_{k=0}^{[d/2]-1}\left(1-\frac{2k}{d-1}\right)(q_{k}-q_{-(k+1)})$ (59)
+and
+$q_{k}=\frac{1}{2d^{3}\sin^{2}\left[\frac{\pi(k+\frac{1}{4})}{d}\right]}\,.$ (60)
+When $m=2$ and $\Delta=d=2$, the only facet inequality reduces to CHSH.
+This inequality is violated when $\omega>\frac{1}{\sqrt{2}}$.
+For $m=2$ and $\Delta=d=3$, the facet inequalities correspond to the CGLMP inequality, which is violated when $\omega>(6\sqrt{3}-9)/2$.
+[…]
+$\mathcal{M}_{\mathrm{HS}}(\rho_{iso})=\sqrt{1-\frac{1}{d^{2}}}\left(\omega-\frac{2}{I_{d}(QM)}\right)$ (61)
+$\mathcal{M}_{\mathrm{Tr}}(\rho_{iso})=\frac{2(d^{2}-1)}{d^{2}}\left(\omega-\frac{2}{I_{d}(QM)}\right)$ (62)
+$\mathcal{M}_{\mathrm{He}}$ $=$ $2-\frac{2}{d}\Bigg[(d^{2}-1)\sqrt{(1-\omega)\left(1-\frac{2}{I_{d}(QM)}\right)}$ (63)
+$+$ $\qquad\sqrt{\big[(d^{2}-1)\omega+1\big]\left[(d^{2}-1)\frac{2}{I_{d}(QM)}+1\right]}\Bigg]$
+$\mathcal{M}_{\mathrm{Re}}$ $=$ $\frac{(d^{2}-1)\omega+1}{d^{2}}\log{\left(\frac{(d^{2}-1)\omega+1}{d^{2}}\right)}+\frac{d^{2}-1}{d^{2}}\log{\left(\frac{1-\omega}{d^{2}}\right)}$ (64)
+$+$ $\frac{(d^{2}-1)\frac{2}{I_{d}(QM)}+1}{d^{2}}\log{\left(\frac{(d^{2}-1)\frac{2}{I_{d}(QM)}+1}{d^{2}}\right)}$
+$+$ $\frac{d^{2}-1}{d^{2}}\log{\left(\frac{1-\frac{2}{I_{d}(QM)}}{d^{2}}\right)}$
+6.2 Bipartite Werner states
+The result obtained in the previous section can also be extended to Werner states in higher dimensions, i.e., states that are invariant under transformations of the form $U\otimes U$ [10]:
+Proposition 4. For any Bell inequality and for any jointly convex and unitarily invariant functional $D$, the closest local state to a Werner state is again a Werner state.
+The proof follows exactly the same reasoning as for isotropic states.
+On the other hand, unlike the case of isotropic states, the precise value of the parameter that separates the nonlocal Werner states from the local ones is at present unknown.
+As a consequence, considering for instance states that are assured to be either CGLMP-local or CGLMP-nonlocal, our quantifiers provide, respectively, upper and lower bounds on the exact geometric measure of CGLMP nonlocality.
+In general, the full characterization of the set of local states for two-qudit systems in arbitrary Hilbert-space dimension requires a very ample and computationally complex effort, which will be the subject of a subsequent work.
+6.3 Multipartite systems
+Examples of Bell inequalities are also known in multipartite systems, such as the Mermin inequality [57] and the Mermin-Ardehali-Belinskii-Klyshko (MABK) inequalities [56].
+In the case of $n$ qubits and two dichotomic measurements per party ($m=2$, $\Delta=2$), a complete characterization of all possible Bell inequalities has been obtained [58].
+Note that the results of the previous propositions also hold for generalized Werner states, defined as those states that are invariant under local unitaries $V\equiv\bigotimes_{i=1}^{n}U$, since the proof follows directly for the same reasons discussed above.
+Nevertheless, for $n>2$, Werner states are no longer simply given by a convex combination of their symmetric and antisymmetric parts [60].
+Therefore, the study of Bell nonlocality in the multipartite case requires a separate and dedicated analysis which will be the subject of a forthcoming work.
+7 Discussion and outlook
+We introduced a geometric measure for Bell nonlocality valid in any scenario, whose defining elements are the state vectors in Hilbert space and their distances from the set of local states.
+For any Bell scenario, we have identified the basic traits of the geometric framework, including the set of local states and the free operations [61, 62, 63, 64, 65].
+The geometric approach is methodologically relevant because it enables a direct characterization and quantification of nonlocality in Hilbert space, thus placing it on the same conceptual footing as other fundamental measures of quantumness such as coherence, entanglement, steering, and quantum discord.
+Besides enabling the direct comparisons between nonlocality and different quantum resources in terms of inequalities and other hierarchical relations, the geometric framework allows in principle to investigate the problem of establishing whether sets of local operations exist that, by acting on maximally nonlocal states, allow to recover the entire Hilbert state space, in analogy with the case of coherence and entanglement.
+The quantification of Bell nonlocality for quantum states in Hilbert space is naturally introduced in terms of geometric measures, in close methodological analogy with the standard theory of geometric entanglement [38, 40] and the more recent geometric approaches to the quantification of other types of nonlocal correlations such as, e.g., the quantum discord [41, 42].
+Geometric measures define a structure and an ordering in the space of quantum states, and we have provided a detailed investigation of such structure in general Bell scenarios for any number of parties, arbitrary finite Hilbert space dimension, and different classes of contractive distances (trace, Hellinger, Bures).
+In particular, we proved rigorously that for contractive distances the closest local state to a Werner (isotropic) state is still a Werner (isotropic) state, both in the bipartite and in the multipartite case.
+We have shown that an analogous result holds for two-qubit Bell-diagonal states.
+We have confronted the ordering provided by contractive geometric measures of nonlocality with the non-contractive Hilbert-Schmidt distance, and we have introduced the relative entropy of nonlocality as a further benchmark.
+In a hierarchy of increasing complexity, moving forward from the CHSH two-qubit scenario, we have determined the explicit expression of the geometric measures of nonlocality for isotropic two-qudit states relative to the CGLMP inequality in any Hilbert-space dimension, and we have outlined the general features of a geometric theory of multipartite nonlocality relative to the Mermin and the MABK inequalities.
+In forthcoming works, for the bipartite case we plan to derive explicit analytical and numerical expressions of the geometric nonlocality of two-qudit Werner states in arbitrary dimension relative to the CGLMP inequality as well as in more general scenarios.
+For the multipartite case, we intend to compute the geometric measure of nonlocality of various classess of multipartite states, including generalized Werner states, in the Mermin, MABK, and more general scenarios [58, 4].
+Further generalizations and applications are conceivable along different directions, such as the identification of the set of local states and the ordering of nonlocal states in finite-dimensional Hilbert spaces of higher dimension, the extension of the geometric methods to different Bell inequalities both in the bipartite and in the multipartite case, the generalization to the field-theoretical setting of elementary particle physics, and the use of the Hilbert space-based measures of Bell nonlocality as tools in the investigation and characterization of quantum collective phenomena and quantum phase transitions.
+Indeed, the problem of characterizing the set of local states in arbitrary dimensions is known to be computationally hard, but by focusing on analytically tractable cases and leveraging known results on specific scenarios, it may be possible to provide concrete results and proofs that go beyond a conceptual proposal whenever partial knowledge of the local set is available or can be obtained by controlled approximation schemes.
+Concerning the investigation of quantum matter, a characterization of the ground states of quantum many-body systems in terms of Hilbert-space-based measures of Bell nonlocality could provide novel and potentially profound insights into the degree of quantumness and quantumness hierarchies between different systems of quantum matter, in analogy with and complementing the very successful use of entanglement in condensed matter theory [66, 67, 68].
+Indeed, some recent preliminary studies on the structure of Bell nonlocality in quantum many-body systems hint at complex patterns of quantum nonlocality beyond entanglement [69, 70], whose precise entropic and geometric quantification might provide important insights into the structure and hierarchy of quantum collective phenomena.
+The Hilbert-space characterization and quantification that we have introduced pave the way for the study of several important questions concerning quantum nonlocality.
+In entanglement theory, it is known that different entanglement measures can give rise to different orderings of quantum states; for instance, the entropy-based entanglement of formation and the entanglement negativity based on the PPT criterion induce a different ordering on the set of entangled Gaussian states [71].
+Similar ordering issues might arise when determining the degree of Bell nonlocality of quantum states, as quantified by the different geometric or entropic measures that we have introduced in the present work.
+A further important question concerns the characterization and quantification of extremal nonlocality in mixed states.
+In entanglement theory, maximally entangled mixed states are typically defined at fixed energy, or fixed degrees of local and global entropies [72, 73].
+An investigation along similar lines could lead to the identification of the maximally nonlocal mixed states at fixed global and/or local purities.
+In our opinion, a particularly challenging open problem whose eventual solution would bear far-reaching consequences revolves around the question of whether the analogs of Werner states, i.e., mixed states that are local and yet entangled, exist in infinite-dimensional Hilbert spaces.
+In particular, it would be very interesting to determine the existence of Werner Gaussian states of two- and multimode continuous-variable systems, and to characterize their properties.
+Acknowledgments
+W.R.
+and M.T.
+are supported by JST Moonshot R&D Grant No.
+JPMJMS2061 and JST COI-NEXT Grant No.
+JPMJPF2221.
+F.I.
+acknowledges funding from the Italian Ministry of University and Research, call PRIN PNRR 2022, project “Harnessing topological phases for quantum technologies”, code P202253RLY, CUP D53D23016250001, and PNRR-NQSTI project ECoN: “End-to-end long-distance entanglement in quantum networks”, CUP J13C22000680006.
+Appendix A Two-qubit Bell-diagonal states and Werner states
+A.1 Bell-diagonal states
 Given the Bell basis formed by the four maximally entangled Bell states
@@
 their convex combinations define the Bell-diagonal states:
-$\rho_{BD}=\displaystyle\sum_{i=1}^{4}e_{i}\ket{\Psi_{i}}\bra{\Psi_{i}}\,,$ (38)
+$|,\rho_{\mathrm{BD}}=\displaystyle\sum_{i=1}^{4}e_{i}\ket{\Psi_{i}}\bra{\Psi_{i}}\,$, (66)
 where $\{0\leq e_{i}\leq 1\}$, with $\sum_{i=1}^{4}e_{i}=1$. By resorting to the notation of Eq. (26), a generic Bell-diagonal state Eq. (66) can be rewritten as
-$\rho_{BD}=\frac{1}{4}\bigg{[}\mathcal{I}_{2}\otimes\mathcal{I}_{2}+\displaystyle\sum_{i=1}^{3}a_{i}\sigma_{i}\otimes\sigma_{i}\bigg{]}\,,$ (39)
-where $\mathcal{I}_{2}$ denotes the identity matrix in two dimensions. Matrix $\alpha$ of Eq. (32) can then be recast in the form
-$\alpha_{BD}=\begin{bmatrix}1&0&0&0\\ 0&a_{1}&0&0\\ 0&0&a_{2}&0\\ 0&0&0&a_{3}\end{bmatrix}\,.$ (40)
+$\rho_{\mathrm{BD}}=\frac{1}{4}\bigg[\mathds{1}_{2}\otimes\mathds{1}_{2}+\displaystyle\sum_{i=1}^{3}a_{i}\sigma_{i}\otimes\sigma_{i}\bigg]\,$, (67)
+where $\mathds{1}_{2}$ denotes the identity matrix in two dimensions. Matrix $\alpha$ of Eq. (27) can then be recast in the form
+$\alpha_{\mathrm{BD}}=\left[\begin{array}[]{cccc}1&0&0&0\\ 0&a_{1}&0&0\\ 0&0&a_{2}&0\\ 0&0&0&a_{3}\end{array}\right]\,.$ (68)
 Figure 4: Semitransparent tetrahedron: set of Bell-diagonal states.
@@
 Collecting the above results, for any Bell-diagonal state it is rather straightforward to derive the following formula relating the coefficients $\vec{a}=(1,a_{1},a_{2},a_{3})^{T}$ of the matrix $\alpha$ and the eigenvalues $\vec{e}=(e_{1},...,e_{4})^{T}$ of the density matrix:
-$\begin{bmatrix}1\\ a_{1}\\ a_{2}\\ a_{3}\end{bmatrix}=\begin{bmatrix}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ -1&1&1&-1\end{bmatrix}\begin{bmatrix}e_{1}\\ e_{2}\\ e_{3}\\ e_{4}\end{bmatrix}\,.$ (41)
+$\left[\begin{array}[]{c}1\\ a_{1}\\ a_{2}\\ a_{3}\end{array}\right]=\left[\begin{array}[]{cccc}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ -1&1&1&-1\end{array}\right]\left[\begin{array}[]{c}e_{1}\\ e_{2}\\ e_{3}\\ e_{4}\end{array}\right]\,.$ (69)
 The matrix in the above expression is unitary upon multiplication by $1/2$. It follows that
-$\begin{bmatrix}e_{1}\\ e_{2}\\ e_{3}\\ e_{4}\end{bmatrix}=\frac{1}{4}\begin{bmatrix}1&1&1&-1\\ 1&1&-1&1\\ 1&-1&1&1\\ 1&-1&-1&-1\end{bmatrix}\begin{bmatrix}1\\ a_{1}\\ a_{2}\\ a_{3}\end{bmatrix}\,,$ (42)
-so that positivity and normalization of the sum of the eigenvalues allow to represent the set of all Bell-diagonal states parameterized by the three-dimensional vector $(a_{1},a_{2},a_{3})$ as shown in Fig.
+$\left[\begin{array}[]{c}e_{1}\\ e_{2}\\ e_{3}\\ e_{4}\end{array}\right]=\frac{1}{4}\left[\begin{array}[]{cccc}1&1&1&-1\\ 1&1&-1&1\\ 1&-1&1&1\\ 1&-1&-1&-1\end{array}\right]\left[\begin{array}[]{c}1\\ a_{1}\\ a_{2}\\ a_{3}\end{array}\right]\,.$ (70)
+so that the positivity and normalization of the sum of the eigenvalues allow us to represent the set of all Bell-diagonal states parameterized by the three-dimensional vector $(a_{1},a_{2},a_{3})$ as shown in Fig.
 (4).
-V.2 Werner states
+A.2 Werner states
 A particular subclass of the Bell-diagonal states are the Werner states $\rho_{w}$, described by one free parameter $|a_{1}|=|a_{2}|=|a_{3}|\equiv w$.
@@
 F.
-Werner Werner1989 , is important as it provides the paramount example of a set of states that for $\frac{1}{3}\leq w\leq\frac{1}{\sqrt{2}}$ are entangled but nevertheless local.
-V.3 CHSH-local states and their properties
-Let $(d_{1},d_{2},d_{3})$ be the singular values in the decomposition Eq.
-(35) and, without loss of generality let $d_{1}\geq d_{2}\geq d_{3}$.
-As proved in Ref.
-Horodecki1995 , a state is CHSH-local, i.e.
-it does not violate the CHSH inequality Eq.
-(3), if and only if
-$d_{1}^{2}+d_{2}^{2}\leq 1.$ (43)
-We prove the following:
-Lemma 1. If $\rho$ is a CHSH-local state with given $(\alpha_{0i},\alpha_{i0},A)$ then the state $\rho^{-}$ given by $(-\alpha_{0i},-\alpha_{i0},A)$ is also CHSH-local.
-Proof. It suffices to consider a local transformation that for $i=1,2,3$ changes $\sigma_{i}$ into $-\sigma_{i}$.
-Lemma 2. For any CHSH-local state $\rho$ with given $(\alpha_{0i},\alpha_{i0},A)$ there exists a CHSH-local state $\rho^{\prime}$ with the same $A$ but with $\alpha_{0i}=\alpha_{i0}=0$
-Proof. Consider the convex combination
-$\rho^{\prime}=\frac{1}{2}\rho+\frac{1}{2}\rho^{-}.$ (44)
-Indeed, for such a combination we have
-$\alpha=\begin{bmatrix}\alpha_{00}&0\\ 0&A\end{bmatrix}.$ (45)
-Now, convex combinations of local states are local: by virtue of the property of shared randomness, if there exists a hidden variable model for each of the states in the combination, then there must exist a hidden variable model for their classical statistical mixture.
-V.4 Properties of symmetric states
-Proposition 1. For any jointly convex and unitarily invariant functional $D$ of pairs of two-qubit density matrices
-$\min_{\rho^{loc}\in\sigma}D(\rho_{w},\rho^{loc})=\min_{\rho^{loc}_{w}\in\sigma}D(\rho_{w},\rho^{loc}_{w})\,,$ (46)
-where $\rho_{w}$ is a Werner state, $\rho^{loc}$ is a CHSH-local state, and $\rho^{loc}_{w}$ is a CHSH-local Werner state.
-Proof. The proof is based on the following observations:
-  • Obs. 1. Invariance of Werner states under the same local unitary transformations on both subsystems:
-    $\forall\ {\rm unitary}\ U:\qquad\left(U\otimes U\right)\rho_{w}\left(U^{\dagger}\otimes U^{\dagger}\right)=\rho_{w}.$ (47)
-  • Obs.
-2.
-Local unitary transformations are resource non-generating (RNG) for nonlocality; in particular, $\left(U\otimes U\right)\rho^{loc}\left(U^{\dagger}\otimes U^{\dagger}\right)$ remains local.
-  • Obs. 3. Any convex combination of local states is a local state.
-  • Obs. 4. Twirling always projects on Werner states. Therefore, we have that for any given class of density matrices $\rho^{a}$
-[…]
-    where the integral is taken over the Haar measure on the set of unitary transformations.
-From the unitary invariance of $D$ we have
-$D(\rho_{w},\rho^{loc})$ $=$ $D\left(U\otimes U\rho_{w}U^{\dagger}\otimes U^{\dagger}\,,\,U\otimes U\rho^{loc}U^{\dagger}\otimes U^{\dagger}\right)$ (49)
-[…]
-[…]
-$=$ $D(\rho_{w},\rho^{loc}_{w})\,.$ (52)
-Here the first and second lines come from unitary invariance of $D$; the third follows from the joint convexity; and the fourth from Obs. 1 - 4 above.
-Similar propositions hold for general Bell-diagonal states; however, in the general case the proof holds for fairly smaller symmetry classes.
-Proposition 2. For any jointly convex and unitarily invariant functional $D$ of pairs of two-qubit density matrices, the closest CHSH-local state to a Bell-diagonal state is Bell-diagonal:
-$\min_{\rho_{loc}\in\sigma}f(\rho_{BD},\rho_{loc})=\min_{\rho_{BD,loc}\in\sigma}f(\rho_{BD},\rho_{BD,loc})\,,$ (53)
-where $\rho_{BD}$ denotes a Bell-diagonal state, $\rho_{loc}$ a CHSH-local state, and $\rho_{BD,loc}$ a CHSH-local Bell-diagonal state.
-Proof.
-The symmetry class that we consider contains the simultaneous local $\pi$-rotations of qubits around the three Pauli axes.
-Rotation $R_{1}$ around $\sigma_{1}$ changes $\sigma_{2}$ and $\sigma_{3}$ to $-\sigma_{2}$ and $-\sigma_{3}$ respectively.
-Rotations $R_{2}$ and $R_{3}$ operate analogously.
-These transformations are local and unitary, therefore they cannot change the locality of a quantum state.
-Moreover, these transformations preserve Bell-diagonal states.
-Indeed, let us consider the general state representation Eq.
-(31).
-Under rotation $R_{1}$ we have the following transformation:
-$\alpha=\begin{pmatrix}1&a_{01}&a_{02}&a_{03}\\ a_{10}&a_{11}&a_{12}&a_{13}\\ a_{20}&a_{21}&a_{22}&a_{23}\\ a_{30}&a_{31}&a_{32}&a_{33}\end{pmatrix}\,\,\longrightarrow\,\,\alpha_{1}=\begin{pmatrix}1&a_{01}&-a_{02}&-a_{03}\\ a_{10}&a_{11}&-a_{12}&-a_{13}\\ -a_{20}&-a_{21}&a_{22}&a_{23}\\ -a_{30}&-a_{31}&a_{32}&a_{33}\end{pmatrix}.$ (54)
-The average $\rho^{\prime}=1/2(\rho+R_{1}(\rho))$ enjoys the representation
-$\alpha_{1}^{\prime}=\begin{pmatrix}1&a_{01}&0&0\\ a_{10}&a_{11}&0&0\\ 0&0&a_{22}&a_{23}\\ 0&0&a_{32}&a_{33}\end{pmatrix}.$ (55)
-Applying rotation $R_{2}$ to $\rho^{\prime}$ and averaging the result with $\rho^{\prime}$ leads to the state with diagonal representation Eq.
-(39), i.e.
-a Bell-diagonal state.
-The total transformation that projects a given state to a Bell-diagonal state takes the compact form
-$\rho\,\,\longrightarrow\,\,\frac{1}{2}\left(\frac{1}{2}(\rho+R_{1}(\rho))+R_{2}\left(\frac{1}{2}(\rho+R_{1}(\rho))\right)\right).$ (56)
-This transformation obviously preserves the Bell-diagonal states, does not change the locality, and is a composition of joint local unitaries and averaging.
-The proof of Proposition 2 is thus immediately completed by straightforwardly adapting the reasoning followed in the proof of Proposition 1.
-VI CHSH nonlocality of two-qubit states
-Exploiting the results derived in the previous section, we can proceed to quantify the CHSH nonlocality of two-qubit states using the Hilbert space based distance and relative entropy measures with respect to the set of CHSH-local two-qubit states.
-Explicit examples will include Werner states, Bell-diagonal states, and reduced convex combinations of Bell states.
-These important classes of states are routinely considered in the study of quantum resource theories, starting from the pioneering investigations on entanglement and discord Wootters1997 ; Wootters1998 ; Henderson_Vedral2001 .
-VI.1 Werner states
-VI.1.1 Hilbert-Schmidt distance
-Consider the squared Hilbert-Schmidt (HS) distance defined on a generic pair $\{\rho_{1},\rho_{2}\}$ of two-qubit states as induced by the HS norm Eq.
-(20).
-Using Eq.
-(31) and the properties of the Pauli matrices:
-$\|\rho_{1}-\rho_{2}\|_{HS}\,=\,\|\alpha_{1}-\alpha_{2}\|_{HS}\,.$ (57)
-It follows from Proposition 1, Eq.
-(46), that the closest CHSH-local state to a Werner state is also a Werner state with $w\leq 1/\sqrt{2}$.
-Hence, for CHSH-nonlocal Werner states, i.e.
-with $w\geq 1/\sqrt{2}$ :
-$\mathcal{M}_{HS}(\rho_{w})=\min_{\rho^{loc}_{w}\in\sigma}\left\|\rho_{w}-\rho^{loc}\right\|_{HS}^{2}=\left\|\frac{1}{\sqrt{2}}\mathbf{1}-w\mathbf{1}\right\|_{HS}^{2}=3\left(\frac{1}{\sqrt{2}}-w\right)^{2}\,,$ (58)
-where $\mathbf{1}$ is the identity matrix and $w\in\left(1/\sqrt{2},1\right]$. Correctly, the maximum is achieved by the maximally entangled Bell states corresponding to $w=1$ :
-$\mathcal{M}_{HS}^{max}\,=\,\frac{3}{2}\left(3-2\sqrt{2}\right)\,.$ (59)
-The behavior of the normalized HS geometric measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{HS}=\mathcal{M}_{HS}(\rho_{w})/\mathcal{M}_{HS}^{max}$ as a function of the $w$ parameter is reported in Fig.
-2.
-The plot is a parabola, as the the HS distance is simply the Euclidean distance in the three-dimensional space of Fig.
-1.
-VI.1.2 Hellinger distance
-Recalling again the consequences of Proposition 1, Eq. (46), the measure of CHSH nonlocality for Werner states based on the (squared) Hellinger distance, Eq. (23), reads
-$\mathcal{M}_{He}(\rho_{w})=\min_{\rho_{w}^{loc}\in\sigma}D^{2}_{He}(\rho_{w},\rho_{w}^{loc})=2-\frac{1}{2}\left(3\sqrt{(1-w)\left(1-\frac{1}{\sqrt{2}}\right)}+\sqrt{(1+3w)\left(1+\frac{3}{\sqrt{2}}\right)}\right)\,,$ (60)
-where $w\in(1/\sqrt{2},1]$.
-Here we have used the relation between the eigenvalues of the state density matrix and the Werner parameter $w$ which is derived directly from Eq.
-(42).
-As $|a_{1}|=|a_{2}|+|a_{3}|=w$, the signs of the matrix elements $a_{i}$ identify the states corresponding to different corners of the state tetrahedron in Fig.
-1.
-The Hellinger measure is maximized by the Bell states at $w=1$ :
-$\mathcal{M}_{He}^{max}=2-\sqrt{1+\frac{3}{\sqrt{2}}}\,\,.$ (61)
-The behavior of the normalized Hellinger geometric measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{He}=\mathcal{M}_{He}(\rho_{w})/\mathcal{M}_{He}^{max}$ as a function of the $w$ parameter is shown in Fig.
-2.
-VI.1.3 Bures distance
-In the case of commuting operators the Bures distance and the Hellinger distance coincide.
-This is indeed the case for Werner density matrices.
-Therefore, the Bures and Hellinger measures of CHSH nonlocality coincide on Werner states $\rho_{w}$:
-$\mathcal{M}_{Bu}(\rho_{w})=\mathcal{M}_{He}(\rho_{w})\,.$ (62)
-VI.1.4 Trace distance
-The geometric measure of CHSH nonlocality of Werner states based on the trace distance Eq. (28) reads:
-$\mathcal{M}_{Tr}(\rho_{w})=\min_{\rho_{w}^{loc}\in\sigma}D_{Tr}(\rho_{w},\rho_{w}^{loc})=\frac{3}{2}\left|w-\frac{1}{\sqrt{2}}\right|=\frac{3}{2}\left(w-\frac{1}{\sqrt{2}}\right)\,,$ (63)
-where the last equality follows from the fact that without loss of generality we can limit the analysis to the interval $w\in(1/\sqrt{2},1]$. The maximum value is achieved at $w=1$ by the Bell states:
-$\mathcal{M}_{Tr}^{max}=\frac{3}{2}\left(1-\frac{1}{\sqrt{2}}\right)\,.$ (64)
-The normalized trace measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{Tr}=\mathcal{M}_{Tr}(\rho_{w})/\mathcal{M}_{Tr}^{max}$ as a function of the $w$ parameter is shown in figure 2.
-VI.1.5 Relative entropy
-It is straightforward to verify that for a nonlocal Werner state in the interval $w\geq 1/\sqrt{2}\,$ its relative entropy Eq.
-(14) with respect to the set $\sigma$ of local states is minimized by the local state with $w=1/\sqrt{2}$ , so that:
-$\mathcal{M}_{Re}(\rho_{w})=\min_{\rho_{w}^{loc}\in\sigma}S(\rho_{w}|\rho_{w}^{loc})$ $=$ $\frac{3(1-w)}{4}\log_{2}\left(\frac{3(1-w)}{4}\right)+\frac{1+3w}{4}\log_{2}\left(\frac{1+3w}{4}\right)+$ (65)
--- $\frac{3(1-w)}{4}\log_{2}\left(\frac{\sqrt{2}-1}{4\sqrt{2}}\right)-\frac{1+3w}{4}\log_{2}\left(\frac{\sqrt{2}+3}{4\sqrt{2}}\right)\,.$
-The maximum value achieved by the Bell states at $w=1$ reads
-$\mathcal{M}_{Re}^{max}=2-\log_{2}{\bigg{(}1+\frac{3}{\sqrt{2}}\bigg{)}}\,.$ (66)
-The normalized relative entropy measure of CHSH nonlocality for Werner states ~$\widetilde{\mathcal{M}}_{Re}=\mathcal{M}_{Re}(\rho_{w})/\mathcal{M}_{Re}^{max}$ as a function of the $w$ parameter is reported in Fig.
-2.
-Figure 2: Normalized geometric measures of CHSH nonlocality for Werner states as a function of the Werner parameter $w$ in the interval $\frac{1}{\sqrt{2}}\leq w\leq 1$.
-Black dashed straight line: Trace measure ~$\widetilde{\mathcal{M}}_{Tr}$.
-Grey dashed curve: Hilbert-Schmidt measure ~$\widetilde{\mathcal{M}}_{HS}$.
-Black solid curve: Relative entropy measure ~$\widetilde{\mathcal{M}}_{Re}$.
-Grey solid curve: Hellinger measure ~$\widetilde{\mathcal{M}}_{He}$.
-All plotted quantities are dimensionless.
-VI.2 Bell-diagonal states
-VI.2.1 Hilbert-Schmidt distance
-All Bell-diagonal states can be represented as points in a tetrahedron, as shown in Fig.
-1, with vertices representing Bell states.
-The points are given by three dimensional vectors $\vec{\alpha}$ that appear in Eq.
-(57).
-This equation implies that the Euclidean distance in the space represented in Fig.
-1 coincides with the HS distance in the space of density matrices of Bell-diagonal states.
-Moreover, for all Bell-diagonal states the closest local state is a Bell-diagonal state, as proved in Proposition 2, Eq.
-(53).
-The set of local states is represented by points in the solid figure inscribed inside the tetrahedron, see Fig.
-1.
-Hence, the Euclidean distance between the points representing nonlocal Bell-diagonal states and the boundary of the solid figure yields the HS-based geometric measure of CHSH nonlocality.
-Exploiting the Horodecki locality criterion Eq.
-(43), the separation surface between local and nonlocal Bell-diagonal state is identified by the solutions of the equation
-$\max_{i,ji\neq j}\{a_{i}^{2}+a_{j}^{2}\}=1\,.$ (67)
+Werner [10], is important as it provides the paramount example of a set of states that for $\frac{1}{3}\leq w\leq\frac{1}{\sqrt{2}}$ are entangled but nevertheless local.
+Appendix B Bell-diangonal states
+B.1 Hilbert-Schmidt distance
 Combining the above with Proposition 2, Eq. (32), the minimum of the squared HS distance is determined as
-$\min_{\rho_{BD}^{loc}\in\sigma}\left\|\rho_{BD}-\rho_{BD}^{loc}\right\|_{HS}^{2}=\min_{a_{i}^{loc}}\frac{1}{2}\displaystyle\sum_{i=1}^{3}(a_{i}-a_{i}^{loc})^{2}\,,$ (68)
-with the provision that the states $\rho_{BD}$ belong to the surface Eq. (67).
+$\min_{\rho_{\mathrm{BD}}^{loc}\in\mathcal{L}}\left\|\rho_{\mathrm{BD}}-\rho_{\mathrm{BD}}^{loc}\right\|_{2}=\min_{a_{i}^{loc}}\frac{1}{2}\sqrt{\displaystyle\sum_{i=1}^{3}(a_{i}-a_{i}^{loc})^{2}}\,$, (71)
+with the provision that the states $\rho_{\mathrm{BD}}$ belong to the surface Eq. (50).
 Considering first the case $|a_{1}^{loc}|,|a_{2}^{loc}|>|a_{3}^{loc}|$, we can proceed to determine the minimum by introducing a Lagrange multiplier $\lambda$.
-The associated Lagrange variational function reads
+The associated Lagrange variational (squared) function reads
 $\Gamma(a_{1}^{loc},a_{2}^{loc},a_{3}^{loc},\lambda)\equiv\displaystyle\sum_{i=1}^{3}(a_{i}-a_{i}^{loc})^{2}+\lambda[(a_{1}^{loc})^{2}+(a_{2}^{loc})^{2}-1]\,.$ (72)
-Vanishing of the gradient of $\Gamma$ leads to
-$\begin{cases}(a_{1}-a_{1}^{loc})+\lambda a_{1}^{loc}=0,\\ (a_{2}-a_{2}^{loc})+\lambda a_{2}^{loc}=0,\\ (a_{3}-a_{3}^{loc})=0,\\ (a_{1}^{loc})^{2}+(a_{2}^{loc})^{2}-1=0.\end{cases}$
+The vanishing of the gradient of $\Gamma$ leads to
+$\left\{\begin{array}[]{rcl}(a_{1}-a_{1}^{loc})+\lambda a_{1}^{loc}&=&0,\\ (a_{2}-a_{2}^{loc})+\lambda a_{2}^{loc}&=&0,\\ (a_{3}-a_{3}^{loc})&=&0,\\ (a_{1}^{loc})^{2}+(a_{2}^{loc})^{2}-1&=&0.\end{array}\right$.
+(73)
 Denoting the minima by $\tilde{a}_{i}^{loc}$, the solution reads
-$a_{1}^{loc}=\pm\frac{a_{1}}{\sqrt{a_{1}^{2}+a_{2}^{2}}}{2},$
-$a_{2}^{loc}=\pm\frac{a_{2}}{\sqrt{a_{2}^{2}+a_{3}^{2}}},$
+$a_{1}^{loc}=\pm\frac{a_{1}}{\sqrt{a_{1}^{2}+a_{2}^{2}}},$
+$a_{2}^{loc}=\pm\frac{a_{2}}{\sqrt{a_{2}^{1}+a_{2}^{2}}},$
 $a_{3}^{loc}=a_{3}.$ (74)
 By Eq. (71) the (un-normalized) HS geometric measure of CHSH nonlocality for Bell-diagonal states is thus
-$\mathcal{M}_{HS}(\rho_{BD})=\frac{3}{4}\sum\limits_{i=x,y,z}(a_{i}-a_{i}^{loc})^{2}=\frac{3}{4}\left(\sqrt{a_{1}^{2}+a_{2}^{2}}\mp 1\right)^{2}$ (71)
-whose minimum clearly correspond to the minus sign (the plus sign in (71)), since $t_{x}^{2}+t_{y}^{2}\geq 1$.
-The remaining cases $|a_{2}^{loc}|,|a_{3}^{loc}|>|a_{1}^{loc}|$ and $|a_{1}^{loc}|,|a_{3}^{loc}|>|a_{2}^{loc}|$ can be immediately found by symmetry and the global minimum is determined over all possible cases.
-For the maximally nonlocal Bell states the minimum is attained when $|a_{1}^{loc}|=|a_{2}^{loc}|=|a_{3}^{loc}|\equiv w^{2}$, with $w$ the Werner parameter.
-The constraints in Eq.
-(69) thus reduce to $12w^{2}=$ 1, i.e.
-$w=1/\sqrt{2}$, and $\mathcal{M}_{HS}$ achieves its maximum $\mathcal{M}_{HS}^{max}$, Eq.
-(59).
-Note that this analytical expression is not always valid, as no constraint has been imposed on the positivity of the density operator associated with the set $\{a_{i}^{loc}\}$.
-Therefore, in order to use expression (71), one must also take such constraints into account.
-If the local state does not satisfy these constraints, the distance must be computed numerically.
+$\mathcal{M}_{\mathrm{HS}}(\rho_{\mathrm{BD}})=\frac{1}{2}\sqrt{\sum\limits_{i=x,y,z}(a_{i}-a_{i}^{loc})^{2}}=\frac{1}{2}\left(\sqrt{a_{1}^{2}+a_{2}^{2}}\mp 1\right)$ (75)
+whose minimum clearly corresponds to the minus sign (the plus sign in (74)), since $a_{x}^{2}+a_{y}^{2}\geq 1$.
+It should be noted that the solution (74) is acceptable only if it satisfies the condition $|a_{1}^{loc}|,|a_{2}^{loc}|>|a_{3}^{loc}|$.
+The remaining cases $|a_{2}^{loc}|,|a_{3}^{loc}|>|a_{1}^{loc}|$ and $|a_{1}^{loc}|,|a_{3}^{loc}|>|a_{2}^{loc}|$ can be immediately found by symmetry, and the global minimum is determined over all possible cases, provided that the corresponding $|a_{i}^{loc}|$ are correctly ordered.
+It may happen that none of the three cases gives rise to a coherent solution; in this case, it is not possible to express the HS measure in the form (75).
+This is the case of Werner states, where $|a_{i}|=w$, but the solution (74) implies $|a_{1}^{loc}|,|a_{2}^{loc}|=1/\sqrt{2}<|a_{3}^{loc}|$.
+In this case the correct solution is indeed different from (75) (see (41)).
 Using Eqs.
 (69) and (70) the transformation from the matrix elements $\{a_{i}\}$ to the eigenvalues $\{e_{i}\}$ of the density matrix follows immediately.
-The behavior of the normalized HS measure of CHSH nonlocality ~$\widetilde{\mathcal{M}}_{HS}=\mathcal{M}_{HS}/\mathcal{M}_{HS}^{max}$ for Bell-diagonal states as a function of the eigenvalues $e_{1}$, $e_{2}$ in the case of $e_{4}=0$ is reported in Fig.
+The behavior of the normalized HS measure of CHSH nonlocality $\widetilde{\mathcal{M}}_{\mathrm{HS}}=\mathcal{M}_{\mathrm{HS}}/\mathcal{M}_{\mathrm{HS}}^{max}$ for Bell-diagonal states as a function of the eigenvalues $e_{1}$, $e_{2}$ in the case of $e_{4}=0$ is reported in Fig.
 3.
-Figure 3: Normalized Hilbert-Schmidt measure ~$\widetilde{\mathcal{M}}_{HS}$ of CHSH nonlocality for Bell-diagonal states as a function of the two free parameters $e_{1}$ and $e_{2}$, with $e_{4}=0$, $e_{3}=1-e_{1}-e_{2}$, and normalization coefficient $\mathcal{M}_{HS}^{max}$.
-The maximum ~$\widetilde{\mathcal{M}}_{HS}=1$ is reached in the three cases $(e_{1},e_{2},e_{3})=(1,0,0)$; $(e_{1},e_{2},e_{3})=(0,1,0)$; and $(e_{1},e_{2},e_{3})=(0,0,1)$.
-Each case corresponds to a different Bell state.
-The central region of minima is composed by local states, i.e.
-states with ~$\widetilde{\mathcal{M}}_{HS}=0$.
-All plotted quantities are dimensionless.
-VI.2.2 Hellinger distance
+B.2 Hellinger distance
 Each of the eigenvalues $\{e_{i}^{k}\}$ of a density matrix $\rho_{k}$ is the probability that a system in state $\rho_{k}$ collapses in the eigenstate relative to that eigenvalue.
-For a Bell-diagonal state $\rho_{BD}$ the eigenstates are obviously the four Bell states.
+For a Bell-diagonal state $\rho_{\mathrm{BD}}$ the eigenstates are obviously the four Bell states.
 Recalling Proposition 2, Eq.
 (32), and Eq.
-(23) the minimization process on the Hellinger distance squared yields the Hellinger geometric measure of CHSH nonlocality $\mathcal{M}_{He}(\rho_{BD})$ for Bell-diagonal states:
-$\mathcal{M}_{He}(\rho_{BD})=\min_{\rho_{BD}^{loc}\in\sigma}D^{2}_{He}(\rho_{BD},\rho_{BD}^{loc})=\min_{\rho^{loc}\in\sigma}\left\{\displaystyle\sum_{i=1}^{4}\bigg{(}\sqrt{e_{i}}-\sqrt{e_{i}^{loc}}\bigg{)}^{2}\right\}\,.$ (72)
+(11) the minimization process on the Hellinger distance squared yields the Hellinger geometric measure of CHSH nonlocality $\mathcal{M}_{\mathrm{He}}(\rho_{\mathrm{BD}})$ for Bell-diagonal states:
+$\mathcal{M}_{\mathrm{He}}(\rho_{\mathrm{BD}})=\min_{\rho_{\mathrm{BD}}^{loc}\in\mathcal{L}}D^{2}_{\mathrm{He}}(\rho_{\mathrm{BD}},\rho_{\mathrm{BD}}^{loc})=\min_{\rho^{loc}\in\mathcal{L}}\left\{\displaystyle\sum_{i=1}^{4}\bigg(\sqrt{e_{i}}-\sqrt{e_{i}^{loc}}\bigg)^{2}\right\}\,.$ (76)
 Using Eq. (70) we can rewrite the condition Eq. (50) in terms of the eigenvalues $\{e_{i}^{loc}\}$. Fixing, for instance, the case $|a_{1}|,|a_{2}|>|a_{3}|$, one has:
 $(e_{1}-e_{3})^{2}+(e_{2}-e_{4})^{2}\,>\,\frac{1}{2}\,.$ (77)
-For probabilities, their positivity $e_{i}>0$ and normalization condition $\sum_{i}e_{i}=1$ define a clove of a unitary $4-$sphere in the $\sqrt{e_{i}}-$space.
+For probabilities, their positivity $e_{i}>0$ and normalization condition $\sum_{i}e_{i}=1$ define a clove of a unitary 44-sphere in the $\sqrt{e_{i}}$-space.
 Therefore, the set of nonlocal states with $|a_{1}|,|a_{2}|>|a_{3}|$ is represented in the $\sqrt{e_{i}}$-space by the intersection of the hypervolume Eq.
@@
 We can thus proceed by the method of Lagrange multipliers and introduce the following Lagrangian function
-$\Gamma(e_{i}^{loc},\lambda_{1},\lambda_{2})\equiv\displaystyle\sum_{i=1}^{4}\sqrt{e_{i}e_{i}^{loc}}+\lambda_{1}\bigg{[}(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}-\frac{1}{2}\bigg{]}+\lambda_{2}\left(\displaystyle\sum_{i=1}^{4}e_{i}^{loc}-1\right)\,.$ (75)
+$\Gamma(e_{i}^{loc},\lambda_{1},\lambda_{2})$ $\equiv$ $\displaystyle\sum_{i=1}^{4}\sqrt{e_{i}e_{i}^{loc}}+\lambda_{1}\bigg[(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}-\frac{1}{2}\bigg]$ (79)
+$+$ $\lambda_{2}\left(\displaystyle\sum_{i=1}^{4}e_{i}^{loc}-1\right)\,.$
 Taking the gradient of $\Gamma$ and imposing its vanishing yields
-$\begin{cases}\frac{1}{2}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}+2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0\\ \vspace{0.2cm}\par\frac{1}{2}\sqrt{\frac{e_{2}}{e_{2}^{loc}}}+2\lambda_{1}(e_{2}^{loc}-p_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{1}{2}\sqrt{\frac{e_{3}}{e_{3}^{loc}}}-2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\par\frac{1}{2}\sqrt{\frac{e_{4}}{e_{4}^{loc}}}-2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\ \vspace{0.2cm}\par\par(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}\,.\end{cases}$
+$\left\{\begin{array}[]{l}\frac{1}{2}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}+2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\[6.0pt] \frac{1}{2}\sqrt{\frac{e_{2}}{e_{2}^{loc}}}+2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\[6.0pt] \frac{1}{2}\sqrt{\frac{e_{3}}{e_{3}^{loc}}}-2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\[6.0pt] \frac{1}{2}\sqrt{\frac{e_{4}}{e_{4}^{loc}}}-2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\[6.0pt] e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\[6.0pt] (e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\end{array}\right$.
+(80)
 From the above relations one immediately determines the Lagrange multipliers:
@@
 so that
-$\begin{cases}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}+\sqrt{\frac{e_{3}}{e_{3}^{loc}}}=\sqrt{\frac{e_{2}}{e_{2}^{loc}}}+\sqrt{\frac{e_{4}}{e_{4}^{loc}}}\,,\\ \vspace{0.2cm}\par\par\bigg{(}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}-\sqrt{\frac{e_{3}}{e_{3}^{loc}}}\bigg{)}(e_{2}^{loc}-e_{4}^{loc})=\bigg{(}\sqrt{\frac{e_{2}}{e_{2}^{loc}}}-\sqrt{\frac{e_{4}}{e_{4}^{loc}}}\bigg{)}(e_{1}^{loc}-e_{3}^{loc})\,,\\ \vspace{0.2cm}\par\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1\,,\\ \vspace{0.2cm}\par\par(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}\,.\end{cases}$
-The remaining cases can be treated symmetrically; the absolute minimum is determined by the smallest value with respect to all the possible cases.
+$\left\{\begin{array}[]{l}\sqrt{\frac{e_{1}}{e_{1}^{loc}}}+\sqrt{\frac{e_{3}}{e_{3}^{loc}}}=\sqrt{\frac{e_{2}}{e_{2}^{loc}}}+\sqrt{\frac{e_{4}}{e_{4}^{loc}}}\,,\\[6.0pt] \bigg(\sqrt{\frac{e_{1}}{e_{1}^{loc}}}-\sqrt{\frac{e_{3}}{e_{3}^{loc}}}\bigg)(e_{2}^{loc}-e_{4}^{loc})=\bigg(\sqrt{\frac{e_{2}}{e_{2}^{loc}}}-\sqrt{\frac{e_{4}}{e_{4}^{loc}}}\bigg)(e_{1}^{loc}-e_{3}^{loc})\,,\\[6.0pt] e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1\,,\\[6.0pt] (e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}\,.\end{array}\right$.
+(83)
+The remaining cases can be treated symmetrically; the absolute minimum is determined by the smallest value with respect to all possible cases.
 In the case of Bell states, e.g.
 $e_{1}=1$, $e_{2}=e_{3}=e_{4}=0$, it is straightforward to verify that the minimum is reached for $e_{1}^{loc}=\frac{1+3t}{4}$, $e_{2}^{loc}=e_{3}^{loc}=e_{4}^{loc}=\frac{1-t}{4}$ , with $t=\frac{1}{\sqrt{2}}$.
-Therefore, as it must be, we find that the maximum value $\mathcal{M}_{He}^{max}$ of the Hellinger geometric measure of CHSH nonlocality Eq.
+Therefore, as it must be, we find that the maximum value $\mathcal{M}_{\mathrm{He}}^{max}$ of the Hellinger geometric measure of CHSH nonlocality Eq.
 (76) coincides with that of the Hellinger measure for Werner states, Eq.
-(60), with Werner parameter $w=1$.
+(43), with the Werner parameter $w=1$.
 On the other hand, we see that the above minimization procedure cannot always yield a closed formula of the Hellinger measure of CHSH nonlocality for generic Bell-diagonal states.
-This is implied by the type of bound on the set of local states and the form of the function one needs to minimize.
-Indeed, the above algebraic systems yields that in order to find the minimum one must determine the roots of a polynomial of the fifth order.
-Therefore, from the Abel-Ruffini-Seralian theorem it follows that it is impossible to obtain a closed analytical formula by quadratures for the roots in terms of the coefficients of the polynomial.
+This is implied by the type of bound on the set of local states and the form of the function that one needs to minimize.
+Indeed, the above algebraic systems yield that in order to find the minimum one must determine the roots of a polynomial of the fifth order.
+Therefore, from the Abel-Ruffini-Seralian theorem, it follows that it is impossible to obtain a closed analytical formula by quadratures for the roots in terms of the coefficients of the polynomial.
 Of course, the solutions can be determined to any desired degree of precision by standard approximation procedures such as, e.g., the Newton-Raphson method or the Laguerre iterative scheme.
-Alternatively, we can proceed analytically in special but relevant instances, such as the Bell states or, as we will show later on, the Bell-diagonal states with one free parameter.
-VI.2.3 Bures distance
-In analogy with the case of Werner states, also in the case of Bell-diagonal states the density operators commute and the Bures and Hellinger distances coincide, so that
-$\mathcal{M}_{Bu}(\rho_{BD})=\mathcal{M}_{He}(\rho_{BD})\,.$ (78)
-VI.2.4 Relative entropy
-In the case of the relative entropy we assume that in order to determine the minimum in Eq.
+Alternatively, we can proceed analytically in special but relevant instances, such as the Bell states or the Bell-diagonal states with one free parameter.
+B.3 Bures distance
+In analogy with the case of Werner states, in the case of Bell-diagonal states as well, the density operators commute, and the Bures and Hellinger distances coincide, so that
+$\mathcal{M}_{\mathrm{Bu}}(\rho_{\mathrm{BD}})=\mathcal{M}_{\mathrm{He}}(\rho_{\mathrm{BD}})\,.$ (84)
+B.4 Relative entropy
+In the case of the relative entropy, we assume that in order to determine the minimum in Eq.
 (20) we can continue to impose the constraint Eq.
 (78).
-In other words, we expect that also for the relative entropy of nonlocality of Bell-diagonal states the closest local state belongs to the same surface as for the Hellinger and Bures measures.
+In other words, we expect that for the relative entropy of nonlocality of Bell-diagonal states, the closest local state belongs to the same surface as for the Hellinger and Bures measures.
 Assuming the above hypothesis, the minimum in Eq.
@@
 achieves its maximum.
-Fixing as usual, to begin with, the case $|a_{1}|,|a_{2}|>|a_{3}|$, the Lagrange function associated to the Lagrange multipliers takes the form
-$\Gamma(e_{i}^{loc},\lambda_{1},\lambda_{2})\equiv\displaystyle\sum_{i=1}^{4}e_{i}\log_{2}{e_{i}^{loc}}+\lambda_{1}[(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}-\frac{1}{2}]+\lambda_{2}\left(\displaystyle\sum_{i=1}^{4}e_{ij}^{loc}-1\right)\,,$ (80)
+Fixing, as usual, to begin with the case $|a_{1}|,|a_{2}|>|a_{3}|$, the Lagrange function associated with the Lagrange multipliers takes the form
+$\Gamma(e_{i}^{loc},\lambda_{1},\lambda_{2})$ $\equiv$ $\displaystyle\sum_{i=1}^{4}e_{i}\log_{2}{e_{i}^{loc}}+\lambda_{1}[(e_{1}^{loc}-e_{3}^{loc})^{2}+(e_{2}^{loc}-e_{4}^{loc})^{2}-\frac{1}{2}]$ (86)
+$+$ $\lambda_{2}\left(\displaystyle\sum_{i=1}^{4}e_{ij}^{loc}-1\right)\,$,
 which leads to
-$\begin{cases}\frac{e_{1}}{e_{1}^{loc}\ln{2}}+2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{p_{01}}{e_{2}^{loc}\ln{2}}+2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{e_{3}}{e_{3}^{loc}\ln{2}}-2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par\frac{e_{4}}{e_{4}^{loc}\ln{2}}-2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\ \vspace{0.2cm}\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\ \vspace{0.2cm}\par(e_{1}^{loc}-e_{2}^{loc})^{2}+(e_{3}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\end{cases}$
+$\left\{\begin{array}[]{l}\displaystyle\frac{e_{1}}{e_{1}^{loc}\ln 2}+2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\[6.0pt] \displaystyle\frac{p_{01}}{e_{2}^{loc}\ln 2}+2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\[6.0pt] \displaystyle\frac{e_{3}}{e_{3}^{loc}\ln 2}-2\lambda_{1}(e_{1}^{loc}-e_{3}^{loc})+\lambda_{2}=0,\\[6.0pt] \displaystyle\frac{e_{4}}{e_{4}^{loc}\ln 2}-2\lambda_{1}(e_{2}^{loc}-e_{4}^{loc})+\lambda_{2}=0,\\[6.0pt] e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\[6.0pt] (e_{1}^{loc}-e_{2}^{loc})^{2}+(e_{3}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\end{array}\right$.
 It is straightforward to show that the above relations reduce to
-$\begin{cases}\frac{e_{1}}{e_{1}^{loc}}+\frac{e_{3}}{e_{3}^{loc}}=\frac{e_{2}}{e_{2}^{loc}}+\frac{e_{4}}{e_{4}^{loc}},\\ \vspace{0.2cm}\par\bigg{(}\frac{e_{3}}{e_{3}^{loc}}-\frac{e_{1}}{e_{1}^{loc}}\bigg{)}(e_{2}^{loc}-e_{4}^{loc})=\bigg{(}\frac{e_{4}}{e_{4}^{loc}}-\frac{e_{2}}{e_{2}^{loc}}\bigg{)}(e_{1}^{loc}-e_{3}^{loc}),\\ \vspace{0.2cm}\par e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\ \vspace{0.2cm}\par(e_{1}^{loc}-e_{2}^{loc})^{2}+(e_{3}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\par\end{cases}$
-In analogy with the case of the Hellinger measure, the above algebraic system is of higher order and not amenable to a solution by elementary quadratures.
-Therefore, in the following subsection we will consider the particular instance of one-parameter Bell-diagonal states that allows for the full analytical evaluation of the entire set of geometric and entropic measures of CHSH nonlocality.
-VI.3 Bell-diagonal states with one free parameter
-Let us consider Bell-diagonal states that are convex combinations of only two of the four Bell states.
-For instance, let us choose $e_{2}=e_{4}=0$ , $e_{1}=p$ , and $e_{3}=1-p$ , with $p$ free parameter in the interval $[0,1]$.
-Such Bell-diagonal states represent all possible convex combinations of the first and third Bell states of Eq.
-(37).
-Moving to symplectic space and recalling Eqs.
-(42), this instance corresponds to $a_{1}=a$, $a_{2}=-a$, and $a_{3}=1$, with $a=2p-1$.
-For this class of states the locality condition Eq.
-(43) is fulfilled only for $p=\frac{1}{2}$; therefore, such states are nonlocal for $0\leq p\leq 1$, with the exception of $p=\frac{1}{2}$.
-The behavior of the four normalized measures of CHSH nonlocality as functions of the convex combination parameter $p$ in the half-interval $p\geq\frac{1}{2}$ is reported in Fig.
-4.
-We see that all measures vanish at $p=\frac{1}{2}$ and reach their maximum at $p=1$, i.e.
-on the first Bell state.
-For all measures, the normalization coefficients turn out to be the same as the ones of the nonlocality measures for Werner states, in agreement with Proposition 1, Eq.
-(46), and Proposition 2, Eq.
-(53).
-The case $p\leq\frac{1}{2}$ is readily obtained by symmetry.
-It is straightforward to verify that the above-described behavior is generic, i.e. it does not depend on the pair of Bell states that are chosen as entries of the convex combination.
-Figure 4: Normalized measures of CHSH nonlocality for the convex combination of two Bell states as functions of the convex combination parameter $p$ in the interval $p\geq\frac{1}{2}$.
-The same behavior holds in the symmetric case $p\leq\frac{1}{2}$, with the maximum achieved in $p=0$.
-Black dashed curve: Trace measure ~$\widetilde{\mathcal{M}}_{Tr}$.
-Black solid curve: Relative entropy measure ~$\widetilde{\mathcal{M}}_{Re}$.
-Grey dashed curve: Hellinger measure ~$\widetilde{\mathcal{M}}_{He}$.
-Grey solid curve: Hilbert-Schmidt measure ~$\widetilde{\mathcal{M}}_{HS}$.
-All plotted quantities are dimensionless.
-VII Discussion and outlook
-We introduced a geometric measure for Bell nonlocality in a fixed scenario, whose defining elements are the state vectors in Hilbert space and their distances from the set of local states.
-In the case of the CHSH inequality, we have identified the basic traits of such approach, including the set of local states and the free operations, within the framework based on Local Operations and Shared Randomness (LOSR), and we have discussed the concept of maximally nonlocal states Buscemi2012 ; Wolfe2020 ; Schmid2020b ; Rosset2020 ; Lipka2021 ; Gallego2017 .
-The quantification of Bell nonlocality for quantum states in Hilbert space is naturally introduced in terms of geometric measures, in close methodological analogy with the standard theory of geometric entanglement Wei2003 ; Uyanik2010 and the more recent geometric approaches to the quantification of other types of nonlocal correlations such as, e.g., the quantum discord Geomdiscord2016 ; Geomdiscord2017 .
-Geometric measures define a structure and an ordering in the space of quantum states, and we have provided a detailed investigation of such structure in the case of the CHSH inequality for different classes of contractive distances (trace, Hellinger, Bures), in the case of two-qubit states, by proving and exploiting the result that for contractive distances the closest local state to a Werner state is still a Werner state, and that an analogous result holds for Bell-diagonal states.
-We have confronted the ordering provided by contractive geometric measures of CHSH nonlocality with the non-contractive Hilbert-Schmidt distance, and we have introduced the relative entropy of nonlocality as a further benchmark.
-An important aspect of our approach is that, within a fixed Bell scenario (defined by the number of measurements, outcomes, and a Bell inequality), we perform a maximization over all possible measurement settings compatible with the given constraints.
-This procedure ensures that geometric measures of nonlocality depend only on the quantum state and the structure of the chosen Bell scenario, and not on the particular experimental setting.
-On the other hand, one could envisage an alternative strategy, in which also the measurement settings are fixed from the outset.
-In this case, the set of local states would be defined relative to these fixed measurements, and the resulting geometric measure would capture the nonlocality of a state specifically with respect to the selected experimental configuration, without any apriori extremization.
-Such a fixed-measurements approach might offer advantages in operational contexts, where the measurements are experimentally constrained or predetermined.
-Generalizations and applications of the conceptual framework introduced in the present work are conceivable along different directions, such as the identification of the set of local states and the ordering of nonlocal states in finite-dimensional Hilbert spaces of higher dimension and generalized Bell inequalities, the extension to field-theoretical settings of elementary particle physics, and the use of the Hilbert space-based measures of Bell nonlocality as tools in the investigation and characterization of quantum collective phenomena and quantum phase transitions.
-Concerning the latter, a characterization of the ground states of quantum many-body systems in terms of Hilbert-space-based measures of Bell nonlocality could provide novel and potentially profound insights on the degree of quantumness and quantumness hierarchies between different systems of quantum matter, in analogy with and complementing the very successful use of entanglement in condensed matter theory Laflorencie2016 ; Illuminati2022 ; Illuminati2023 .
-Indeed, some recent preliminary studies on the structure of Bell nonlocality in quantum many-body systems hint at complex patterns of quantum nonlocality beyond entanglement Nonlocmanybody2015 ; Belldepthmanybody2019 whose precise entropic and geometric quantifications might provide important insights on the structure and hierarchy of quantum collective phenomena.
-The Hilbert-space characterization and quantification that we have introduced paves the way to the study of several important questions concerning quantum nonlocality.
-In entanglement theory, it is known that different entanglement measures can give rise to different orderings of quantum states; for instance, the entropy-based entanglement of formation and the entanglement negativity based on the PPT criterion induce a different ordering on the set of entangled Gaussian states Adesso2005 .
-Similar ordering issues might arise when determining the degree of Bell nonlocality of quantum states as quantified by the different geometric or entropic measures that we have introduced in the present work.
-A further important question concerns the characterization and quantification of extremal nonlocality in mixed states.
-In entanglement theory, maximally entangled mixed states are typically defined at fixed energy, or fixed degree of local and global entropies Adesso2003 ; Adesso2004 .
-An investigation along similar lines could lead to the identification of the maximally nonlocal mixed states at fixed global and/or local purities.
-In our opinion, a particularly challenging open problem whose eventual solution would bear far-reaching consequences revolves around the question whether the analogs of Werner states, i.e., mixed states that are local and yet entangled, exist in infinite-dimensional Hilbert spaces.
-In particular, it would be very interesting to determine the existence of Werner Gaussian states of two- and multimode continuous-variable systems, and to characterize their properties.
-Acknowledgements.
-W.R.
-and M.T.
-are supported by JST Moonshot R&D Grant No.
-JPMJMS2061 and JST COI-NEXT Grant No.
-JPMJPF2221.
-F.I.
-acknowledges funding from the Italian Ministry of University and Research, call PRIN PNRR 2022, project “Harnessing topological phases for quantum technologies”, code P202253RLY, CUP D53D23016250001, and PNRR-NQSTI project ”ECoN: End-to-end long-distance entanglement in quantum networks”, CUP J13C22000680006;
+$\left\{\begin{array}[]{l}\displaystyle\frac{e_{1}}{e_{1}^{loc}}+\frac{e_{3}}{e_{3}^{loc}}=\frac{e_{2}}{e_{2}^{loc}}+\frac{e_{4}}{e_{4}^{loc}},\\[6.0pt] \displaystyle\bigg(\frac{e_{3}}{e_{3}^{loc}}-\frac{e_{1}}{e_{1}^{loc}}\bigg)(e_{2}^{loc}-e_{4}^{loc})=\bigg(\frac{e_{4}}{e_{4}^{loc}}-\frac{e_{2}}{e_{2}^{loc}}\bigg)(e_{1}^{loc}-e_{3}^{loc}),\\[6.0pt] e_{1}^{loc}+e_{2}^{loc}+e_{3}^{loc}+e_{4}^{loc}=1,\\[6.0pt] (e_{1}^{loc}-e_{2}^{loc})^{2}+(e_{3}^{loc}-e_{4}^{loc})^{2}=\frac{1}{2}.\end{array}\right$.
+In analogy with the case of the Hellinger measure, the above algebraic system is of a higher order and not amenable to a solution by elementary quadratures.
+Therefore, in the following subsection, we will consider the particular instance of one-parameter Bell-diagonal states that allow for the full analytical evaluation of the entire set of geometric and entropic measures of CHSH nonlocality.
+Appendix C Numerical Methods for the Estimation of the CHSH-Nonlocality Measures
+C.1 Geometric measures
+In this section, we describe the procedures used to estimate the geometric and entropic nonlocality measures for Bell-diagonal states in cases where closed analytical formulas are not available.
+Recall that Bell-diagonal states are convex combinations of the four Bell states and are therefore characterized by three free parameters, namely the independent probabilities $e_{1}$, $e_{2}$, and $e_{3}$, since the fourth probability is fixed by normalization as $e_{4}=1-e_{1}-e_{2}-e_{3}$.
+Alternatively, one can adopt the parametrization (30) in terms of the vector $\mathbf{a}=(a_{1},a_{2},a_{3})$.
+For the numerical analysis, the choice of parametrization can be made according to convenience, since for different distances a suitable parametrization can simplify the minimization procedure.
+Symmetric families.
+For Werner and isotropic states, we prove that the closest local state remains within the same family.
+This property allows us to derive explicit analytical expressions for the Hilbert-Schmidt, Hellinger, Bures, trace, and relative entropy measures of CHSH nonlocality (see Sections 5 and 6 of the main text).
+The formulas obtained for the two-qubit case are also employed as benchmarks to validate the numerical procedures described below.
+Bell-diagonal states.
+For general Bell-diagonal states, the minimization cannot be solved analytically, since it is not guaranteed a priori that the candidate minimum $\rho_{\mathrm{loc}}$ is a physical density operator.
+In particular, its positivity must be explicitly verified.
+The local set $\sigma$ is convex, so in principle the problem is convex once expressed in a suitable parametrization.
+In practice, however, the locality constraints lead to nonlinear and piecewise boundary conditions, which make direct global minimization unreliable.
+To provide concrete estimation methods, we implemented a numerical strategy based on the SciPy optimization library. Specifically:
+  • The parameters of Bell-diagonal states are expressed either in terms of the probabilities $e_{i}$ or the correlators $(a_{1},a_{2},a_{3})$, with explicit routines to convert between the two parameterizations.
+  • Locality constraints are encoded through inequalities on the $t_{i}$ or $e_{i}$, ensuring that the candidate $\rho^{loc}\in\mathcal{L}$ corresponds to a valid density operator.
+  • Since the boundary of $\sigma$ consists of disjoint surfaces, the minimization is performed separately on each surface, and the smallest value is retained.
+This piecewise approach improves robustness.
+  • Optimization is carried out with minimize (SLSQP or trust-constr), with bounds [1,1][-1,1] on correlators and $[0,1]$ on probabilities.
+Multiple initial seeds (symmetric choices and random perturbations) are used to avoid local traps.
+  • For entropic measures, the minimization is performed directly in the probability space, with constraints enforcing normalization and positivity.
+C.2 Numerical procedure
+For Bell-diagonal states, we considered two different restricted families.
+First, we studied convex combinations of two Bell states, which depend on a single free parameter.
+Second, we analyzed convex combinations of three Bell states, which involve two free parameters, with the fourth probability fixed at zero by construction.
+In the numerical minimization, the Hilbert-Schmidt measure is evaluated in the $(a_{1},a_{2},a_{3})$ parametrization, while for the other measures (Hellinger, Bures, trace, and relative entropy) the probability parametrization $(e_{i})$ is used.
+Since Werner states form a subclass of Bell-diagonal states in the two-qubit case, we use the available analytical formulas for this family as benchmarks to validate our numerical procedures.
+This strategy provides concrete and stable estimates of the nonlocality measures and enables the generation of contour plots over the corresponding parameter spaces.
+C.3 Validation and accuracy
+To ensure correctness:
+  1. 1.
+    We benchmark against the analytical solutions for Werner and isotropic states, reproducing the exact values within tolerance $10^{-6}$.
+  2. 2.
+    We check grid convergence by refining the discretization (e.g. $n=10,20,50$ points per parameter) and verifying stability of the estimates.
+  3. 3.
+    We verify that different initializations lead to the same minimum, confirming consistency of the optimization.
+All Python codes used for these computations are available on GitHub: https://github.com/2zanfardino/geometric_nonlocality.
+Typical execution times for two-parameter contour plots (grid size $50\times 50$) are a few minutes on a standard workstation.